Enriched Functors and Stable Homotopy Theory

Bjørn Ian Dundas, Oliver Röndigs, Paul Arne Østvær

Received: September 24, 2002 Revised: August 7, 2003

Communicated by Ulf Rehmann

Abstract. In this paper we employ enriched category theory to construct a convenient model for several stable homotopy categories. This is achieved in a three-step process by introducing the pointwise, homotopy functor and stable model category structures for enriched functors. The general setup is shown to describe equivariant stable homotopy theory, and we recover Lydakis’ model category of simplicial functors as a special case. Other examples – including motivic homotopy theory – will be treated in subsequent papers.

2000 Mathematics Subject Classification: 55P42, 55U35, 18D20, 55P91 Keywords and Phrases: model structures for enriched functor categories, stable homotopy theory, equivariant stable homotopy theory

1 Introduction

An appropriate setting to study stable phenomena in algebraic topology is the stable homotopy theory of spectra described in [1]. More recently, much research has been focused on rebuilding the foundation of stable homotopy theory. There are now several equivalent model categories to study structured ring spectra and their modules. These frameworks are important in many aspects and make powerful tools from algebra applicable to “brave new rings”. For the purpose of this paper, the relevant constructions are those of symmetric spectra [9] and simplicial functors [11].

In [8], Hovey considers the notion of spectra for general model categories. This level of generality allows one to use techniques from stable homotopy theory in traditionally unrelated subjects. Of particular interest, where further applications are expected, is algebraic geometry and Voevodsky’s motivic stable homotopy category [16]. We are interested in an approach to this subject where all coherence problems which arise when one tries to make a smash product are encoded in the underlying category. This is different from the popular means of attack through symmetric spectra, where the controlling categories are much more restricted. Our point of view is analogous to Lydakis’ work [11] on simplicial functors as a model for ordinary spectra. But the theory we develop here is complicated by the fact that we do not assume properties which are particular to simplicial sets. Let us state a tentative version of the main theorem in this paper. Due to their technical nature, we defer on listing all the required assumptions. The basic input is a monoidal model category V and a finitely presentable cofibrant object T in V, the 1-sphere. See Sections 6 and 7 for precise statements.

Theorem 1. There is a monoidal model category (F, ∧, 𝕀) which satisfies the monoid axiom, and a right Quillen equivalence from F to the stable model category of T -spectra.

A T -spectrum is a sequence (E_{0}, E_{1}, …) of objects in V together with structure maps T ⊗ E_{n} E_{n+1}. An object in F is a functor X from a category of finitely presentable objects in V to V, which is “continuous” or enriched in the sense that for finitely presentable objects v and w there is a natural map

v ⊗ X(w) X(v ⊗ w).

Using this map, it follows that any enriched functor yields a T -spectrum by evaluating at spheres T^{⊗n}. We show that the induced functor from F to the category of T -spectra is a right Quillen equivalence. The monoidal structure is a special case of a result due to Day [4]. By construction, the sphere spectrum or unit 𝕀 is the inclusion of the subcategory of finitely presentable objects. For V the category of simplicial sets, Lydakis [11] has shown that F models the classical stable homotopy category. Our theorem extends this result to a wide range of model categories. In the sequel [5] we construct a model for Voevodsky’s motivic stable homotopy category. Motivic cohomology has a natural description as an algebra in this model. The monoid axiom implies that also categories of algebras and modules in F have model structures [15].

As a guide to this paper, it seems appropriate to summarize the content of each section. In Section 2 we recall categorical precursors and Day’s smash product for enriched functors. This material is included to make the paper reasonably self-contained and to set notation. Next we record a general isomorphism between enriched functor categories build from spheres and symmetric spectra. Moreover, under this isomorphism the corresponding smash products are shown to agree. Section 3 recalls some frequently used notions in homotopical algebra. An expert could skip most of this part. We introduce a class of model categories dubbed weakly finitely generated and show that weak equivalences and fibrant objects are closed under filtered colimits. Such a model structure is cofibrantly generated, with additional finiteness conditions on the generating cofibrations and acyclic cofibrations which are satisfied in many cases of interest. This introductory part ends with a discussion of fibrant replacement functors. Quillen’s small object argument is the usual device for replacing objects in a model category by fibrant objects. Some modifications are necessary in the enriched setting. If the monoid axiom holds and the model category is weakly finitely generated, we construct enriched fibrant replacements. Our constructions are primarily of a technical interest and might be omitted on a first cursory reading. However, we should remark that much of the following relies on this input. In the remaining sections we study homotopical algebra for enriched functor categories. First we construct the pointwise model structure where the fibrations and weak equivalences are defined pointwise. This gives an example of a weakly finitely generated model structure provided some weak assumptions are satisfied. In many cases of interest, we prove the monoid axiom and that smashing with a cofibrant object preserves weak equivalences. The latter result requires further assumptions on F, and is similar to the algebraic fact that tensoring with a projective module preserves short exact sequences. These two results are important for the model structures we construct later on. One drawback with the pointwise model structure is that it has far too many homotopy types. For example, a weak equivalence v ^{∼} w does not necessarily induce a pointwise weak equivalence V(w, −) V(v, −) on the level of representable functors. However, for all fibrant objects u in V the map V(w, u) V(v, u) is a weak equivalence. We therefore enlarge the class of pointwise weak equivalences by looking at fibrant objects as input only. The result is the homotopy functor model structure which has the same cofibrations as the pointwise model structure. The fibrant functors are precisely the pointwise fibrant functors which preserve weak equivalences, thus any enriched functor is weakly equivalent in the homotopy functor model structure to a homotopy functor. It seems to be of considerable interest to discuss a motivic version of Goodwillie’s calculus of functors. Let us remark that the homotopy functor model structure is a first step in this direction. In Section 6, the stable model structure is constructed by means of a general stabilization process. Theorem 6.26 lists conditions for the stable model structure to exist. The stable fibrations are characterized by pointwise homotopy pullback squares, and stable acyclic fibrations are precisely the pointwise acyclic fibrations. To prove these results we compare with spectra [8]. The stabilization we use does not coincide with the usual stabilization for spectra, and it requires some cruel details to compare them. These can be found in Appendix A. We note that the monoid axiom holds under an additional assumption on the source of the functor category. For a particular choice of the source category, which is explained in Section 7, the evaluation functor is the right adjoint in a Quillen equivalence. It follows that the highly structured category of enriched functors describes the same homotopy theory as spectra in many cases of interest. In Section 8, we give a short summary of the important algebraic consequences of the previous sections. In the last section we discuss equivariant homotopy theory for finite groups and we prove the following theorem. The general machinery gives deeper structure than stated, but we refer the reader to Section 9 for more details.

Theorem 2. Let G be a finite group. Then there is a monoidal model category (GF, ∧, 𝕊^{G}) satisfying the monoid axiom and a right Quillen equivalence from GF to the category of G-spectra.

The general framework may seem abstract, but we obtain a common footing for applications. A project in progress suggests that the approach anticipated in the present paper is relevant for the theory of motives. We hope the reader finds results herein which he or she can prove to have further applications.

2 Enriched Categories

This section contains an introduction to enriched categories, Day’s work on enriched functor categories [4], and simplicial homotopies in categories enriched over simplicial sets. In the last part we show that spectra and symmetric spectra are isomorphic to enriched functor categories build from spheres.

2.1 Introduction

Entry points to the literature on enriched category theory include [2] and [12]. A monoidal category consists of a category V, and

• a functor ⊗: V × V V and natural associativity isomorphisms

α_{A,B,C}: (A ⊗ B) ⊗ C A ⊗ (B ⊗ C)

subject to the coherence law [2, 6.1],

• an object e of V called the unit, and natural unit isomorphisms

l_{A}: e ⊗ A  A and r_{A}: A ⊗ e  A

such that [2, 6.2] holds.

The functor ⊗ is the tensor or monoidal product of V. A monoidal category is symmetric monoidal if there is a natural isomorphism σ_{A,B}: A ⊗ B B ⊗ A subject to the coherence laws [2, 6.3, 6.4, 6.5]. A symmetric monoidal category (V, ⊗, e) is closed if there exists a right adjoint Hom_{V}(A, −): V V to the endofunctor − ⊗ A for every object A of V. The categories of sets Set and pointed sets Set_{∗} are both closed symmetric monoidal categories. Let Δ denote the simplicial category. Its objects are the finite ordered sets [n] = {0 < 1 < ⋯ < n} for n ≥ 0, and morphisms are order-preserving maps. Consider sSet = Fun(Δ^{op}, Set), the category of simplicial sets. Its monoidal product is the categorical product, formed degree-wise in Set. If K, L ∈ sSet, the simplicial set of maps Hom_{sSet}(K, L) has n-simplices the set of maps from K × Δ^{n} to L. Here Δ^{n} is the simplicial set represented by [n]. The unit of the product is the terminal object Δ^{0}. Let (V, ⊗, e) be a closed symmetric monoidal category. Then a V-category C, or a category enriched over V, consists of a class Ob C of objects and

• for any pair (a, b) of objects in C, an object V_{C}(a, b) of V called the Vobject of maps in C,

• a composition V_{C}(b, c) ⊗ V_{C}(a, b) V_{C}(a, c), an identity or unit map e V_{C}(a, a) subject to the associativity and unit coherence laws listed in [2, 6.9 and 6.10].

Categories in the usual sense are the Set-categories. If C is a category, let Set_{C}(a, b) denote the set of maps in C from a to b. A closed symmetric monoidal category V is a V-category due to its internal Hom objects [2, 6.2.6]. Let V(A, B) denote the V-object Hom_{V}(A, B) of maps in V. Any V-category C defines a Set-category UC. Its class of objects is Ob C, the morphism sets are Set_{UC}(a, b) = Set_{V}(e, V_{C}(a, b)). For example, the Set-category obtained from a sSet-category C has morphism sets Set_{sSet}(Δ^{0}, sSet_{C}(a, b)) = sSet_{C}(a, b)_{0} the zero-simplices of the simplicial sets of maps. A V-functor F from C to D is an assignment from Ob C to Ob D together with morphisms hom_{a,b}^{F}: V_{C}(a, b) V_{D}(F (a), F (b)) in V which preserve composition and identities. A small sSet-category defines a simplicial object in the category Cat of small categories. With this description, a sSet-functor is a natural transformation of functors from Δ^{op} to Cat [11, 3.2]. If F and G are V-functors from C to D, a V-natural transformation t: F G consists of the following data: There is a morphism t(a): F (a) G(a) in UD for every a ∈ Ob C, and all the diagrams of the following form commute.

[figure: V_{C}(a, b); hom_{a,b}^{F}; V_{D}(F (a), F (b)); V_{D}(G(a), G(b)); hom_{a,b}^{G}; V_{D}(t(a),G(b)); V_{D}(F (a), G(b)); V_{D}(F (a),t(b))]

The V-natural isomorphisms and V-adjoint pairs of V-functors are defined as for V = Set. The adjoint pair of endofunctors (−⊗A, V(A, −)) on V explicate a V-adjoint pair. Denote the unit of the adjunction by η_{A}: Id_{V} V(A, −⊗A), and the counit by ϵ_{A}: V(A, −) ⊗ A Id_{V}. Details concerning hom^{−⊗A} and hom^{V(A,−)} can be found in Appendix A. Note that any V-functor F: C D gives a functor UF: UC UD with the same effect on objects as F, and similarly for V-natural transformations. That is, one can consider U as a 2-functor from the 2-category of small V-categories, V-functors and V-natural transformations to the 2-category of small categories, functors and natural transformations. A V-category C is called small provided UC is small. In fact, U is the base change along the lax symmetric monoidal functor Set_{V}(e, −): V Set, see [2, 6.4]. If no confusion can arise, we will omit U from the notation. The monoidal product C ⊗ D of two V-categories C and D is the V-category where Ob (C⊗D) := Ob C×Ob D and V_{C⊗D}((a, x), (b, y)) := V_{C}(a, b)⊗V_{D}(x, y). Note that the monoidal product in V induces a V-functor mon: V ⊗ V V. A V-category C is a right V-module if there is a V-functor act: C ⊗ V C, denoted (c, A) c⊘A and a V-natural unit isomorphism r_{c}: act(c, e) c subject to the following conditions.

• There are natural coherent associativity isomorphisms

act(c, A ⊗ B) act(act(c, A), B).

• The isomorphisms act(c, e ⊗ A) [figure] act(c, A) coincide.

A right V-module (C, act, r) is closed if there is a V-functor

coact: V^{op} ⊗ C C

such that for all A ∈ Ob V and c ∈ Ob C, the V-functor act(−, A): C C is left V-adjoint to coact(A, −) and act(c, −): V C is left V-adjoint to V_{C}(c, −). A monoidal V-category consists of a V-category C equipped with a V-functor ⋄: C⊗C C, a unit u ∈ Ob C, a V-natural associativity isomorphism and two V-natural unit isomorphisms satisfying the conditions mentioned for V = Set. Symmetric monoidal and closed symmetric monoidal V-categories are defined similarly.

2.2 Categories of enriched functors

If C is a small V-category, V-functors from C to V and their V-natural transformations form the category [C, V] of V-functors from C to V. If V is complete, then [C, V] is also a V-category. Denote this V-category by F(C), or F if no confusion can arise. The morphism V-object V_{F}(X, Y) is the end

∫_{Ob C} V(X(c), Y (c)).

See [2, 6.3.1] for details. Note that UF is [C, V]. One can compare F with C and V as follows: Given c ∈ Ob C, X X(c) defines the V-functor Ev_{c}: F V called “evaluation at c”. The assignment c V_{C}(c, −) from C to F is again a V-functor C^{op} F, called the V-Yoneda embedding [2, 6.3.6]. V_{C}(c, −) is a representable functor, represented by c.

Lemma 2.1 (Enriched Yoneda lemma). Let V be a complete closed symmetric monoidal category and C a small V-category. For every V-functor X: C V and every c ∈ Ob C, there is a V-natural isomorphism X(c) ≅ V_{F}(V_{C}(c, −), X).

The isomorphism in 2.1 is called the Yoneda isomorphism [2, 6.3.5]. It follows from 2.1 that every V-functor can be expressed as a colimit of representable functors [2, 6.6.13, 6.6.17]:

Lemma 2.2. If V is a bicomplete closed symmetric monoidal category and C is a small V-category, then [C, V] is bicomplete. (Co)limits are formed pointwise.

Corollary 2.3. Assume V is bicomplete, and let C be a small V-category. Then any V-functor X: C V is V-naturally isomorphic to the coend

∫ ^{Ob C} V_{C}(c, −) ⊗ X(c).

See [2, 6.6.18] for a proof of 2.3.

Proposition 2.4. Let V be a closed symmetric monoidal category, and let C be a small V-category. Then F is a closed V-module.

Proof. There is an obvious “pointwise” closed V-module structure. The Vfunctor F ⊗ V F defined by (X, A) (− ⊗ A) ∘ X gives the action of V on F. Next, the assignment (A, X) V(A, −) ∘ X defines the coaction. There are V-natural isomorphisms V_{F}((−⊗A)∘X, Y) ≅ V_{F}(X, V(A, −)∘Y) ≅ V(A, V_{F}(X, Y)) induced from the natural closed V-module structure on V. From this, a routine check finishes the proof.

Recall the notion of left Kan extensions:

Proposition 2.5. Fix a bicomplete closed symmetric monoidal category V, and a V-functor F: C D of small V-categories. For any V-functor X: C V, there exists a V-functor F_{∗}(X): D V and a V-natural isomorphism

V_{F(D)}(F_{∗}X, Y) ≅ V_{F(C)}(X, Y ∘ F).

In other words, there exists a V-adjoint pair of V-functors

F_{∗}: F(C) [figure] F(D): F^{∗}

where F^{∗} denotes pre-composition with F.

See [2, 6.7.7] for a proof of 2.5. The V-functor F_{∗}X is the left Kan extension of X along F. An explicit expression is given by the coend

F_{∗}X = ∫ ^{Ob C} V_{D}(F (c), −) ⊗ X(c).

2.3 Smash product of enriched functors

Let (C, ⋄, u) be a small symmetric monoidal V-category where V is bicomplete. In [4], B. Day constructed a closed symmetric monoidal product ∧ on the category [C, V] of V-functors from C to V. For X, Y ∈ Ob [C, V], there is the V-functor

X\overline{∧}Y: C ⊗ C ^{X⊗Y} V ⊗ V ^{mon} V.

The smash product X ∧ Y ∈ Ob [C, V] is the left Kan extension

⋄_{∗}(X\overline{∧}Y) = ∫ ^{Ob (C⊗C)} V_{C}(c ⋄ d, −) ⊗ (X(c) ⊗ Y (d)): C V.

The next result is a special case of [4, 3.3], cf. [4, 3.6, 4.1].

Theorem 2.6 (Day). Let (V, ⊗, e) be a bicomplete closed symmetric monoidal category and (C, ⋄, u) a small symmetric monoidal V-category. Then the category ([C, V], ∧, V_{C}(u, −)) is closed symmetric monoidal.

The V-category F of V-functors from C to V is also a closed symmetric monoidal V-category. The internal Hom functor, right adjoint to − ∧ X, is given by

F(X, Y)(c) = V_{F}(X, Y (c ⋄ −)) = ∫_{d∈Ob C} V(X(d), Y (c ⋄ d)).

Concerning smash products of representable functors, one has the following result.

Lemma 2.7. The smash product of representable functors is again representable. There is a natural isomorphism V_{C}(c, −) ∧ V_{C}(d, −) ≅ V_{C}(c ⋄ d, −).

In the following, a sub-V-category means a sub-V-category of V. We use 2.7 to define assembly maps if C is a full sub-V-category containing the unit and closed under the monoidal product. In this case, the inclusion 𝕀: C ^{⊂} V can be chosen as the unit of [C, V]. The composition X ∘ Y of two V-functors X, Y: C V is given by 𝕀_{∗}X ∘ Y. Up to coherent natural isomorphisms, the composition is associative with unit 𝕀.

Corollary 2.8. Given functors X and Y in [C, V], there exists a natural assembly map X ∧ Y X ∘ Y = 𝕀_{∗}X ∘ Y which is an isomorphism if Y is representable.

Proof. One can define the assembly map objectwise via the composition

𝕀_{∗}X(c) ⊗ Y (d) ^{sw_{Y (d)}^{𝕀∗X}(c)} 𝕀_{∗}X(Y (d) ⊗ c) ^{𝕀_{∗}X(sw_{c}^{Y} (d))} 𝕀_{∗}X(Y (c ⊗ d))

where sw_{c}^{Z}: Z ⊗ c Z(c ⊗ −) is ‘the’ natural map described in Appendix A. Here is another description via representable functors. Suppose X = V(c, −) and Y = V(d, −), for c, d ∈ Ob C. By 2.7, X ∧ Y is naturally isomorphic to V(c ⊗ d, −), i.e. to V(c, −) ∘ V(d, −). If X is arbitrary, it follows from 2.3 that X ∧ V(d, −) is naturally isomorphic to 𝕀_{∗}X ∘ V(d, −). If also Y is arbitrary, apply 2.3 and consider the natural map

∫ ^{Ob C}(𝕀_{∗}X ∘ V(c, −)) ⊗ Y (c) 𝕀_{∗}X ∘ ∫ ^{Ob C} V(c, −) ⊗ Y (c).

2.4 Categories enriched over simplicial sets

A functor F: V W of monoidal categories (V, ⊗, e) and (W, ⊗′, e′) is lax monoidal if there is a natural transformation t_{A,B}: F (A)⊗′F (B) F (A⊗B) and a morphism e′ F (e), such that the diagrams [2, 6.27, 6.28] commute. The word “lax” is replaced by “strict” if t_{A,B} is a natural isomorphism and e′ F (e) is an isomorphism. F is lax symmetric monoidal if t_{B,A}∘σ_{F A,F B} = F (σ_{A,B}) ∘ t_{A,B}. In this case, every W-category is a V-category by [2, 6.4.3]. We used this fact for the forgetful 2-functor U induced by Set_{V}(e, −). The assembly map makes Id_{[C,V]} into a lax monoidal functor from the monoidal category ([C, V], ∘, 𝕀) to the closed symmetric monoidal category ([C, V], ∧, 𝕀). Suppose that F: sSet V is a lax symmetric monoidal functor. One can then lift the notion of simplicial homotopy equivalence from sSet-categories to V-categories.

Definition 2.9. Let C be a sSet-category and f, f′: Δ^{0} sSet_{C}(c, d) maps in C. Then H: Δ^{1} sSet_{C}(c, d) is a simplicial homotopy from f to f′ if the following diagram commutes, where i_{0} and i_{1} are the canonical inclusions.

[figure: Δ^{0}; i_{0}; Δ^{1}; i_{1}; Δ^{0}; sSet_{C}(c, d); H; f′; f]

The map f is called a simplicial homotopy equivalence if there exists a map g: Δ^{0} sSet_{C}(d, c), and simplicial homotopies from g ∘ f to id_{c} and from f ∘ g to id_{d}. Let the symbol ≃ denote simplicial homotopy equivalences.

Simplicial homotopy equivalence is in general not an equivalence relation. If C is a closed sSet-module with action (c, K) c ⊘ K and coaction (c, K) c^{K}, a simplicial homotopy may also be described by maps c ⊘ Δ^{1} d or c d^{Δ^{1}}. If C has pushouts, the simplicial mapping cylinder factors any map f as follows: Let C_{f} denote the pushout of the diagram

c ⊘ Δ^{1} ^{c⊘i_{1}} c ⊘ Δ^{0} ^{≅} c ^{f} d.

The maps c ⊘ Δ^{1} ^{c⊘s} c ⊘ Δ^{0} ^{≅} c ^{f} d and id_{d} : d d induce the simplicial homotopy equivalence p_{f}: C_{f} d. Its homotopy inverse is the

canonical map. Denote the composition c  c ⊘ Δ^{0} ^{c⊘i_{0}} c ⊘ Δ^{1}  C_{f}

by i_{f}: c C_{f}. Note the factorization p_{f} ∘ i_{f} = f. The relevance of this will become clear in the context of simplicial model categories. In this case, i_{f} is a cofibration provided c is cofibrant. It is easy to prove the next result.

Lemma 2.10. A sSet-functor preserves simplicial homotopies, and therefore simplicial homotopy equivalences.

Corollary 2.11. Assume F: sSet V is a lax monoidal functor. Then any V-functor preserves simplicial homotopy equivalences.

2.5 Spectra as enriched functors

Let (V, ⊗, e) denote a bicomplete closed symmetric monoidal category with initial object ∅. For T ∈ Ob V, one can consider T -spectra in V, see [8, 1.1]. A T -spectrum E is a sequence E_{0}, E_{1}, ⋯ of objects in V, together with structure maps e_{n}: E_{n} ⊗ T E_{n+1} for all n. If E and F are T -spectra, a map of

T -spectra g: E  F is a collection of maps g_{n}: E_{n}  F_{n} such that

[figure: E_{n} ⊗ T; e_{n}; E_{n+1}; F_{n} ⊗ T; g_{n}⊗T; f_{n}; F_{n+1}; g_{n+1}]

commutes for all n. Thus T -spectra in V form a category Sp(V, T), see [8, 1.3]. We claim Sp(V, T) can be viewed as an enriched functor category, cf. [11, 4.3]. Its domain category is the V-category T Sph. The objects in T Sph are the objects T^{n} for n ≥ 0, where T^{0} = e and T^{n} := T ⊗ T^{n−1} for n > 0. The V-objects of morphisms are V_{T Sph}(T^{m}, T^{n}) := T^{n−m} for n ≥ m ≥ 0 and V_{T Sph}(T^{m}, T^{n}) := ∅ for n < m. Note that there are canonical unit maps id_{T^{0}}: T^{0} V_{T Sph}(T^{n}, T^{n}) for all n ≥ 0. It remains to describe the composition. For k, l, m ≥ 0, the map

V_{T Sph}(T^{l+m}, T^{k+l+m}) ⊗ V_{T Sph}(T^{m}, T^{l+m}) V_{T Sph}(T^{m}, T^{k+l+m})

is the associativity isomorphism α_{k,l}: T^{k} ⊗ T^{l} T^{k+l}. In all other cases, the composition is uniquely determined. It follows that T Sph is a V-category, using the associativity and unit coherence laws in V. To elaborate on this definition, let us describe a V-functor π: T Sph V. Define π to be the identity on objects. Concerning morphisms, it suffices to give hom_{T^{m},T^{k+m}}^{π}: V_{T Sph}(T^{m}, T^{k+m}) V(T^{m}, T^{k+m}) for k, m ≥ 0, which we choose as

i_{k,m}: T^{k} ^{η_{T^{m}} T^{k}} V(T^{m}, T^{k} ⊗ T^{m}) ^{V(T^{m},α_{k,m})} V(T^{m}, T^{k+m}).

Associativity coherence and a calculation with adjoints imply that the composition T^{k} ⊗T^{l} ^{α_{k,l}} T^{k+l} ^{i_{k+l,m}} V(T^{m}, T^{k+l+m}) is the same as the composition T^{k} ⊗ T^{l} ^{i_{k,l+m}⊗i_{l,m}} V(T^{l+m}, T^{k+l+m}) ⊗ V(T^{m}, T^{l+m}) ^{comp} V(T^{m}, T^{k+l+m}). Hence π preserves composition, and it clearly preserves identites. In our cases of interest, the maps η_{T^{m}}T^{k} are monomorphisms so that T Sph can be regarded as a sub-V-category.

Proposition 2.12. The categories Sp(V, T) and [T Sph, V] are isomorphic.

Proof. Let X: T Sph V be a V-functor. Define Ψ(X) to be the spectrum with Ψ(X)_{n} := X(T^{n}) and structure maps Ψ(X)_{n} ⊗ T Ψ(X)_{n+1} adjoint to T = V_{T Sph}(T^{n}, T^{n+1}) V(X(T^{n}), X(T^{n+1})). If f: X Y is a Vnatural transformation, let Ψ(f)_{n} := f(T^{n}): X(T^{n}) Y (T^{n}). The diagram

[figure: X(T^{n}) ⊗ T; X(T^{n+1}); Y (T^{n}) ⊗ T; f(T^{n})⊗T; Y (T^{n+1}); f(T^{n+1})]

commutes by V-naturality, and Ψ respects identities and composition. Define Φ: T Sph V by Φ(E)(T^{n}) := E_{n}. If n = m + k, m ≥ 0 and k ≥ 1, hom_{T^{m},T^{n}}^{Φ(E)}: T^{k} = V_{T Sph}(T^{m}, T^{n}) V(E_{m}, E_{n}) is the adjoint of

the composition E_{m} ⊗ T^{k} ^{α_{Em,T,T^{k−1}}^{−1}} (E_{m} ⊗ T) ⊗ T^{k−1} ^{e_{m}⊗T^{k−1}} E_{m+1} ⊗ T^{k−1} ⋯ E_{n}. The maps hom_{T^{n},T^{n}}^{Φ(E)} are determined by the property that Φ(E) has to preserve identities. To prove that Φ(E) is a V-functor, it remains to note that

T^{k} ⊗ T^{l}  \overline{^{α_{k,l}}} T^{k+l}

V(E_{l+m}, E_{k+l+m}) ⊗ V(E_{m}, E_{l+m})

hom_{T l+m,T k+l+m}^{Φ(E)} ⊗hom_{T m,T l+m}^{Φ(E)}

^{comp} V(E_{m}, E_{k+l+m})

^{hom_{T^{m},T^{k+l+m}}^{Φ(E)}}

commutes for m ≥ 0 and k, l ≥ 1. This uses ϵ_{E_{m}}(−⊗E_{m})∘η_{E_{m}}⊗E_{m} = id_{−⊗E_{m}}, associativity coherence, and associativity of composition in V. If g is a map, let Φ(g)(T^{n}) be g_{n}. Then Φ(g) is V-natural, and functoriality of Φ follows. Note that Φ(Ψ(X)) = X on objects, i.e. for all powers of T. The structure maps of X and Φ(Ψ(X)) coincide, since the adjointness isomorphism that defines the structure maps of Φ(Ψ(X)) is inverse to the adjointness isomorphism that defines the structure maps of Ψ(X). The equality Φ(Ψ(f)) = f is obvious, hence Φ ∘ Ψ is the identity functor. Likewise, one finds Ψ ∘ Φ = Id_{Sp(V,T)}.

Remark 2.13. In the following, we will identify Sp(V, T) with [T Sph, V] via 2.12. Then Sp(V, T) is a closed V-module by 2.4. A consequence of A.1 is that the functor “suspension with T ” obtained from the action of V on Sp(V, T) can also be defined as the prolongation [8, 1.5] of the functor −⊗T: V V using the natural transformation t: (− ⊗ T) ∘ (− ⊗ T) (− ⊗ T) ∘ (− ⊗ T), which twists the factors. In detail, the latter is defined by the coherence isomorphism (A⊗T)⊗T ^{α_{A,T,T}} A⊗(T ⊗T) ^{A⊗σ_{T,T}} A⊗(T ⊗T) ^{α_{A,T,T}^{−1}} (A⊗T)⊗T. Another functor “suspension with T ” – which we denote by Σ_{T} – is obtained as the prolongation of −⊗T using the identity natural transformation id_{(−⊗T)∘(−⊗T)}. If X: T Sph V is a V-functor, then the nth structure map in the associated spectrum of X ⊗ T is the left hand side composition in the following diagram.

(X_{n} ⊗ T) ⊗ T

[figure: T ⊗ (X_{n} ⊗ T); α_{T,Xn,T}^{−1}; σ_{Xn⊗T,T}; (T ⊗ X_{n}) ⊗ T; σ_{Xn,T} ⊗T]

[figure: V(X_{n}, X_{n+1}) ⊗ (X_{n} ⊗ T); hom_{n,n+1}^{X}⊗(X_{n}⊗T); α_{V(Xn,X_{n+1}),Xn,T}^{−1}; (V(X_{n}, X_{n+1}) ⊗ X_{n}) ⊗ T; (hom_{n,n+1}^{X}⊗X_{n})⊗T]

[figure: V(X_{n} ⊗ T, X_{n+1} ⊗ T) ⊗ (X_{n} ⊗ T); hom_{Xn,X_{n+1}}^{−⊗T}; ϵ_{Xn⊗T} (X_{n+1}⊗T); X_{n+1} ⊗ T; (ϵ_{Xn} X_{n+1})⊗T]

The right hand side composition is the structure map of the spectrum Σ_{T} X. The lower square commutes by A.1, the middle square commutes by naturality, but the triangle does not commute in general. This will cause some complications in our comparison of stable model categories, cp. Section 6.

The monoidal product ⊗ defines a V-functor mon: V ⊗ V V where

hom_{(A_{1},A_{2})(B_{1},B_{2})}^{mon}: V(A_{1}, B_{1}) ⊗ V(A_{2}, B_{2}) V(A_{1} ⊗ A_{2}, B_{1} ⊗ B_{2})

is the adjoint of the composition

V(A_{1}, B_{1}) ⊗ V(A_{2}, B_{2}) ⊗ A_{1} ⊗ A_{2}

V(A_{1}, B_{1}) ⊗ A_{1} ⊗ V(A_{2}, B_{2}) ⊗ A_{2} ^{V(A_{1},B_{1})⊗ σ_{V(A_{2},B_{2}),A_{1}} ⊗A_{2}}

B_{1} ⊗ B_{2}. ^{ϵ_{A_{1}} B_{1}⊗ ϵ_{A_{2}} B_{2}}

Now suppose that the symmetric monoidal product in V induces a V-functor mon: T Sph ⊗ T Sph T Sph. On objects we have that mon(T^{k}, T^{l}) = T^{k} ⊗ T^{l} = T^{k+l}, while for V-objects of morphisms there is a map f from V_{T Sph}(T^{m}, T^{k}⊗T^{m})⊗V_{T Sph}(T^{n}, T^{l}⊗T^{n}) to V_{T Sph}(T^{m}⊗T^{n}, T^{k}⊗T^{m}⊗T^{l}⊗T^{n}) rendering the following diagram commutative.

[figure: T^{k} ⊗ T^{l}; f; T^{k} ⊗ T^{l}; V(T^{m}, T^{k} ⊗ T^{m}) ⊗ V(T^{n}, T^{l} ⊗ T^{n}); η_{T^{m}} T^{k}⊗η_{T^{n}} T^{l}; hom^{mon}; V(T^{m} ⊗ T^{n}, T^{k} ⊗ T^{m} ⊗ T^{l} ⊗ T^{n}); η_{T^{m}⊗T^{n}} (T^{k}⊗T^{l})] Reverting to adjoints, a straightforward calculation shows that the maps

Documenta Mathematica 8 (2003) 409–488

T^{k+l+m+n} = T^{k}⊗T^{l}⊗T^{m}⊗T^{n} [figure: f⊗T^{m}⊗T^{n}; T^{k}⊗σ_{T^{l},T^{m}} ⊗T^{n}] T^{k}⊗T^{m}⊗T^{l}⊗T^{n} = T^{k+l+n+m}

must coincide. This is only possible if σ_{T,T} = id_{T^{2}} and f = id. In other words, the product ⊗ does not necessarily restrict to a monoidal product of T Sph via π: T Sph V. In the next section, we show – following [11, 5.15] – how to remedy this by enlarging T Sph.

2.6 Symmetric spectra as enriched functors

For ease of notation, we will leave out associativity and unit isomorphisms throughout this section. If n ≥ 1, let \overline{n} be short for \{1, ⋯, n\} and let \overline{0} denote the empty set. Let Inj be the category with objects the sets \overline{n} for all n ≥ 0, and injective maps as morphisms. If m ≤ n, define

Inj(m, n) := ∐_{Set_{Inj}(\overline{m},\overline{n})} T^{0}

where T^{0} is the unit of V. Note that Inj(n, n) is a group object in V, the symmetric group on n letters. By regarding Inj(n, n) as a V-category with a single object, a left Inj(n, n)-action on A ∈ Ob V is a V-functor Inj(n, n) V with value A. One gets a left Inj(n, n)-action on T^{n} = T_{n} ⊗ ⋯ ⊗ T_{2} ⊗ T_{1} using iterations of the commutativity isomorphism σ_{T,T}. A symmetric T -spectrum X in V as defined in [8, 7.2] consists of a sequence X_{0}, X_{1}, ⋯, X_{n}, ⋯, where X_{n} is an object of V with a left Inj(n, n)-action, and with structure maps X_{n} ⊗ T X_{n+1} such that the following composition X_{n} ⊗ T^{m} X_{n+1} ⊗ T^{m−1} ⋯ X_{n+m} is Inj(n, n) ⊗ Inj(m, m)- equivariant. A map of symmetric T -spectra consists of maps X_{n} Y_{n} that are compatible with the Inj(n, n)-action and the structure maps. We will show that the category Sp^{Σ}(V, T) of symmetric T -spectra in V is isomorphic to a category of V-functors with codomain V and domain T Sph^{Σ}. The objects in T Sph^{Σ} are the objects in T Sph, but the morphism objects are different. If n = k + m with m ≥ 0, k ≥ 0, define the V-object V_{T Sph^{Σ}}(T^{m}, T^{n}) to be Inj(m, n) ⊗ T^{k}. If n < m, define V_{T Sph^{Σ}}(T^{m}, T^{n}) to be the initial object. The unit map T^{0} Inj(n, n) = V_{T Sph^{Σ}}(T^{n}, T^{n}) is the canonical map to the summand corresponding to id\overline{n}. Next, to describe the composition, identify Inj(m, k + m) ⊗ T^{k} indexed by β: \overline{m} \overline{k + m} with T_{i_{1}^{β}} ⊗ ⋯ ⊗ T_{i_{k}^{β}} = T^{k}.

Here \{i_{1}^{β}, i_{2}^{β}, ⋯, i_{k}^{β}\} is the reordering of \overline{k + m} \ β(\overline{m}) which satisfies that i_{1}^{β} > i_{2}^{β} > ⋯ > i_{k}^{β}. If n = k + l + m with k, l, m ≥ 0, we define the map

Inj(l + m, n) ⊗ T^{k} ⊗ Inj(m, l + m) ⊗ T^{l} Inj(m, n) ⊗ T^{k+l}

in two steps. First, we identify the source of the map with the coproduct ∐_{Set_{Inj}(\overline{l+m},\overline{n})×Set_{Inj}(\overline{m},\overline{l+m})} T^{k+l}. For the second step, consider the unique isomorphism T^{k+l} T^{k+l} induced by the permutation that reorders the set \{i_{1}^{γ}, i_{2}^{γ}, ⋯, i_{k}^{γ}, γ(i_{1}^{β}), γ(i_{2}^{β}), ⋯, γ(i_{l}^{β})\}, i.e. the set \overline{k + l + m} \ (γ ∘ β)(\overline{m}), as {i_{1}^{γ∘β}, i_{2}^{γ∘β}, ⋯, i_{k+l}^{γ∘β}}. This isomorphism maps the summand T_{i_{1}^{γ}} ⊗ T_{i_{2}^{γ}} ⊗ ⋯ ⊗ T_{i_{k}^{γ}} ⊗ T_{i_{1}^{β}} ⊗ T_{i_{2}^{β}} ⊗ ⋯ ⊗ T_{i_{l}^{β}} = T^{k+l} indexed by (γ, β) to the summand T_{i_{1}^{γ∘β}} ⊗ T_{i_{2}^{γ∘β}} ⊗ ⋯ ⊗ T_{i_{k+l}^{γ∘β}} = T^{k+l} indexed by γ ∘ β.

Lemma 2.14. T Sph^{Σ} is a symmetric monoidal V-category.

Proof. The composition in T Sph^{Σ} is clearly unital. Associativity follows, since the permutation reordering the set {i_{1}^{γ}, i_{2}^{γ}, …, i_{k}^{γ}, γ(i_{1}^{β}), γ(i_{2}^{β}), …, γ(i_{l}^{β})} is unique. On objects, (T^{l}, T^{k}) T^{k+l} defines the monoidal product in T Sph^{Σ}. For morphisms, consider injections β: \overline{n} \overline{l + n} and γ: \overline{m} \overline{k + m}. Define β ▷ γ: \overline{m + n} \overline{k + l + m + n} by concatenation. That is, β ▷ γ(i) is defined as γ(i) for i ≤ m and as β(i − m) + k + m for i > m. Note that the ordered set i_{1}^{β} + k + m > ⋯ > i_{l}^{β} + k + m > i_{1}^{γ} > ⋯ > i_{k}^{γ} coincides with i_{1}^{β▷γ} > ⋯ > i_{k+l}^{β▷γ}. On V-objects of morphisms, the map

Inj(n, l + n) ⊗ T^{l} ⊗ Inj(m, k + m) ⊗ T^{k} Inj(m + n, k + l + m + n) ⊗ T^{k+l}

sends the summand T^{l} ⊗ T^{k} = T_{i_{1}^{β}} ⊗ ⋯ ⊗ T_{i_{l}^{β}} ⊗ T_{i_{1}^{γ}} ⊗ ⋯ ⊗ T_{i_{k}^{γ}} indexed by

(\overline{n} ^{β} \overline{l + n}, \overline{m} ^{γ} \overline{k + m}) via the identity onto the summand T^{k+l} indexed by β ▷ γ. To see this, one can rewrite the indices according to the above equality of ordered sets, that is, changing T_{i_{1}^{β}} ⊗ ⋯ ⊗ T_{i_{l}^{β}} ⊗ T_{i_{1}^{γ}} ⊗ ⋯ ⊗ T_{i_{k}^{γ}} into T_{i_{1}^{β}+k+m} ⊗ ⋯ ⊗ T_{i_{l}^{β}+k+m} ⊗ T_{i_{1}^{γ}} ⊗ ⋯ ⊗ T_{i_{k}^{γ}}.

Let us check that this defines a V-functor mon^{Σ}: T Sph^{Σ} ⊗T Sph^{Σ} T Sph^{Σ}. The reason is essentially that concatenation ▷: Inj × Inj Inj is a functor. Since id\overline{n} ▷ id\overline{m} = id\overline{m+n}, if follows that mon^{Σ} preserves identities. To check compatibility with composition, fix four injective maps α: \overline{m} \overline{l + m}, β: \overline{l + m} \overline{k + l + m =: n}, γ: \overline{r} \overline{q + r} and δ: \overline{q + r} \overline{p + q + r}. The equality (δ ∘ γ) ▷ (β ∘ α) = (δ ▷ β) ∘ (γ ▷ α) implies that it suffices to consider only one summand, say M = {i_{1}^{δ}, ⋯, i_{p}^{δ}, i_{1}^{β}, ⋯, i_{k}^{β}, i_{1}^{γ}, ⋯, i_{q}^{γ}, i_{1}^{α}, ⋯, i_{l}^{α}}. First, by rewriting M we find {i_{1}^{δ}, ⋯, i_{p}^{δ}, δ(i_{1}^{γ}), ⋯, δ(i_{q}^{γ}), i_{1}^{β}, ⋯, i_{k}^{β}, β(i_{1}^{α}), ⋯, β(i_{l}^{α})}, and next by reordering we get {i_{1}^{δ∘γ}, ⋯, i_{p+q}^{δ∘γ}, i_{1}^{β∘α}, ⋯, i_{k+l}^{β∘α}}. The monoidal product rewrites the latter as M′ = {i_{1}^{δ∘γ} + n, ⋯, i_{p+q}^{δ∘γ} + n, i_{1}^{β∘α}, ⋯, i_{k+l}^{β∘α}}

and there is a bijection M ^{≅} M′. On the other hand, the monoidal product rewrites M as {i_{1}^{δ} +n, ⋯, i_{p}^{δ} +n, i_{1}^{β}, ⋯, i_{k}^{β}, i_{1}^{γ} +l+m, ⋯, i_{q}^{γ} +l+m, i_{1}^{α}, ⋯, i_{l}^{α}}, and composition rewrites this set as {i_{1}^{δ} + n, ⋯, i_{p}^{δ} + n, i_{1}^{β}, ⋯, i_{k}^{β}, δ ▷ β(i_{1}^{γ} + l + m), ⋯, δ ▷ β(i_{q}^{γ} + l + m), δ ▷ β(i_{1}^{α}), ⋯, δ ▷ β(i_{l}^{α}).} By definition of ▷, this set coincides with

{i_{1}^{δ} + n, ⋯, i_{p}^{δ} + n, i_{1}^{β}, ⋯, i_{k}^{β}, δ(i_{1}^{γ}) + n, ⋯, δ(i_{q}^{γ}) + n, β(i_{1}^{α}), ⋯, β(i_{l}^{α})}

which is then reordered as M′ = {i_{1}^{(δ▷β)∘(γ▷α)}, ⋯, i_{k+l+p+q}^{(δ▷β)∘(γ▷α)}}. The corre-

sponding bijection M ^{≅} M′ is thus the same as above. Hence mon^{Σ} is a V-functor. By definition, mon^{Σ}(T^{0}, T^{m}) = T^{m} = mon^{Σ}(T^{m}, T^{0}). So T^{0} is a strict unit in T Sph^{Σ}. Similarly, strict associativity holds. The commutativity isomorphism T^{0} V_{T Sph^{Σ}}(T^{m+n}, T^{m+n}) is the canonical map on the summand indexed by the permutation \overline{m + n} \overline{n + m} which interchanges m and n. Next, the coherence conditions [2, 6.3,6.4,6.5] follow by a straightforward calculation with permutations of \overline{l + m + n}. This ends the proof.

To explain the composition in T Sph^{Σ} more throughly, we will define V-functors ν: T Sph T Sph^{Σ} and σ: T Sph^{Σ} V such that σ ∘ ν = π: T Sph V. Both ν and σ are the respectively identities on objects. The map

hom_{T^{m},T^{l+m}}^{ν}: T^{l}  Inj(m, l + m) ⊗ T^{l} ≅ ∐_{Set_{Inj}(\overline{m},\overline{l+m})} T^{l}

hits the summand indexed by the inclusion \overline{m} ^{⊂} \overline{l + m}. It is then clear that ν preserves identities. For the composition, consider

T^{k} ⊗ T^{l}  \overline{^{id}} T^{k+l}

[figure: Inj(l + m, k + l + m) ⊗ T^{k} ⊗ Inj(m, l + m) ⊗ T^{l}; hom_{T^{l+m},T^{k+l+m}}^{ν} ⊗hom_{T^{m},T^{l+m}}^{ν}; comp; Inj(m, k + l + m) ⊗ T^{k+l}; hom_{T^{m},T^{k+l+m}}^{ν}]

and observe that the left vertical map hits the summand

(T_{k+l+m} ⊗ T_{k+l+m−1} ⊗ ⋯ ⊗ T_{l+m+1}) ⊗ (T_{l+m} ⊗ T_{l+m−1} ⊗ ⋯ ⊗ T_{m+1})

indexed by the inclusions (\overline{l + m} ^{⊂}  \overline{k + l + m}, \overline{m} ^{⊂}  \overline{l + m}). Composition

in T Sph^{Σ} maps this summand by the identity to the summand T_{k+l+m} ⊗ ⋯ ⊗ T_{m+1} indexed by the inclusion \overline{m} ^{⊂} \overline{l + m} ^{⊂} \overline{k + l + m}, since the indices are ordered in the prescribed way. So the diagram commutes and ν is a V-functor. This also explains the ordering i_{1}^{β} > i_{2}^{β} > ⋯ > i_{k}^{β} of the set \overline{k + m} \ β(\overline{k}). The adjoint Inj(m, l + m) ⊗ T^{l} ⊗ T^{m} ≅ ∐_{Set_{Inj}(\overline{m},\overline{l+m})} T^{l} ⊗ T^{m} T^{l+m} of hom_{T^{m},T^{l+m}}^{σ}: Inj(m, l + m) ⊗ T^{l} V(T^{m}, T^{l+m}) is defined as follows: Consider the summand T_{i_{1}^{β}} ⊗ ⋯ ⊗ T_{i_{l}^{β}} ⊗ T_{m} ⊗ ⋯ ⊗ T_{1} indexed by β. First rewrite the indices as T_{i_{1}^{β}} ⊗ ⋯ ⊗ T_{i_{l}^{β}} ⊗ T_{β(m)} ⊗ ⋯ ⊗ T_{β(1)}, and then map to T_{l+m} ⊗ ⋯ ⊗ T_{2} ⊗ T_{1} by the unique permutation which reorders the set {i_{1}^{β}, ⋯, i_{l}^{β}, β(m), ⋯, β(1)}. To conclude that σ is a V-functor, one has to check that reordering the set {i_{1}^{γ}, ⋯, i_{k}^{γ}, γ(i_{1}^{β}), ⋯, γ(i_{l}^{β}), γ(β(m)), ⋯, γ(β(1))} is the same as first reordering {i_{1}^{γ}, ⋯, i_{l}^{γ}, i_{1}^{β}, ⋯, i_{k}^{β}, β(m), ⋯, β(1)} as {i_{1}^{γ}, ⋯, i_{l}^{γ}, l + m, ⋯, 2, 1}, and then reordering {i_{1}^{γ}, ⋯, i_{l}^{γ}, γ(l+m), ⋯, γ(2), γ(1)}. However, the monoidal product mon^{Σ} on T Sph^{Σ} is such that the diagram

T Sph^{Σ} ⊗ T Sph^{Σ} \overline{^{σ⊗σ}} V ⊗ V

[figure: T Sph^{Σ}; mon^{Σ}; σ; V; mon]

commutes, so σ is in fact a strict symmetric monoidal V-functor.

Proposition 2.15. The category [T Sph^{Σ}, V] is isomorphic to the category of symmetric T -spectra in V, and ν induces the forgetful functor from symmetric T -spectra to T -spectra. The smash product on [T Sph^{Σ}, V] corresponds to the smash product on Sp^{Σ}(V, T).

Proof. This is similar to 2.12, and some details will be left out in the proof. The functor Φ: [T Sph^{Σ}, V] Sp^{Σ}(V, T) maps X: T Sph^{Σ} V to the sequence XT^{0}, XT^{1}, · · ·. It is clear that XT^{n} has a left Inj(n, n)-action. The adjoint of the structure map XT^{n} ⊗T XT^{n+1} is given by hom_{T^{n},T^{n+1}}^{X} ∘hom_{T^{n},T^{n+1}}^{ν}. More generally, the composition of structure maps XT^{n} ⊗ T^{k} XT^{n+k} is

the adjoint of T^{k} ^{hom_{T^{n},T^{k+n}}^{ν}} Inj(n, k + n) ⊗ T^{k} ^{hom_{T^{n},T^{k+n}}^{X}} V(XT^{n}, XT^{k+n}). This proves the required equivariance. The definition of Φ on V-natural transformations is clear, and also functoriality.

The inverse Ψ: Sp^{Σ}(V, T) [T Sph^{Σ}, V] is harder to define. If X_{0}, X_{1}, ⋯

is a symmetric T -spectrum with structure maps x_{n}^{1}: X_{n} ⊗ T X_{n+1}, define Ψ(X): T Sph^{Σ} V on objects by T^{n} X_{n}. Since X_{n} has a left Inj(n, n)-action, there is the map hom_{T^{n},T^{n}}^{Ψ(X)}: V_{T Sph^{Σ}}(T^{n}, T^{n}) =

Inj(n, n)  V(X_{n}, X_{n}). Choose an injection β: \overline{n}  \overline{k + n}, and define

hom_{T^{n},T^{k+n}}^{Ψ(X)}: V_{T Sph^{Σ}}(T^{n}, T^{k+n}) = Inj(n, k + n) V(X_{n}, X_{k+n})

on the summand T_{i_{1}^{β}} ⊗⋯⊗T_{i_{k}^{β}} as the adjoint map of the following composition

T_{i_{1}^{β}} ⊗ ⋯ ⊗ T_{i_{k}^{β}} ⊗ X_{n} ^{σ_{T^{k},Xn}} X_{n} ⊗ T_{i_{1}^{β}} ⊗ ⋯ ⊗ T_{i_{k}^{β}}  ^{x_{n}^{k}} X_{k+n}  X_{k+n} where

x_{n}^{k} is defined by the structure maps. The right hand map is the isomorphism associated to the permutation of the set \overline{k + n} that changes \{k + n, ⋯, 2, 1\} to {i_{1}^{β}, ⋯, i_{k}^{β}, β(n), ⋯, β(1)} and reorders this as {k + n, ⋯, 2, 1}. For the latter we use the left Inj(k + n, k + n)-action on X_{k+n}. The Inj(n, n)-action on X_{n} is unital, so Ψ(X) preserves identities. To prove

that Ψ(X): T Sph^{Σ}  V is a V-functor, pick injective maps β: \overline{m}  \overline{l + m},

γ: \overline{l + m} \overline{k + l + m} and consider the following diagram.

T_{i_{1}^{γ}} ⊗ ⋯ ⊗ T_{i_{k}^{γ}} ⊗ T_{i_{1}^{β}} ⊗ ⋯ ⊗ T_{i_{l}^{β}} ⊗ X_{m} \overline{^{T^{k}⊗(x_{m}^{l}∘σ_{T^{l},Xm})}} T_{i_{1}^{γ}} ⊗ ⋯ ⊗ T_{i_{k}^{γ}} ⊗ X_{l+m}

[figure: T_{i_{1}^{γ∘β}} ⊗ ⋯ ⊗ T_{i_{k+l}^{γ∘β}} ⊗ X_{m}; f⊗X_{m}; X_{k+l+m}; x_{l+m}^{k}∘σ_{T^{k},X_{l+m}}; x_{m}^{k+l}∘σ_{T^{k+l},Xm}; T_{i_{1}^{γ}} ⊗ ⋯ ⊗ T_{i_{k}^{γ}} ⊗ X_{l+m}; T^{k}⊗b]

[figure: X_{k+l+m}; x_{m}^{k+l}∘σ_{T^{k+l},Xm}; X_{k+l+m}; f′; X_{k+l+m}; x_{l+m}^{k}∘σ_{T^{k},X_{l+m}}]

[figure: X_{k+l+m}; cb; c; cb]

Here b: X_{l+m} X_{l+m} is the isomorphism obtained from reordering the set {i_{1}^{β}, ⋯, i_{k}^{β}, β(m), ⋯, β(1)} as {l + m, ⋯, 2, 1}, and similarly for c and cb. The isomorphism f: T^{k+l} T^{k+l} is induced by the permutation

{i_{1}^{γ}, ⋯, i_{k}^{γ}, γ(i_{1}^{β}), ⋯, γ(i_{l}^{β})} ^{≅} {i_{1}^{γ∘β}, ⋯, i_{k+l}^{γ∘β}}.

Similarly, f′: X_{k+l+m} X_{k+l+m} is induced by the permutation

{i_{1}^{γ}, ⋯, i_{k}^{γ}, γ(i_{1}^{β}), ⋯, γ(i_{l}^{β}), m, ⋯, 1} ^{≅} {i_{1}^{γ∘β}, ⋯, i_{k+l}^{γ∘β}, m, ⋯, 1}.

Since x_{m}^{k+l} is Inj(k + l, k + l) ⊗ Inj(m, m)-equivariant, the left parallelogram commutes. The upper triangle commutes by commutativity coherence. Finally, the right parallelogram commutes, because x_{l+m}^{k} is Inj(k, k)⊗Inj(l +m, l +m)- equivariant, the Inj(k+l+m, k+l+m)-action on X_{k+l+m} is associative and the permutation obtained from reordering {i_{1}^{γ}, ⋯, i_{k}^{γ}, i_{1}^{β}, ⋯, i_{l}^{β}, β(m), ⋯, β(1)} as {i_{1}^{γ}, ⋯, i_{k}^{γ}, γ(l + m), ⋯, γ(1)} and then as {k + l + m, ⋯, 2, 1} equals the permutation obtained from {i_{1}^{γ}, ⋯, i_{k}^{γ}, γ(i_{1}^{β}), ⋯, γ(i_{l}^{β}), γ(β(m)), ⋯, γ(β(1))}. The definition of Ψ on maps is clear and functoriality follows easily. The proof of 2.12 shows that Φ ∘ Ψ = Id_{Sp^{Σ}(V,T)}, but an extra argument is required to prove the equality Ψ ∘ Φ = Id_{[T Sph^{Σ},V]}. The only point which is not obvious is whether the maps hom_{T^{m},T^{k+m}}^{X}: Inj(m, k+m)⊗T^{k} V(XT^{m}, XT^{m+k}) and hom_{T^{m},T^{k+m}}^{Ψ(Φ(X))}: Inj(m, k + m) ⊗ T^{k} V(XT^{m}, XT^{m+k}) coincide. To prove this, we fix an injection β: \overline{m} \overline{k + m} and let ι: \overline{m} ^{⊂} \overline{k + m} denote the inclusion. The permutation γ: \overline{k + m} \overline{k + m} obtained from rewriting the set {k + m, ⋯, 1 + m, m, ⋯, 1} as {i_{1}^{β}, ⋯, i_{k}^{β}, β(m), ⋯, β(1)} and reordering this set as {k + m, ⋯, 2, 1} has the property that γ ∘ ι = β. Since X is a V-functor, the map hom_{T^{m},T^{k+m}}^{X} is determined by its restriction to the summand indexed by \overline{m} ^{⊂} \overline{k + m}. This shows that hom_{T^{m},T^{k+m}}^{X} = hom_{T^{m},T^{k+m}}^{Ψ(Φ(X))}.

The claim concerning [ν, V]: [T Sph^{Σ}, V] [T Sph, V] is clear by the above. It remains to prove compatibility of the smash products. The smash product ∧′ of symmetric T -spectra satisfies F_{m}T^{0} ∧′ F_{n}T^{0} ≅ F_{m+n}T^{0} by the remark below [8, 7.3]. On the other hand, the monoidal product in [T Sph^{Σ}, V] is determined by representable functors, and from 2.7 there is the natural isomorphism

V_{T Sph^{Σ}}(T^{m}, −) ∧ V_{T Sph^{Σ}}(T^{n}, −) ≅ V_{T Sph^{Σ}}(T^{m+n}, −).

This completes the proof.

Corollary 2.16. Let C be a full sub-V-category. Assume C is closed under the monoidal product, and contains the unit and T. Then σ: T Sph^{Σ} V factors over C ^{⊂} V, and the induced V-functor T Sph^{Σ} C induces a lax symmetric monoidal functor [C, V] Sp^{Σ}(V, T) which has a strict symmetric monoidal left adjoint.

Proof. The factorization of σ is clear. Under the isomorphism in 2.15, the closed symmetric monoidal product on [T Sph^{Σ}, V] coincides with the closed symmetric monoidal product on symmetric T -spectra. For formal reasons, the V-functor induced by the first factor of σ is lax symmetric monoidal. By checking on representable functors, it follows that its left V-adjoint obtained by an enriched Kan extension 2.5 is strict symmetric monoidal.

3 Model Categories

The term model category is to be understood in the sense of [7, 1.1.4]. We denote weak equivalences by ^{∼} , fibrations by [figure] and cofibrations by .

3.1 Types of model categories

The model structures we will consider on enriched functor categories require a cofibrantly generated model category (C, I, J) as input. For a definition of this type of model categories and related terminology, consider [7, 2.1]. Maps in I = {i: si ti}_{i∈I} are called generating cofibrations, and maps in J = {j: sj ^{∼} tj}_{j∈J} are called generating acyclic cofibrations. The (co)domains of I and J may have additional properties.

Definition 3.1. An object A ∈ Ob C is finitely presentable if the set-valued Hom-functor Set_{C}(A, −) commutes with all filtered colimits. If C is a Vcategory, A ∈ Ob C is V-finitely presentable if the V-valued Hom-functor V_{C}(A, −) commutes with all filtered colimits.

A set is finitely presentable in the category of sets if and only if it is a finite set. If the unit e in V is finitely presentable, any V-finitely presentable object of V is finitely presentable. See [8, 4.1] for 3.2 and 3.3.

Definition 3.2. A cofibrantly generated model category C is finitely generated if I and J can be chosen such that the (co)domains of the maps in I and J are κ-small for a finite cardinal κ.

Finitely generated model categories are not necessarily closed under Bousfield localization, cf. [8, §4]. The following definition was suggested by Voevodsky.

Definition 3.3. The cofibrantly generated model category C is almost finitely generated if I and J can be chosen such that the (co)domains of the maps in I are κ-small for a finite cardinal κ, and there exists a subset J′ of J for which

• the domains and the codomains of the maps in J′ are κ-small for a finite cardinal κ,

• a map f: A B in C such that B is fibrant is a fibration if and only if it is contained in J′-inj.

The left Bousfield localization with respect to a set with sSet-small domains and codomains preserves the structure of almost finitely generated cellular left proper simplicial model categories [6, Chapters 3,4]. For almost finitely generated model categories, the classes of weak equivalences and fibrant objects are closed under sequential colimits. We require these classes to be closed under filtered colimits, which holds for model categories of the following type.

Definition 3.4. A cofibrantly generated model category V is weakly finitely generated if I and J can be chosen such that the following conditions hold.

• The domains and the codomains of the maps in I are finitely presentable.

• The domains of the maps in J are small.

• There exists a subset J′ of J of maps with finitely presentable domains and codomains, such that a map f: A B in V with fibrant codomain B is a fibration if and only if it is contained in J′-inj.

The choices of the sets I, J and J′ will often be left implicit in the following. A weakly finitely generated model category is almost finitely generated. Examples include simplicial sets, simplicial sets with an action of a finite group, cp. 9.5, and the category of pointed simplicial presheaves on the smooth Nisnevich site. The latter will be discussed in [5].

Lemma 3.5. Assume V is a weakly finitely generated model category. Then the following classes are closed under filtered colimits: weak equivalences, acyclic fibrations, fibrations with fibrant codomain, and fibrant objects.

Proof. Since V is cofibrantly generated, [6, 11.6.1] shows that Fun(I, V) supports a cofibrantly generated model structure for any small category I. Fibrations and weak equivalences are defined pointwise. Any weak equivalence in Fun(I, V) factors as an acyclic cofibration g composed with an acyclic fibration p: T [figure] B. Consider the induced factorization colimp ∘ colimg. Note that colimg is an acyclic cofibration, since colim is a left Quillen functor. Therefore, the second claim will imply the first claim. If I is filtered, then colimp is an acyclic fibration: Let

si \overline{^{α}} colimT

[figure: ti; i; β; colimB; colimp]

be a lifting problem, where i ∈ I. The existence of a lift in this diagram for all choices of α and β is equivalent to surjectivity of the canonical map

ϕ: Set_{V}(ti, colimT) Set_{V}(si, colimT) ×_{Set_{V} (si,colimB)} Set_{V}(ti, colimB).

Since si and ti are finitely presentable by assumption and filtered colimits commute with pullbacks in Set, the canonical map ϕ is the filtered colimit of the canonical maps

ϕ_{d}: Set_{V}(ti, T (d)) Set_{V}(si, T (d)) ×_{Set_{V} (si,B(d))} Set_{V}(ti, B(d))

induced by composition with p(d) and pre-composition with i. Note that ϕ_{d} is surjective since i is a cofibration and p is a pointwise acyclic fibration. It follows that ϕ is surjective and colimp is an acyclic fibration. The proof of the third claim – including the last claim as a special case – is analogous.

Definition 3.6. Let (V, ⊗, e) be a closed symmetric monoidal category, and C a right V-module with action (C, A) C ⊘ A. Consider, for f: C D a map in C and g: A B a map in V, the diagram:

[figure: C ⊘ A; C⊘g; C ⊘ B; D ⊘ A; f⊘A; D⊘f; D ⊘ B; f⊘B]

If C has pushouts, denote the induced map from the pushout of the diagram to the terminal corner by f□g: D ⊘ A ∪_{C⊘A} C ⊘ B D ⊘ B. This is the pushout product map of f and g.

Note that 3.6 applies to (V, ⊗, e) considered as a right V-module. Recall the pushout product axiom [15, 3.1].

Definition 3.7. Let (V, ⊗, e) be a closed symmetric monoidal category and a model category. It is a monoidal model category if the pushout product f□g of two cofibrations f and g is a cofibration, which is acyclic if either f or g is acyclic.

Monoidal model categories with a cofibrant unit are symmetric monoidal model categories in the sense of [7, 4.2.6]. See [15, 3.3] for the following definition.

Definition 3.8. Let V be a monoidal model category. For a class K of maps in V, define K ⊗ V as the class of maps f ⊗ A, where f is a map in K and A ∈ Ob V. Let aCof(V) be the class of acyclic cofibrations in V. The monoid axiom holds if every map in aCof(V) ⊗ V-cell is a weak equivalence.

In the proof of 4.2, we will use the monoid axiom to construct the pointwise model structure on enriched functor categories. Let us end this section with the definition of two other types of model categories.

Definition 3.9. Let V be a monoidal model category, C a closed V-module and a model category. The action of V on C allows us to consider the pushout product of a map in C and a map in V. Then C is a V-model category if the pushout product of a cofibration f in C and a cofibration g in V is a cofibration in C, which is acyclic if either f or g is acyclic.

Note that a simplicial model category is a sSet-model category.

Lemma 3.10. A simplicial homotopy equivalence in a sSet-model category is a weak equivalence.

Proof. This follows from [6, 9.5.16]. Simplicial homotopy equivalences in 2.9 are also simplicial homotopy equivalences as defined in [6, 9.5.8].

Definition 3.11. Let F: V W be a strict symmetric monoidal functor of monoidal model categories. If F is a left Quillen functor, W is called a monoidal V-model category.

A monoidal V-model category is clearly a V-model category.

3.2 Homotopy pullback squares

Homotopy pullback squares will be used to characterize fibrations. Definition 3.12 is equivalent to [7, 7.1.12].

Definition 3.12. Let C be a model category. A commutative diagram

[figure: A; B; C; g; D; f]

in C is a homotopy pullback square if for any commutative diagram

[figure: C; g; D; f; B; C′; ∼; D′; ∼; B′; ∼]

where C′ and D′ are fibrant and B′ [figure] D′ is a fibration, the canonical map A C′ ×_{D′} B′ is a weak equivalence.

This definition seems to by asymmetric, but one of the squares

[figure: A; f; B; A; g; C]

or

[figure: C; g; i; D; h; B; f; h; D; i]

is a homotopy pullback square if and only if the other is a homotopy pullback square. We list some elementary properties.

Lemma 3.13. All diagrams below are commutative diagrams in C.

1. The diagram

[figure: A; f; B; C; ∼; D]

is a homotopy pullback square if and only if f is a weak equivalence.

2. Consider a natural transformation f: A B of diagrams

[figure: A_{0}; A_{1}; B_{0}; B_{1}]

and

[figure: A_{2}; A_{3}; B_{2}; B_{3}]

which is a pointwise weak equivalence. That is, f_{i}: A_{i} B_{i} is a weak equivalence for all i ∈ {0, 1, 2, 3}. Then A is a homotopy pullback square if and only if B is a homotopy pullback square.

3. Let

[figure: A_{0}; A_{1}; A_{2}]
(1)  (2)
[figure: B_{0}; B_{1}; B_{2}]

be a diagram where (2) is a homotopy pullback square. Then the composed square (12) is a homotopy pullback square if and only (1) is a homotopy pullback square.

Lemma 3.14. Assume C is right proper. Then

[figure: A; B; C; D; f]

is a homotopy pullback square if and only if for some factorization of f as a weak equivalence B ^{∼} E followed by a fibration E [figure] D, the induced map A C ×_{D} E is a weak equivalence.

Proof. Since C is right proper, the dual of the gluing lemma holds. The statement follows easily.

3.3 Fibrant replacement functors

In every model category, any object may be replaced by a fibrant object in a functorial way up to an acyclic cofibration. It is often desirable to explicate fibrant replacement functors. Quillen’s small object argument is the classical method 3.3.1. We place emphasis on enriched fibrant replacement functors 3.3.2. Another fibrant replacement functor is constructed in 3.3.3 as a certain filtered colimit.

3.3.1 Classical

Fix a cocomplete category V and a set K = {sk tk}_{k∈K} of maps in V with finitely presentable domains and finitely presentable codomains. The set K gives rise to a natural transformation of endofunctors on V, namely

∐_{k∈K} ∐_{f∈Set_{V} (sk,−)} sk ^{∐_{k∈K} ∐_{f∈Set_{V} (sk,−)} k} ∐_{k∈K} ∐_{f∈Set_{V} (sk,−)} tk.

Consider also

∐_{k∈K} ∐_{f∈Set_{V} (sk,−)} sk ^{∐_{k∈K} ∐_{f∈Set_{V} (sk,−)} f} ∐_{k∈K} ∐_{f∈Set_{V} (sk,−)} Id_{V} ^{codiagonal} Id_{V},

and take the pushout of the two natural transformations. Let ι_{1}^{K}: Id_{V} F_{1}^{K} denote the canonically induced map. This construction can be iterated. Suppose there is a natural transformation ι_{n}^{K}: F_{n−1}^{K} F_{n}^{K} of endofunctors of V. Next, define F_{n+1}^{K} as the pushout of

∐_{k∈K} ∐_{f∈Set_{V} (sk,F_{n}(−))} tk  ∐_{k∈K} ∐_{f∈Set_{V} (sk,F_{n}(−))} sk  F_{n}.

The colimit of the sequence Id_{V} = F_{0} ^{ι_{1}^{K}} F_{1}^{K} ^{ι_{2}^{K}} ⋯ is denoted as F^{K}: V V, and ι^{K}: Id_{V} F^{K} is the canonical natural transformation. The following statement is a special case of [7, 2.1.14].

Lemma 3.15. For every object A in V, the map ι^{K}(A): A F^{K}(A) is in K-cell, and the map F^{K}(A) ∗ is in K-inj.

Corollary 3.16. Suppose that V is a weakly finitely generated model category. Then ι^{J′}: Id_{V} F^{J′} is a fibrant replacement functor, i.e. F^{J′}(A) is fibrant and the natural map ι^{J′}(A): A F^{J′}(A) is an acyclic cofibration for all A ∈ Ob V.

This fibrant replacement functor yields “big” objects. We will make this more precise after recalling some definitions from [6, Chapter 10].

Definition 3.17. A relative K-cell complex f: A B is called finite if f is a composition A = B_{0} ^{f_{1}} B_{1} ^{f_{2}} ⋯ ^{f_{n}} B_{n} = B where each f_{m} is a cobase change of a map in K, i.e. f is obtained by attaching finitely many cells from K.

Definition 3.18. A relative K-cell complex f: A B is called presented if there is an explicit choice of the data [6, 10.6.3]. In detail, one chooses a limit ordinal λ, a λ-sequence A = B_{0} ^{f_{1}} B_{1} ⋯ ^{f_{β+1}} B_{β+1} ⋯ whose sequential composition is f, and f_{β+1}: B_{β} B_{β+1} is gotten from the pushout of the following diagram for every β < λ:

[figure: ∐_{m∈M_{β}} ti_{m}; ∐ i_{m}; ∐_{m∈M_{β}} si_{m}; B_{β}]

We omit the choice from the notation. Let f: A B be a presented relative K-cell complex. A subcomplex of f is a presented K-cell complex g: A C relative to A such that the explicit choice relevant for g is a subset of the explicit choice for f, see [6, 10.6.7]. In particular, there exists a map h: C B in K-cell such that h ∘ g = f. A subcomplex is called finite if it is a finite K-cell complex relative to A, using the explicit choice.

Any relative K-cell complex can be turned into a presented one. If we consider ι^{K}(A) as a presented relative K-cell complex, we will use the explicit choices appearing in its construction.

Lemma 3.19. Let f: A B be a finite relative K-cell complex. Then f has the structure of a finite subcomplex of ι^{K}(K): A F^{K}(A).

Proof. We will prove this by induction on the number of cells. By convention, a map obtained by attaching no cells is an identity map. Suppose the following is true. If f: A B is obtained by attaching n cells from K, with n ≥ 0, then f has the structure of a finite subcomplex of the presented relative K-cell complex

A ^{ι_{1}^{K}(A)} ⋯ ^{ι_{n}^{K}(A)} F_{n}^{K}(A). Assume f: A B is obtained by attaching n+1 cells from K. By definition, there is a factorization A ^{g} C ^{h} B where g is obtained by attaching n cells from K and h is the cobase change of some k ∈ K along a map α: sk C. From the induction hypothesis, α induces a

map sk  C  F_{n}A, hence an element in Set_{V}(sk, F_{n}A). It follows that

there is a map β: tk F_{n+1}A rendering the diagram

[figure: sk; α; C; F_{n}^{K}(A); tk; k ∼; β; F_{n+1}^{K}(A); ι_{n+1}^{K}(A)]

commutative. Since h is the cobase change of k along α, there is a unique induced map B F_{n+1}^{K}(A) which gives f the structure of a subcomplex of

the complex A ^{ι_{1}^{K}(A)} ⋯ ^{ι_{n+1}^{K}(A)} F_{n+1}^{K}(A).

3.3.2 Enriched

Suppose that V is a cocomplete and closed symmetric monoidal category, and that K is a set of maps with finitely presentable domains and codomains. The fibrant replacement functor defined in 3.3.1 is a priori not a V-functor, but one can remedy this as follows. Each k in K induces a V-natural transformation V(sk, −)⊗k: V(sk, −)⊗sk V(sk, −)⊗tk of endo-V-functors. On the other hand, the counit ϵ_{sk}: V(sk, −)⊗sk Id_{V} is also a V-natural transformation. By taking the coproduct over all k ∈ K, one gets the diagram of V-functors

∐_{k∈K} V(sk, −) ⊗ tk ^{∐ V(sk,−)⊗k} ∐_{k∈K} V(sk, −) ⊗ sk Id_{V}.

Denote the pushout by R_{1}^{K}. It is clear that one can iterate this construction. Given a V-natural transformation ρ_{n}^{K}: R_{n−1}^{K} R_{n}^{K} of endo-V-functors of V, let R_{n+1}^{K} be the pushout of

∐_{k∈K} V(sk, R_{n}^{K}(−)) ⊗ tk ^{∐_{k∈K} V(sk,R_{n}^{K}(−))⊗k} ∐_{k∈K} V(sk, R_{n}^{K}(−)) ⊗ sk R_{n}^{K}.

The colimit of the diagram Id_{V} = R_{0} ^{ρ_{1}^{K}} R_{1}^{K} ^{ρ_{2}^{K}} ⋯ is called R^{K}: V V. Let ρ^{K}: Id_{V} R^{K} be the canonical V-natural transformation.

Lemma 3.20. Given any object A in V, the map ρ^{K}(A): A R^{K}(A) is in K ⊗ V-cell, and the map R^{K}(A) ∗ is in K-inj.

Proof. The first statement is obvious. To prove that R^{K}(A) ∗ is in K-inj, consider a lifting problem for k ∈ K:

[figure: sk; f; R^{K}(A); tk; k ∼; ∗]

Note that f factors as sk ^{g} R_{n}^{K}(A) R^{K}(A) for some n, since sk is finitely presentable and R is a sequential colimit. The adjoint of g, tensored with sk, induces the map h: sk V(sk, R_{n}^{K}(A))⊗sk such that the following diagram commutes.

\begin{aligned}& V(sk, R_{n}^{K}(A)) ⊗ sk \\ sk & \overline{^{g}} & ^{^{h}} & R_{n}^{K}(A) & ^{ϵ_{sk}(R_{n}^{K}(A))} \end{aligned}

Let h′ denote the canonical map to the coproduct given by h. Then the diagram

[figure: sk; h′; ∐_{k∈K} V(sk, R_{n}^{K}(A)) ⊗ sk; R_{n}^{K}(A); tk; k ∼; ∐_{k∈K} V(sk, R_{n}^{K}(A)) ⊗ tk; R_{n+1}^{K}(A)]

shows that tk  ∐_{k∈K} V(sk, R_{n}^{K}(A))  R_{n+1}^{K}(A)  R^{K}(A) solves

the lifting problem above. This proves that R^{K}(A) ∗ is in K-inj.

Corollary 3.21. Let V be a weakly finitely generated monoidal model category satisfying the monoid axiom. Then ρ^{J′}: Id_{V} R^{J′} is a fibrant replacement V-functor, i.e. R^{J′}(A) is fibrant and ρ^{J′}(A) is a weak equivalence for all A.

Proof. It remains to prove that the natural map ρ_{A}^{J′}: A R^{J′}(A) is a weak equivalence for every A ∈ Ob V. Note that ρ_{A}^{J′} is contained in aCof(V)⊗V-cell. The monoid axiom for V implies that ρ_{A}^{J′} is a weak equivalence.

Remark 3.22. The map ρ^{J′}(A) is not a cofibration in general, even if A is cofibrant. But if all objects in V are cofibrant, then ρ^{J′}(A) is a cofibration.

3.3.3 Filtered

Let V denote a cocomplete closed symmetric monoidal category, and K a set of maps with finitely presentable domains and codomains. We will define for each object A of V three categories ac^{K}(A), ac^{K}(A, R) and ac^{K}(A, F), and functors U_{R}: ac^{K}(A, R) ac^{K}(A), U_{F}: ac^{K}(A, F) ac^{K}(A). Objects in ac^{K}(A) are finite K-cell complexes β: A B relative to A. The morphisms in ac^{K}(A) from β: A B to γ: A C are finite K-cell complexes τ: B C for which τ ∘β = γ. The identity id_{A} is the initial object of ac^{K}(A). Consider the functor Ψ_{A}: ac^{K}(A) V which sends β: A B to B and τ: B C to τ. The colimit of Ψ_{A} will define the desired fibrant replacement of A, up to isomorphism. Objects in ac^{K}(A, R) are pairs (β: A B, β′: B R^{K}(A)) such that β is a finite K-cell complex relative to A, and β′ ∘ β = ρ^{K}(A) holds. A map

from (β: A  B, β′: B  R^{K}(A)) to (γ: A ^{∼} C, γ′: C ^{∼} R^{K}(A)) is

a finite K-cell complex τ: B C such that

B
[figure: A; β; R^{K}(A); β′]
[figure: C; τ; γ′; γ]

commutes. The initial object of ac^{K}(A, R) is (id_{A}, ρ_{A}). Denote the forgetful functor ac^{K}(A, R) ac^{K}(A) which maps the pair (β, β′) to β by U_{R}. The category ac^{K}(A, F) has objects finite subcomplexes of the relative K-cell complex ι^{K}(A): A F^{K}(A). Such a subcomplex β: A B comes with a map β′: B F^{K}(A) in K-cell such that β′ ∘ β = ι^{K}(A), whence the objects in ac^{K}(A, F) are also denoted (β, β′). Maps are defined as for ac^{K}(A, R). Here (id_{A}, ι(A)) is the initial object. Let U_{F}: ac^{K}(A, F) ac^{K}(A) be the forgetful functor. Recall from [12, IX.3] the notion of a final functor.

Lemma 3.23. If maps in K-cell are monomorphisms, then U_{R} and U_{F} are final functors, and ac^{K}(A, F) is a small filtered preorder for all A ∈ Ob V.

Proof. To prove that U_{R} is final, let β: A B be an object of ac^{K}(A). Since R^{K}(A) ∗ is in K-inj, there exists a lift in the following diagram.

[figure: A; ρ^{K}(A); R^{K}(A); B; β; ∗]

Hence the comma category β ↓ U_{R} is nonempty. Consider objects (γ, γ′) and

(δ, δ′) in ac^{K}(A, R), and maps σ: β  γ, τ: β  δ in ac^{K}(A). The pushout

[figure: B; τ; D; C; σ; C ∪_{B} D]

yields a finite K-cell complex α: A C ∪_{B} D. The maps γ′ and δ′ induce a map α′: C ∪_{B} D R^{K}(A) turning (α, α′) into an object of ac^{K}(A, R). The cobase changes of σ and τ are maps (γ, γ′) (α, α′) and (δ, δ′) (α, α′) in ac^{K}(A, R). This implies that β ↓ U_{R} is connected. Hence U_{R} is final. Next we consider U_{F}. Let β: A B be an object of ac^{K}(A). By 3.19, β ↓ U_{F} is nonempty. Since maps in K-cell are monomorphisms, the union of two finite subcomplexes is again a finite subcomplex [6, 12.2.1]. Connectness follows. A category is a preorder if there is at most one map between any two objects.

Suppose that there exist two different maps σ: B  C and τ: B  C

from (β: A  B, β′: B  F^{K}(A)) to (γ: A  C, γ′: C  F^{K}(A)).

By definition, we have γ′ ∘ σ = β′ = γ′ ∘ τ. Since γ′ is a monomorphism by assumption, it follows that σ = τ. Note that ac^{K}(A, F) is nonempty, since it contains the initial object (id_{A}, ι_{A}). Hence to prove that ac^{K}(A, F) is filtered, it remains to observe that any two objects have a common upper bound, given by the union of subcomplexes, cf. [6, 12.2.1]

The colimits of Ψ_{A}, Ψ_{A}∘U_{R} and Ψ_{A}∘U_{F} are isomorphic via the canonical maps colimΨ_{A} ∘ U_{R} colimΨ_{A} colimΨ_{A} ∘ U_{F} by Theorem 1 of [12, IX.3]. To have a natural comparison with the V-functor R^{K}, we let Φ^{K}(A) be the colimit of the functor Ψ_{A} ∘ U_{R}: ac^{K}(A, R) V. There is a canonical map φ_{A}: A Φ^{K}A induced by the object (id_{A}, ρ^{K}(A)) of ac^{K}(A, R). Likewise there is a canonical map ω_{A}: Φ^{K}(A) R^{K}(A) induced by the maps β′.

Proposition 3.24. Suppose that relative K-cell complexes are monomorphisms. Then Φ^{K}(A) ∗ is in K-inj, and φ_{A}: A Φ^{K}(A) is a filtered colimit of finite relative K-cell complexes. Moreover, the assignment A Φ^{K}(A) defines a functor Φ^{K}, and the maps A ^{φ_{A}} Φ^{K}(A) and Φ^{K}(A) ^{ω_{A}} R^{K}(A) are natural transformations Id_{V} ^{φ} Φ^{K} ^{ω} R^{K} such that ω ∘ φ = ρ^{K}.

Proof. The object Φ^{K}(A) is isomorphic to Ã := colimΨ_{A} ∘ U_{F}. Hence to prove the first claim, it suffices to consider a lifting problem

[figure: sk; f; Ã; tk; k ∼; ∗]

The object sk is finitely presentable, and Ã is the colimit of a filtered diagram

by 3.23. Hence f factors as sk ^{g} a  Ã for some finite subcomplex
(β: A  B, β′: B  F^{K}(A)) in ac^{K}(A, F). Take the pushout of

[figure: sk; g; B; tk; k; h; C; τ]

and define γ := τ ∘ β: A C. Since τ is the cobase change of k ∈ K, 3.19

implies that γ: A  C is a finite subcomplex of ι_{A} such that τ: β  γ is
a map in ac^{K}(A, F). The canonical map C  Ã belongs to γ, and the map
tk ^{h} C  Ã solves the lifting problem. Hence Ã  ∗ is in K-inj.

To prove the second assertion, let ac^{K}(A, F) π_{0}ac^{K}(A, F) be the canonical projection onto the set of connected components of ac^{K}(A, F). This functor is final, and the constant diagram c_{A}: ac^{K}(A, F) V with value A factors it. Since ac^{K}(A, F) is connected, it follows that colimc_{A} ≅ A. There is a natural transformation of diagrams c_{A} Ψ_{A} ∘ U_{F} with value β at (β, β′). The induced map colimc_{A} Φ^{K}(A) coincides with the canonical map φ_{A}: A Φ^{K}(A) up to the isomorphism above. The proof of the remaining claim is clear and will be left to the reader.

Corollary 3.25. Let V be a weakly finitely generated monoidal model category satisfying the monoid axiom. If relative J′-cell complexes are monomorphisms, then

A ^{φ_{A}} Φ^{J′}(A) ^{ω_{A}} R^{J′}(A)

are weak equivalences with fibrant codomains.

If V is pointed by ∗, any endo-V-functor on V maps the point to the point. On the other hand, ι_{∗} is not an isomorphism and there may be non-trivial finite subcomplexes of ι_{∗}. So Φ^{K} is not a V-functor in general. To define the stable model structure we need Φ to be “enriched with respect to spheres”. This requires a natural map T ⊗ Φ(A) Φ(T ⊗ A) for some finitely presentable object T, such that

[figure: T ⊗ Φ^{K}(A); Φ^{K}(T ⊗ A); T ⊗ R^{K}(A); T ⊗ω_{A}; sw_{T}^{R}(A); R^{K}(T ⊗ A); ω_{T ⊗A}]

is commutative. Here the lower horizontal map is the adjoint of

T ^{η_{A}(T)} V(A, T ⊗ A) ^{hom_{A,T ⊗A}^{R^{K}}} V(R^{K}(A), R^{K}(T ⊗ A)).

More details can be found in Appendix A. Suppose T ⊗ − maps finite K-cell complexes relative to A to finite K-cell complexes relative to T ⊗ A, for all A ∈ Ob V. Then tensoring with T defines a functor Θ_{A}: ac^{K}(A, R) ac^{K}(T ⊗

A, R) by sending (β: A  B, β′: B  R^{K}(A)) to
(T ⊗ A ^{T ⊗β} T ⊗ B, T ⊗ B ^{T ⊗β′} T ⊗ R^{K}(A) ^{r(T,A)} R^{K}(T ⊗ A)).

This functor induces a map

T ⊗ Φ^{K}(A) = T ⊗ colim_{ac^{K}(A,R)}  Ψ_{A} ∘ U_{R}
≅  colim_{ac^{K}(A,R)} (T ⊗ −) ∘ Ψ_{A} ∘ U_{R}
=  colim_{ac^{K}(A,R)} Ψ_{T ⊗A} ∘ U_{R} ∘ Θ_{A}
\overline{}colim_{ac(T ⊗A,R)} Ψ_{T ⊗A} ∘ U_{R}
=  Φ^{K}(T ⊗ A)

via the identity natural transformation (T ⊗−)∘Ψ_{A} ∘U_{R} Ψ_{T ⊗A} ∘U_{R} ∘Θ_{A}.

Lemma 3.26. Let V be a monoidal model category satisfying the monoid axiom. Suppose that V is weakly finitely generated, and that relative J′-cell complexes are monomorphisms. If A is an object of V, the map θ_{A}: T ⊗ Φ^{J′}(A) Φ^{J′}(T ⊗ A) is a weak equivalence making the diagram

[figure: T ⊗ Φ^{J′}(A); θ_{A}; Φ^{J′}(T ⊗ A); T ⊗ R^{J′}(A); T ⊗ω_{A}; r(T,A); R^{J′}(T ⊗ A); ω_{T ⊗A}]

commutative. Furthermore, θ_{A} is natural in A.

Proof. Commutativity and naturality follow by the construction of θ_{A}. Note that ρ_{T ⊗A} and T ⊗ρ_{A}^{J′} ∈ aCof(V)⊗V-cell are weak equivalences. It follows that r(T, A) is a weak equivalence since r(T, A) ∘ T ⊗ ρ_{A}^{J′} = ρ_{T ⊗A}^{J′}. The map φ_{T ⊗A} is a weak equivalence, and T ⊗ φ_{A} is a filtered colimit of weak equivalences. Hence the vertical maps in the diagram are weak equivalences.

Finally, we relate V-functors and Φ^{K} in the case where 𝕀: C ^{⊂} V is a full sub-V-category and all objects in C are V-finitely presentable.

Lemma 3.27. Suppose that relative K-cell complexes are monomorphisms. Let X: C V be a V-functor and let A ∈ Ob C. There is an isomorphism

𝕀_{∗}X(Φ^{K}(A)) ≅ colim(X ∘ Ψ_{A} ∘ U_{R}: ac^{K}(A, R) V)

which is natural in X.

Proof. To prove this, use the canonical expression of a V-functor as a coend of representables 2.3, 3.23 and V-finiteness of the objects in C. Then

𝕀_{∗}X(Φ^{K}(A)) = ∫ ^{Ob C} V(c, Φ^{K}(A))⊗X(c) ≅ colim ∫ ^{Ob C} V(c, Ψ_{A} ∘U_{R})⊗X(c),

and the claim follows.

4 The pointwise model structure

Let V be a weakly finitely generated monoidal model category. If C is a small V-category and the monoid axiom holds in V, we introduce the pointwise model structure on [C, V]. Of particular interest are the cases where C is a full sub-V-category, and C satisfies the following properties.

f0 Every object of V is a filtered colimit of objects in C.

f1 Every object of C is V-finitely presentable.

f2 The unit e is in C, and C is closed under the monoidal product in V.

4.1 The general case

Our pointwise notions of weak equivalences, fibrations and cofibrations are as follows.

Definition 4.1. A morphism f in [C, V] is a

• pointwise weak equivalence if f(c) is a weak equivalence in V for all c ∈ Ob C,

• pointwise fibration if f(c) is a fibration in V for all c ∈ Ob C,

• cofibration if f has the left lifting property with respect to all pointwise acyclic fibrations.

Theorem 4.2. Let V be a weakly finitely generated monoidal model category, and let C be a small V-category. Suppose the monoid axiom holds in V. Then [C, V], with the classes of maps in 4.1, is a weakly finitely generated model category.

Proof. We will use [7, 2.1.19]. The category [C, V] is bicomplete by 2.2. The class of pointwise weak equivalences is closed under retracts and satisfies the “two out of three” or saturation axiom. Let I be the generating cofibrations in V, and J the generating acyclic cofibrations in V. Let P_{I} be the set of maps

{V_{C}(c, −) ⊗ si ^{V_{C}(c,−)⊗i} V_{C}(c, −) ⊗ ti | i ∈ I, c ∈ Ob C}.

Likewise, let P_{J} denote the set of maps

{V_{C}(c, −) ⊗ sj ^{V_{C}(c,−)⊗j} V_{C}(c, −) ⊗ tj | j ∈ J, c ∈ Ob C}.

Since V is cofibrantly generated, it follows from adjointness that P_{J}-inj coincides with the class of pointwise fibrations, and P_{I}-inj coincides with the class of pointwise acyclic fibrations. If A is finitely presentable (small) in V, then V_{C}(c, −) ⊗ A is finitely presentable (small) in [C, V] for any c ∈ Ob C, since colimits are computed pointwise according to 2.2. Hence the smallness conditions listed in 3.4 are satisfied. It remains to show that maps in P_{J}-cell are pointwise weak equivalences. Every map in P_{J}-cell is pointwise a map in J ⊗ V-cell, and the latter class consists of weak equivalences by the monoid axiom.

We refer to the model structure in 4.2 as the pointwise model structure. Note that the evaluation functor preserves fibrations and acyclic fibrations.

Lemma 4.3. Suppose the pointwise model structure exists. Then the functor from V to [C, V] which maps A to V_{C}(c, −) ⊗ A is a left Quillen functor, with right adjoint Ev_{c} for all c ∈ Ob C.

If the unit in V is cofibrant, then the representable functors are cofibrant in the pointwise model structure.

Theorem 4.4. Consider V and C as in 4.2. Then the pointwise model structure gives [C, V] the structure of a V-model category. Likewise, [C, V] is a monoidal V-model category provided C is a symmetric monoidal V-category, and the monoid axiom holds.

Proof. Recall from 2.4 that [C, V] is a closed V-module. By [7, 4.2.5] it suffices to check the following conditions.

• Let V_{C}(c, −) ⊗ i: V_{C}(c, −) ⊗ si V_{C}(c, −) ⊗ ti be a map in P_{I}, and let j: sj tj be a map in I. Then the pushout product (V_{C}(c, −) ⊗ i)□j is a cofibration.

• If either i or j in the above sentence are generating acyclic cofibrations, then (V_{C}(c, −) ⊗ i)□j is a pointwise acyclic cofibration.

Since V_{C}(c, −) ⊗ − is a left adjoint, the pushout product map in question is of the form V_{C}(c, −) ⊗ (i□j). Hence the conditions hold, because V is a monoidal model category and 4.3 holds. The monoidality statement is proven similarly using [7, 4.2.5]. Note from 2.7 and the compatibility of ∧ and ⊗, that the pushout product map of V_{C}(c, −)⊗i and V_{C}(d, −) ⊗ j is isomorphic to V_{C}(c ⋄ d, −) ⊗ (i□j) where ⋄ denotes the monoidal product in C. Let u be the unit of C. Then V_{C}(u, −) ⊗ − is a strict symmetric monoidal functor and a left Quillen functor by 4.3. It remains to prove the monoid axiom. Abbreviate [C, V] by F. Since F is cofibrantly generated, it suffices to check that every map in the class P_{J} ∧ Fcell is a pointwise weak equivalence. Let c ∈ Ob C, j ∈ J and X: C V a V-functor. Then (V(c, −) ⊗ j) ∧ X coincides up to isomorphism with the map (V(c, −)∧X)⊗j. In particular, ((V(c, −)∧X)⊗j)(d) = (V(c, −)∧X)(d)⊗j is contained in J ⊗ V for every d ∈ Ob C, X ∈ Ob F. For a map f in P_{J} ⊗ F-cell, f(d) belongs to J ⊗ V-cell because colimits are formed pointwise. Since the monoid axiom holds in V, f is a pointwise weak equivalence.

Remark 4.5. Via 2.12, the pointwise model structure on T -spectra corresponds to the pointwise model structure on [T Sph, V].

For a discussion of properness of the pointwise model structure, we introduce the following definition. Let Cof(V) denote the class of cofibrations in V.

Definition 4.6. A monoidal model category V is strongly left proper if the cobase change of a weak equivalence along any map in Cof(V) ⊗ V-cell is again a weak equivalence.

Strongly left proper monoidal model categories are left proper. If a model category has only cofibrant objects, it is left proper. If a monoidal model category has only cofibrant objects, it is strongly left proper. The relevance of 4.6 is explained by the following lemma.

Lemma 4.7. Consider V and C as in 4.2. If f is a cofibration in [C, V], then f(c) is a retract of a map in Cof(V) ⊗ V-cell for every c ∈ Ob C.

Proof. Any cofibration in [C, V] is a retract of a relative P_{I}-cell complex, and V_{C}(d, c) ⊗ i is a map in I ⊗ V-cell for every c ∈ Ob C. The claim follows.

Corollary 4.8. The pointwise model structure on [C, V] is right proper if V is right proper, and left proper if V is strongly left proper.

4.2 The subcategory case

The goal in this section is to develop conditions under which smashing with cofibrant V-functors preserves pointwise weak equivalences. This fact will be used to prove the monoid axiom for the stable model structure. Recall that if C is a full sub-V-category of V, the left Kan extension along the inclusion functor 𝕀: C ^{⊂} V is

𝕀_{∗}X = ∫ ^{Ob C} V(c, −) ⊗ X(c).

Lemma 4.9. Assume V is a weakly finitely generated monoidal model category, and C is a small and full sub-V-category satisfying f0 and f1. If f is a pointwise weak equivalence in [C, V], then so is 𝕀_{∗}f.

Proof. We have to check that, for any A ∈ Ob V, 𝕀_{∗}f(A) is a pointwise weak equivalence. Write A as the colimit of C: I C for I filtering by f0. Since coends commute with colimits and f1 holds, it follows that

𝕀_{∗}f(A)  =  ^{∫ Ob (C)} V(B, A) ⊗ f(B)
≅  ^{∫ Ob (C)}V(B, colim_{i∈I}C) ⊗ f(B)
≅ colim_{i∈I}^{∫ Ob (C)} V(B, C(i)) ⊗ f(B)
≅ colim_{i∈I} f ∘ C.

Note that, since f is a pointwise weak equivalence, 𝕀_{∗}f(A) is a filtered colimit of weak equivalences and hence a weak equivalence by 3.5.

Corollary 4.10. Let V and C be as in 4.9. If C satisfies f2 and f is a pointwise weak equivalence in [C, V], then f ∧ V(c, −) is a pointwise weak equivalence for all c ∈ Ob C.

Proof. Axiom f2 implies that the smash product of two enriched functors exists. By 2.8, f ∧ V(c, −) is isomorphic to 𝕀_{∗}f ∘ V(c, −). Since 𝕀_{∗}f is a pointwise weak equivalence by 4.9, 𝕀_{∗}f(V(c, d)) is a weak equivalence for all d ∈ Ob C.

Theorem 4.11. Let V and C be as in 4.10. Assume V satisfies the monoid axiom and is strongly left proper, and tensoring with the domains and the codomains of the generating cofibrations in V preserves weak equivalences. Then smashing with a cofibrant object in [C, V] preserves pointwise weak equivalences.

Proof. Let f: X Y be a pointwise weak equivalence, and let Z be a cofibrant V-functor. Since Z is cofibrant, it is a retract of some V-functor Z′, such that ∗ Z′ is the sequential composition of a sequence

∗ = Z′_{0} ^{g_{0}} Z′_{1} ^{g_{1}} ⋯ ^{g_{n−1}} Z′_{n} ^{g_{n}} ⋯,

where g_{n} is the cobase change of a coproduct of maps in P_{I}. It suffices to consider sequences indexed by the natural numbers, since the domains of the maps in P_{I} are finitely presentable. The map f ∧ Z is a retract of f ∧ Z′, hence it remains to prove that the latter is a pointwise weak equivalence. We will prove this by induction on n. Consider the following diagram.

[figure: X ∧ ∐_{m∈M} V(c_{m}, −) ⊗ ti_{m}; X∧∐ V(c_{m},−)⊗i_{m}; X ∧ ∐_{m∈M} V(c_{m}, −) ⊗ si_{m}; X ∧ Z_{n}; Y ∧ ∐_{m∈M} V(c_{m}, −) ⊗ ti_{m}; f∧ ∐ V(c_{m},−)⊗ti_{m}; X∧∐ V(c_{m},−)⊗i_{m}; Y ∧ ∐_{m∈M} V(c_{m}, −) ⊗ si_{m}; f∧ ∐ V(c_{m},−)⊗si_{m}; Y ∧ Z_{n}.; f∧Z_{n}]

The map induced on the pushouts of the upper and lower row is f ∧ Z_{n+1}. Suppose f ∧ Z′_{n} is a pointwise weak equivalence. By 4.10 and the hypothesis on I, it follows that both of the other vertical maps are pointwise weak equivalences. The horizontal maps on the left hand side are not necessarily cofibrations. However, evaluation at any object gives maps in Cof(V) ⊗ Vcell. Since V is strongly left proper, this implies that the map induced on the pushouts – which is computed pointwise – is a pointwise weak equivalence.

Remark 4.12. Let V be a weakly finitely generated monoidal model category. Suppose that − ⊗ si and − ⊗ ti preserve weak equivalences for every i ∈ I, and V is strongly left proper. Then − ⊗ A preserves weak equivalences for any cofibrant object A in V, cp. 4.11. We say a strongly left proper monoidal model category is strongly monoidal if −⊗A preserves weak equivalences for A either cofibrant or a domain or codomain of the generating cofibrations, if they exist. A monoidal model category in which every object is cofibrant satisfies this condition and also the monoid axiom.

5 The homotopy functor model structure

Suppose F: C D is a functor of categories with chosen subclasses of weak equivalences. If F maps weak equivalences to weak equivalences, then F is called a homotopy functor. As a first step towards the stable model structure on enriched functors, we define a model structure in which every enriched functor is weakly equivalent to a homotopy functor. Let V be a weakly finitely generated strongly monoidal sSet-model category. Additionally, assume the following for V: the monoid axiom holds, the unit is cofibrant, Δ^{1} is finitely presentable in V, filtered colimits commute with pullbacks, and cofibrations are monomorphisms. The simplicial structure is used for the simplicial mapping cylinder construction. Let 𝕀: fV ^{⊂} V be a small full sub-V-category such that the following axioms hold.

f1 Every object of fV is V-finitely presentable.

f2 The unit e is in fV, and fV is closed under the monoidal product.

f3 If tj ^{∼} sj v is a diagram in V where v ∈ Ob fV and j ∈ J′, then the pushout tj ∪_{sj} v is in fV.

Consider also the following additional axioms.

f0 Every object of V is a filtered colimit of objects in fV.

f4 All objects in fV are cofibrant.

f5 The simplicial mapping cylinder exists in fV.

Objects in fV will usually be denoted by small letters, since fV should be thought of as a category of small objects. Let F be short for [fV, V]. Recall the left Kan extension 𝕀_{∗}X of X ∈ Ob F along 𝕀: fV ^{⊂} V, cp. 2.5.

5.1 Equivalences of homotopy functors

Let Φ: fV V denote the functor induced by the fibrant replacement functor Φ^{J′} from 3.3.3. In general it is not a V-functor. Denote by ℏ(X): fV V the composition 𝕀_{∗}X ∘ Φ, and by ℏ: F Fun(fV, V) the induced functor. There is a natural transformation X ℏ(X) induced by the canonical maps φ_{v}: v Φ(v) where v varies through the set of objects in fV.

Lemma 5.1. The functor ℏ commutes with colimits and the action of V. The natural transformations X ℏ(X) define a natural transformation from the forgetful functor F Fun(fV, V) to ℏ.

Definition 5.2. A map f is an hf-equivalence if ℏ(f)(v) is a weak equivalence in V for all v ∈ Ob fV.

Lemma 5.3. Any pointwise weak equivalences is an hf-equivalence. The class of hf-equivalences is saturated.

Proof. The first statement follows as in 4.9, since ℏ(f)(v) = 𝕀_{∗}(f)(Φ(v)) is a filtered colimit of weak equivalences provided f is a pointwise weak equivalence. The second statement follows from 5.1 and the analogous fact in V.

Lemma 5.4. Let f be a cofibration in F. Then ℏ(f)(v) is a retract of a map in Cof(V) ⊗ V-cell for every v ∈ Ob fV.

Proof. If V(v, −) ⊗ i is a generating cofibration, then ℏ(V(v, w) ⊗ i) coincides with V(v, Φ(w)) ⊗ i, which is in Cof(V) ⊗ V-cell. The general case follows since ℏ commutes with colimits.

5.2 Fibrations of homotopy functors

If ϕ: v w is an acyclic cofibration in fV, the simplicial mapping cylinder factors V(ϕ, −): V(w, −) V(v, −) as a cofibration c_{ϕ}: V(w, −) C_{ϕ} followed by a simplicial homotopy equivalence. This uses that V(w, −) is a cofibrant functor (since the unit in V is cofibrant), and that V (and hence F by 4.4) is a sSet-model category. Take a generating cofibration i: si ti in V and form the pushout product

c_{ϕ}□i: V(w, −) ⊗ ti ∪_{V(w,−)⊗si} C_{ϕ} ⊗ si C_{ϕ} ⊗ ti.

Let H denote the set {c_{ϕ}□i}, where ϕ runs through the set of acyclic cofibrations in fV and i runs through the set I of generating cofibrations in V.

Definition 5.5. A map is an hf-fibration if it is a pointwise fibration having the right lifting property with respect to H.

Lemma 5.6. Let f: X Y be a pointwise fibration. Then f is an hf-fibration if and only if the following diagram is a homotopy pullback square in V for every acyclic cofibration ϕ: v ^{∼} w in fV.

[figure: X(v); X(ϕ); X(w); Y (v); f(v); Y (ϕ); Y (w); f(w)]

Proof. Let V_{F}(X, Y) denote the V-object of maps in F from X to Y. For a map of V-functors f: X Y, the square in the statement of the lemma is naturally isomorphic, by the Yoneda lemma 2.1, to the square

[figure: V_{F}(V(v, −), X); V_{F} (V(ϕ,−),X); V_{F}(V(w, −), X); V_{F}(V(v, −), Y); V_{F} (V(v,−),f); V_{F} (V(ϕ,−),Y); V_{F}(V(w, −), Y).; V_{F} (V(w,−),f)]

The factorization of V(ϕ, −) as a cofibration c_{ϕ}: V(w, −) C_{ϕ} followed by a simplicial homotopy equivalence C_{ϕ} V(v, −) induces a factorization of the square above into two squares. Since V_{F}(−, X) preserves simplicial homotopy equivalences by 2.11, which are pointwise weak equivalences by 3.10, the square above is a homotopy pullback square if and only if

[figure: V_{F}(C_{ϕ}, X); V_{F} (c_{ϕ},X); V_{F}(V(w, −), X); V_{F}(C_{ϕ}, Y); V_{F} (C_{ϕ},f); V_{F} (c_{ϕ},Y); V_{F}(V(w, −), Y); V_{F} (V(w,−),f)] is a homotopy pullback square. If f: X Y is a pointwise fibration, the induced map g: V_{F}(C_{ϕ}, X) V_{F}(C_{ϕ}, Y) ×_{V_{F} (V(w,−),Y)} V_{F}(V(w, −), X) is a fibration in V. Here we use f2, so F is a monoidal model category when equipped with the pointwise model structure, see 4.4. The square in 5.6 is therefore a homotopy pullback square if and only if g has the right lifting property with respect to the generating cofibrations in V. By adjointness, this holds if and only if f has the right lifting property with respect to H.

Lemma 5.7. Let f: X Y be an hf-fibration. Then

[figure: X(v); ℏ(X)(v); Y (v); f(v); ℏ(Y)(v); ℏ(f)(v)]

is a homotopy pullback square in V for every object v of fV.

Proof. Let f: X [figure] Y be a pointwise fibration of pointwise fibrant functors. From 3.27, ℏ(f)(v) is a filtered colimit of fibrations of fibrant objects in V, and therefore a fibration of fibrant objects by 3.5. This uses properties f1 and f3. The square in the lemma is therefore a homotopy pullback square if and only if X(v) Y (v) ×_{colimY (a)} colimX(a) is a weak equivalence in V. Up to isomorphism, the colimit is taken over the category of finite subcomplexes (α: v ^{∼} a, α′: a ^{∼} F v) in Ob ac^{J′}(v, F). Note that the colimit is filtered by 3.23. Filtered colimits commute with pullbacks in V by assumption, so the map in question is a filtered colimit of maps X(v) Y (v) ×_{Y (a)} X(a) for

acyclic cofibrations α: v ^{∼} a in fV. If f is an hf-fibration, then by 5.6 the map in question is a filtered colimit of weak equivalences in V, and hence a weak equivalence. If f: X [figure] Y is any pointwise fibration, use the factorizations in the pointwise model structure to construct a commutative square

[figure: X; \begin{matrix}∼ \\ g \end{matrix}; X′; Y; f; \begin{matrix}∼ \\ h \end{matrix}; Y′; f′]

for f′ a pointwise fibration of pointwise fibrant functors, and g and f pointwise acyclic cofibrations. Note that f is an hf-fibration if and only f′ is, see 3.13 and 5.6. The maps g and h are pointwise weak equivalences, hence ℏ(g)(v) and ℏ(h)(v) are weak equivalences in V for every v. The square in question is therefore a homotopy pullback square by the previous case.

Corollary 5.8. A map is an hf-fibration and an hf-equivalence if and only if it is a pointwise acyclic fibration.

Proof. If f is an hf-fibration, it is a pointwise fibration by definition. If f is also an hf-equivalence, it is a pointwise weak equivalence by 3.13 and 5.7. A pointwise acyclic fibration is an hf-fibration by 3.13 and 5.6.

5.3 The homotopy functor theorem

Before we prove the existence of the homotopy functor model structure, let us first consider the maps in H-cell.

Lemma 5.9. The maps in H-cell are hf-equivalences.

Proof. Let ϕ: v ^{∼} w be an acyclic cofibration in fV. Then the induced map V(ϕ, −): V(w, −) V(v, −) is an hf-equivalence, because ℏ(V(ϕ, −))(v) is naturally isomorphic to V(ϕ, Φ(v)), ϕ is an acyclic cofibration and Φ(v) is fibrant. Pointwise weak equivalences are hf-equivalences by 5.3, so the map c_{ϕ}: V(w, −) C_{ϕ} is an hf-equivalence. Let i: si ti be a generating cofibration in V. Consider the diagram

[figure: V(w, −) ⊗ si; c_{ϕ}⊗si; C_{ϕ} ⊗ si; V(w, −) ⊗ ti; V(w,−)⊗i; c_{ϕ}⊗ti; C_{ϕ} ⊗ ti; C_{ϕ}⊗i]

and the pushout product map c_{ϕ}□i. The functor ℏ commutes with pushouts and the action of V by 5.1, so ℏ(c_{ϕ}□i)(u) is the pushout product map obtained from ℏ(c_{ϕ})(u) and i. Since V is strongly monoidal, it follows that ℏ(c_{ϕ})(u)□i is a weak equivalence. Hence the maps in H are hf-equivalences. The general case of a map in H-cell follows using similar arguments and 5.4.

Theorem 5.10. Let V be a weakly finitely generated monoidal sSet-model category, and let fV be a full sub-V-category satisfying f1, f2 and f3. Suppose the monoid axiom holds in V, pullbacks commute with filtered colimits in V, and Δ^{1} is finitely presentable in V. Suppose also that V is strongly monoidal, and that cofibrations in V are monomorphisms. Then F is a weakly finitely generated model category, with hf-equivalences as weak equivalences, hf-fibrations as fibrations, and cofibrations as cofibrations.

Proof. Again we use [7, 2.1.19]. The set of generating cofibrations is P_{I}, and the set of generating acyclic cofibrations is the union P_{J} ∪H. It is clear that the class of hf-equivalences is saturated and closed under retracts. It is also clear that the domains of the maps in H are finitely presentable, because finitely presentable objects are closed under pushouts and tensoring with finitely presentable objects. Here we use that Δ^{1} is finitely presentable in V. The other properties which have to be checked are either obvious or follow from 5.8, 5.9 and the corresponding fact for the pointwise model structure 4.2.

The model structure in 5.10 is called the homotopy functor model structure. To emphasize the model structure, we use the notation F_{hf}. An hf-equivalence is denoted by ^{∼hf} and an hf-fibration by ^{hf} [figure] . Likewise, we use the notations F_{pt}, ^{∼pt} and ^{pt} [figure] for the pointwise model structure.

Lemma 5.11. The identity induces a left Quillen functor Id_{F}: F_{pt} F_{hf}.

Lemma 5.12. Let V and fV be as in 5.10, and assume fV satisfies f4. Then F_{hf} is a monoidal F_{pt}-model category.

Proof. Condition f2 is used to construct the smash product on F, and 4.4 holds. To prove that the homotopy functor model structure is monoidal, it suffices to show that the pushout product map of a map c_{ϕ}□i in H (where ϕ: v ^{∼} w) and a generating cofibration V(u, −) ⊗ j is an hf-equivalence. It is straightforward to check that this pushout product map coincides with the pushout product map c_{ϕ⊗u}□f where f is the pushout product map in V of i and j. Since V is a monoidal model category and fV satisfies f2 and f4, ϕ ⊗ u is an acyclic cofibration in fV. Hence the map in question is an hf-equivalence.

Lemma 5.13. The homotopy functor model structure is left proper. If V is right proper, then the homotopy functor model structure is right proper.

Proof. For left properness, let i: Y Z ∪_{X} Y be the cobase change of an hf-equivalence g: X ^{∼hf} Z along a cofibration f: X Y. Factor g as a cofibration h: X T, followed by a pointwise acyclic fibration p: T ^{∼} [figure] Z. Then g is an hf-equivalence, hence an acyclic cofibration in the homotopy functor model structure. These maps are closed under cobase changes, so i factors as an acyclic cofibration, followed by the cobase change of p. The latter is a pointwise weak equivalence, since the cobase change of f along g is a cofibration and the pointwise model structure is left proper by 4.8 provided V is strongly left proper. A slightly stronger property than right properness holds. Consider the maps f: X ^{pt} [figure] Z and g: Y ^{∼hf} Z. We claim the base change i of g along f is an hf-equivalence. Let us shorten the notation by setting R = R^{J′}. To prove that i is an hf-equivalence, factor Rf: RX RZ as h: RX ^{∼ pt} T followed by p: T ^{pt} [figure] RZ. Then ℏ(p)(v) is a fibration of fibrant objects for any v by 3.5. Moreover, ℏ(h)(v) and ℏ(ρ_{Z(v)}): ℏ(Z)(v) ℏ(RZ)(v) are weak equivalences for any v. Hence the base change of the weak equivalence ℏ(ρ_{Z(v)}) ∘ ℏ(g)(v) along ℏ(b)(v) is a weak equivalence, using that V is right proper. Note that the base change map factors as ℏ(h)(v) composed with the base change of ℏ(g)(v) along ℏ(f)(v), i.e. ℏ(i)(v) since pullbacks commute with filtered colimits. It follows that i is an hf-equivalence.

5.4 Homotopy functors

We end this section with a discussion of homotopy functors. If f4 and f5 hold, then a fibrant V-functor X in F_{hf} is a homotopy functor for the following reasons: By 5.6, X maps acyclic cofibrations to weak equivalences. Using f4 and f5, every weak equivalence in fV can be factored as an acyclic cofibration in fV, followed by a simplicial homotopy equivalence in fV. It follows from 2.11 that X preserves arbitrary weak equivalences in fV. Conversely, any V-functor which is pointwise fibrant and a homotopy functor is fibrant in F_{hf}. Therefore we regard the expressions “pointwise fibrant homotopy functor” and “fibrant in F_{hf}” as synonymous. Next we define a fibrant replacement functor in F_{hf} which allows to replace our definition of hf-equivalence by a better one.

Definition 5.14. For X ∈ Ob F, define X^{h} as the composition 𝕀_{∗}X ∘R, where 𝕀: fV ^{⊂} V is the inclusion and R := R^{J′} is the fibrant replacement V-functor constructed in 3.3.2.

For pointed simplicial sets, Lydakis [11, 8.6] uses the singular complex applied to the geometric realization as an enriched fibrant replacement functor. This functor preserves fibrations, weak equivalences and finite limits.

Lemma 5.15. The map X X^{h} is a V-natural transformation of Vfunctors, and extends to a natural transformation Id_{F} (−)^{h}. The functor (−)^{h} commutes with colimits and the action of V.

Proof. The first two statements follow from 3.25 and properties of enriched Kan extension. For the last statement, use that coends commute with colimits and the action of V, which are pointwise constructions.

Lemma 5.16. Assume f4 and let X ∈ Ob F be cofibrant. For every object v in fV, the weak equivalence ω_{v}: Φ(v) ^{∼} R(v) induces a weak equivalence ℏ(X)(v) X^{h}(v).

Proof. Since V is monoidal, f4 implies that V(w, Φv) V(w, Rv) induced by the canonical weak equivalence ω_{v}: Φv ^{∼} R(v) is a weak equivalence for every w ∈ fV. Now express X as a retract of a P_{I}-cell complex. The lemma follows then by induction, because V is strongly monoidal.

Corollary 5.17. Assume f4 holds and X is cofibrant. Then R ∘ X^{h} is a pointwise fibrant homotopy functor and X R ∘X^{h} is an hf-equivalence. A map f of cofibrant functors is an hf-equivalence if and only if f^{h} is a pointwise weak equivalence.

Proof. The second claim follows directly from 5.16, while the first claim requires just a slight variation of the proof of 5.16.

Lemma 5.18. Suppose C satisfies f0 and f4. For any X ∈ Ob V, the canonical map ℏ(X)(v) X^{h}(v) induced by the weak equivalence ω_{v}: Φ(v) ^{∼} R(v) is then a weak equivalence for all v ∈ Ob fV.

Proof. Note that 4.9 holds. Hence, by a cofibrant replacement in F_{pt}, it suffices to consider cofibrant functors. This case follows from 5.16.

Corollary 5.19. Assume f0 and f4 hold. A map f is then an hf-equivalence if and only if f^{h} is a pointwise weak equivalence. Furthermore, for any X ∈ Ob F, the functor R∘X^{h} is a pointwise fibrant homotopy functor and the natural map X R ∘ X^{h} is an hf-equivalence.

Proof. As in 5.17, using 4.9.

Remark 5.20. If ρ: Id_{V} R has the property that its restriction to fV takes values in cofibrant objects, then X^{h} is always a homotopy functor. The reason is that in a sSet-model category, any weak equivalence of fibrant and cofibrant objects is a simplicial homotopy equivalence, and V-functors preserve them.

6 The stable model structure

We will construct the stable model structure on the category F = [fV, V] with respect to some cofibrant object T of fV. For this, assume V and fV are as in 5.10. In addition, V has to be right proper and cellular. We also require that fV satisfies f4, in order to have a well-behaved fibrant replacement in F_{hf}. Finally, we assume the adjoint pair (−⊗T, V(T, −)) is a Quillen equivalence on the stable model structure on spectra described in 6.16. Since T is contained in fV and f2 holds, the canonical functor π: T Sph V factors over the inclusion as i: T Sph fV. Let (i_{∗}, ev) denote the corresponding adjoint pair of functors.

6.1 Stable equivalences

We start by describing the stabilization process. For every object v in fV, the composition of the counit ϵ_{T} V(v, −): V(T, V(v, −)) ⊗ T V(v, −) and the natural isomorphism V(T, V(v, −)) ≅ V(T ⊗ v, −) define a morphism τ_{v}: V(T ⊗ v, −) ⊗ T V(v, −) which is natural in v. If X is a V-functor, then the induced map V_{F}(τ_{v}, X): V_{F}(V(v, −), X) V_{F}(V(T ⊗ v, −) ⊗ T, X) is natural in v and X. Using the enriched Yoneda lemma 2.1, one obtains a map t_{X}(v): X(v) V(T, X(T ⊗ v)). Let Sh: F F denote the ‘shift’ functor obtained by pre-composing with the V-functor T ⊗ −: fV fV. Define 𝕋: F F to be the composition V(T, −) ∘ Sh, so that 𝕋(X)(v) = V(T, X(T ⊗ v)). The collection of the maps t_{X}(v) is a V-natural transformation t_{X}: X 𝕋(X). Let 𝕋^{∞}(X) denote the colimit of the sequence X ^{t(X)} 𝕋(X) ^{𝕋(t(X))} 𝕋(𝕋(X)) · · ·. The canonical map t_{X}^{∞}: X 𝕋^{∞}(X) yields a natural transformation t^{∞}: Id_{F} 𝕋^{∞}. The definition of stable weak equivalences uses the fibrant replacement functor Φ^{J′} considered in 3.3.3. Let Φ be short notation for Φ^{J′}, and similarly for the other fibrant replacement functors R and F. Recall that ℏ(X) is not necessarily a V-functor for every X ∈ Ob F. But it can be stabilized, since there are natural weak equivalences θ_{v}: T ⊗ Φ(v) Φ(T ⊗ v) according to 3.26. Let 𝕋′: Fun(fV, V) Fun(fV, V) be the functor that maps X to the composition V(T, −) ∘ X ∘ (T ⊗ −): fV V. Define the map t′_{ℏ(X)}: ℏ(X) 𝕋′(ℏ(X)) pointwise as the adjoint of

𝕀_{∗}X(Φ(v)) ⊗ T  𝕀_{∗}X(T ⊗ Φ(v))  𝕀_{∗}X(Φ(T ⊗ v)).

The map on the left hand side in this composition is adjoint to the composition

T ^{η_{Φ(v)}(T)} V(Φ(v), T ⊗ Φ(v)) ^{hom_{Φ(v),T ⊗Φ(v)}^{𝕀∗X}} V(𝕀_{∗}X(Φ(v)), 𝕀_{∗}X(T ⊗ Φ(v))), and the map on the right hand side is 𝕀_{∗}X(θ_{v})

Lemma 6.1. There is a natural transformation t′_{ℏ}: ℏ 𝕋′ ∘ ℏ. The natural transformation u: U ℏ, where U: F Fun(fV, V) is the forgetful functor, makes the following diagram commutative.

[figure: U ∘ Id_{F}; U∘t; U ∘ T = 𝕋′ ∘ U; ℏ; u; t′_{ℏ}; 𝕋′ ∘ ℏ.; 𝕋′∘u]

Proof. The claim follows since t_{X}(v): X(v) 𝕋(X)(v) can be defined as the adjoint (under tensoring with T) of the adjoint (under tensoring with X(v)) of

hom_{v,T ⊗v}^{X} ∘ η_{v}(T): T  V(v, T ⊗ v)  V(X(v), X(T ⊗ v)), cp. A.8.

Denote the colimit of ℏ(X) ^{t′_{ℏ(X)}} 𝕋′(ℏ(X)) ^{𝕋′(t′_{ℏ(X)})} 𝕋′(𝕋′(ℏ(X))) ⋯ by 𝕋′^{∞}(ℏ(X)), and let t′_{ℏ(X)}^{∞} be the canonical map ℏ(X) 𝕋′^{∞}(ℏ(X)).

Definition 6.2. A map f in F is a stable equivalence if 𝕋′^{∞}(ℏ(Rf))(v) is a weak equivalence in V for every object v of fV.

Lemma 6.3. Every hf-equivalence is a stable equivalence. The class of stable equivalences is saturated.

There are canonical maps X(v)  RX(v)  ℏ(RX)(v) for all v ∈ Ob fV.

Consider the induced map 𝕋^{∞}(X) 𝕋′^{∞}(ℏ(RX)). The latter is sometimes a pointwise weak equivalence.

Lemma 6.4. Assume X ∈ Ob F is a pointwise fibrant homotopy functor. Then 𝕋^{∞}(X)(v) 𝕋′^{∞}(ℏ(RX))(v) is a weak equivalence in V for all v ∈ Ob fV.

Proof. Note that X(v) ℏ(RX)(v) is a weak equivalence of fibrant objects. V(T, −) preserves weak equivalences of fibrant objects since T is cofibrant and V is a monoidal model category. The map is question is hence a sequential colimit of weak equivalences, so 3.5 concludes the proof.

Corollary 6.5. A map f between pointwise fibrant homotopy functors is a stable equivalence if and only if 𝕋^{∞}(f) is a pointwise weak equivalence.

6.2 Stable fibrations

Let τ_{v}: V(T ⊗ v, −) ⊗ T V(v, −) be the canonical map of V-functors described in section 6.1. The simplicial mapping cylinder factors τ_{v} as a cofibration d_{v}: V(T ⊗v, −)⊗T D_{v} followed by a simplicial homotopy equivalence. Take a generating cofibration i: si ti ∈ I in V and form the pushout product d_{v}□i. The set D of generating acyclic cofibrations for the class of stable equivalences is {d_{v}□i}, where v ∈ Ob fV and i ∈ I.

Definition 6.6. A map is called a stable fibration if it is an hf-fibration having the right lifting property with respect to the set D.

Lemma 6.7. An hf-fibration f: X Y is a stable fibration if and only if

[figure: X(v); t_{X}(v); 𝕋(X)(v); Y (v); f(v); t_{Y} (v); 𝕋(Y)(v); 𝕋(f)(v)]

is a homotopy pullback square in V for every object v of fV.

Proof. The proof is formally the same as for 5.6.

The rest of this section is devoted to prove that a stable fibration which is also a stable equivalence is a pointwise weak equivalence.

Lemma 6.8. Assume V is right proper, and that filtered colimits commute with pullbacks in V. Let f: X Y be a stable fibration. Then

[figure: ℏ(RX)(v); t′_{ℏ(RX)}(v); 𝕋′(ℏ(RX))(v); ℏ(RY)(v); ℏ(Rf)(v); t′_{ℏ(RY )}(v); 𝕋′(ℏ(RY))(v); 𝕋′(ℏ(Rf))(v)]

is a homotopy pullback square in V for all v ∈ Ob fV.

The proof of 6.8 uses 6.10, 6.11 and 6.12. We start with a general fact about model categories.

Lemma 6.9. Let G: C D be a functor between right proper model categories which preserves pullbacks, fibrations and acyclic fibrations. Suppose

[figure: A; ∼; B; C; f; ∼; D]

is a commutative diagram in C, such that the horizontal maps are weak equivalences with fibrant targets, and such that f: A [figure] C is a fibration. Then the image of this square under G is a homotopy pullback square.

Proof. By Ken Brown’s lemma, we can assume that the horizontal maps are in fact acyclic cofibrations. Factor the composition A D as an acyclic cofibration i: A ^{∼} E followed by a fibration p: E [figure] D. Then the map B D factors as B ^{∼} E ^{p} [figure] D, by choosing a lift in the diagram

[figure: A; \begin{matrix}∼ \\ i \end{matrix}; E; B; ∼; D.; p]

Define P to be the pullback of C ^{∼} D [figure]^{p} E and call the map induced by f and i h: A P. Since C ^{∼} D and i are weak equivalences and C is right proper, h is a weak equivalence. Using the assumptions on G and right properness of D (3.14), we have to prove that G(h) is a weak equivalence.

Let Q be the pullback of A ^{f} [figure] C [figure]^{p′} P, where p′ is the base change of p. The maps id_{A} and h induce a map A Q which can be factored as A \begin{matrix}s \\ ∼ \end{matrix} W ^{q} [figure] Q. After all, we have a factorization of h as A \begin{matrix}s \\ ∼ \end{matrix} W ^{f′∘q} [figure] P, where f′ is the base change of the fibration f. The map h is a weak equivalence, thus f′ ∘ q is an acyclic fibration. In particular, G(f′ ∘ q) is an acyclic fibration. The map p′′ ∘ q: W A is a fibration (where p′′ is the base change of the fibration p′) and has s: A ^{∼} W as a section. Hence p′′ ∘ q is an acyclic fibration, and so is G(p′′ ∘ q). Now G(s) is a section of G(p′′ ∘ q), implying that G(s), and therefore G(h = (f′ ∘ q) ∘ s), is a weak equivalence.

Corollary 6.10. Let f: X Y be a pointwise fibration, and assume V is right proper. Then the following is a homotopy pullback square in V for all v ∈ Ob fV.

[figure: V(T, X(v)); V(T,ρ_{X(v)}); V(T, RX(v)); V(T, Y (v)); V(T,f(v)); V(T,ρ_{Y (v)}); V(T, RY (v)); V(T,Rf(v))]

Proof. Follows from 6.9, since V(T, −): V V is a right Quillen functor.

Corollary 6.11. Suppose V is right proper and f: X Y is a stable fibration. Then the following is a homotopy pullback square for all v ∈ Ob fV.

[figure: RX(v); t_{RX}(v); 𝕋(RX)(v); RY (v); Rf(v); t_{RY} (v); 𝕋(RY)(v); 𝕋(Rf)(v)]

Proof. Consider the following commutative diagram.

X(v)  \overline{^{t_{X}(v)}} 𝕋(X)(v)

[figure: RX(v); t_{RX}(v); ρ_{X(v)}; 𝕋(f)(v); 𝕋(RX)(v); 𝕋(ρ_{X(v)})]

[figure: Y (v); f(v); t_{Y} (v); 𝕋(Y)(v)]

[figure: RY (v); Rf(v); t_{RY} (v); ρ_{Y (v)}; 𝕋(RY)(v); 𝕋(Rf)(v); 𝕋(ρ_{Y (v)})]

The right hand square of the cube is a homotopy pullback square by 6.10. Likewise for the square in the back using the assumption on f. In the left hand square the horizontal maps are weak equivalences. Hence the square in question is a homotopy pullback square by 3.13.

Lemma 6.12. Assume f: X Y is a stable fibration of pointwise fibrant functors, and filtered colimits commute with pullbacks in V. Then

[figure: ℏ(X)(v); t′_{ℏ(X)}(v); 𝕋′(ℏ(X))(v); ℏ(Y)(v); ℏ(f)(v); t′_{ℏ(Y )}(v); 𝕋′(ℏ(Y))(v); 𝕋′(ℏ(f))(v)]

is a homotopy pullback square for all v ∈ Ob fV.

Proof. Up to isomorphism, the square above decomposes into two squares:

[figure: colim_{ac(v,R)}X(a); colimt_{X}(a); colim_{ac(v,R)}𝕋(X)(a); c_{X}; colim_{ac(T ⊗v,R)}V(T, X(b))]

[figure: colim_{ac(v,R)}Y (a); colimf(a); colimt_{Y} (a); colim_{ac(v,R)}𝕋(Y)(a); colim𝕋(f)(a); c_{Y}; colim_{ac(T ⊗v,R)}V(T, Y (b)); colim V(T,f(b))]

Recall from 3.3.3 that the objects in ac(T ⊗v, R) are T ⊗v ^{∼} b R(T ⊗v). The maps c_{X} and c_{Y} are obtained from Θ_{v}: ac(v, R) ac(T ⊗ v, R) which maps v ^{∼} a R(v) to T ⊗v ^{∼} T ⊗a T ⊗R(v) R(T ⊗v), and the natural transformation V(T, X(T ⊗ Φ_{v})) V(T, X(Φ_{T ⊗v} ∘ Θ_{v})) which consists of identity maps. From 3.23, one can replace the indexing categories ac(v, R) and ac(T ⊗ v, R) by filtered ones, namely ac(v, F) and ac(T ⊗ v, F). Then all the vertical maps are fibrations of fibrant objects, because f is a pointwise fibration of pointwise fibrant functors and 3.5 holds. It follows that the left hand square is a homotopy pullback square if and only if the canonical map g from colim_{ac(v,F)}X(a) to the pullback is a weak equivalence. Filtered colimits commute with pullbacks, so g is the filtered colimit of the canonical maps induced by the squares

[figure: X(a); t_{X}(a); 𝕋(X)(a); Y (a); f(a); t_{Y} (a); 𝕋(Y)(a); 𝕋(f)(a)]

for finite sub-complexes v ^{∼} a of v ^{ι_{v}} F v. These squares are all homotopy pullback squares: f is a stable fibration, and the vertical maps are fibrations of fibrant objects. It follows that g is a filtered colimit of weak equivalences, so the left hand square is a homotopy pullback square. That the right hand square is a homotopy pullback depends on whether f is an hf-fibration. We claim that

X(T ⊗ a) \overline{}colim_{ac(T ⊗v,R)}X(b)

[figure: Y (T ⊗ a); f(T ⊗a); colim_{ac(T ⊗v,R)}Y (b); colim_{ac(T ⊗v,R)}f(b)]

is a homotopy pullback square for all v \begin{matrix}∼ \\ α \end{matrix} a R(v) in ac(v, R). Denote

the full subcategory of ac(T ⊗ v, R) consisting of T ⊗ v \begin{matrix}∼ \\ β \end{matrix} b R(T ⊗ v),

where β factors as T ⊗ v ^{T ⊗α} T ⊗ a b, by ac(T ⊗ v, R)_{a}. This category is a final subcategory of ac(T ⊗ v, R). Hence we may assume the colimit in the square above is indexed by ac(T ⊗ v, R)_{a}. As in the proof of 5.7, it follows that the square above is a homotopy pullback square for all v \begin{matrix}∼ \\ α \end{matrix} a R(v) in ac(v, R). Here we use that f is an hf-fibration. The right hand square in the main diagram is then a homotopy pullback square, using the by now standard argument for filtered colimits of homotopy pullback squares.

A proof of 6.8 follows:

Proof. By 6.11, we may assume f is a stable fibration of pointwise fibrant functors. The result is therefore a consequence of 6.12.

Lemma 6.13. Suppose that V is right proper and filtered colimits commute with pullbacks in V. Let f: X Y be a stable fibration. Then

[figure: X(v); 𝕋′^{∞}(ℏ(RX))(v); Y (v); f(v); 𝕋′^{∞}(ℏ(RY))(v); 𝕋′^{∞}(ℏ(Rf))(v)]

is a homotopy pullback square in V for all v ∈ Ob fV.

Proof. Let f: X Y be a stable fibration. Factor Rf as RX ^{∼ pt} Z followed by g: Z ^{pt} [figure] RY. Here g is a stable fibration of pointwise fibrant functors by 5.6 and 6.11. The claim in 6.13 is equivalent to the statement that

[figure: Z(v); 𝕋′^{∞}(ℏ(Z))(v); RY (v); g(v); 𝕋′^{∞}(ℏ(RY))(v); 𝕋′^{∞}(ℏ(g))(v)]

is a homotopy pullback square. The vertical maps are fibrations of fibrant objects, and there is the decomposition

[figure: Z(v); ℏ(Z)(v); 𝕋′^{∞}(ℏ(Z))(v); RY (v); g(v); ℏ(RY)(v); ℏ(g)(v); 𝕋′^{∞}(ℏ(RY))(v); 𝕋′^{∞}(ℏ(g))(v)]

The left hand square is a homotopy pullback square by 5.7. The functor V(T, −) preserves homotopy pullback squares provided the vertical maps are fibrations of fibrant objects. It follows, using 6.8, that the right hand square is a homotopy pullback square.

Corollary 6.14. A map is a stable fibration and a stable equivalence if and only if it is a pointwise acyclic fibration.

Proof. Let f: X Y be a stable fibration and a stable equivalence. Stable fibrations are in particular pointwise fibrations, so it remains to prove that f is a pointwise weak equivalence. By 6.13, the following diagram is a homotopy pullback square in V for all v ∈ Ob fV.

[figure: X(v); 𝕋′^{∞}(ℏ(RX))(v); Y (v); f(v); 𝕋′^{∞}(ℏ(RY))(v); 𝕋′^{∞}(ℏ(Rf))(v)]

Since f is a stable equivalence, the right hand vertical map is a weak equivalence. It follows that f is a pointwise weak equivalence. Consider the other implication. A pointwise acyclic fibration f: X Y is a stable equivalence according to 6.3. So 5.8 implies that f is an hf-fibration. By 6.7, it remains to prove that

[figure: X(v); t(X)(v); 𝕋(X)(v); Y (v); f(v); t(Y)(v); 𝕋(Y)(v); 𝕋(f)(v)]

is a homotopy pullback square in V for all v ∈ fV. The maps f(v) and f(T ⊗v) are acyclic fibrations and V(T, −) preserves acyclic fibrations. This implies that 𝕋(f)(v) is an acyclic fibration.

6.3 Comparison with T -spectra

To proceed with the stable model structure we will compare the stabilizations of enriched functors 6.1 and spectra [8, §4]. Recall the “suspension with T ” functors − ⊗ T and Σ_{T} on Sp(V, T) from 2.13. Let Ev_{n} := Ev_{T^{n}} denote the functor evaluating a spectrum on T^{n}. If n ≥ 0, there is the commutative diagram:

Sp(V, T) \overline{^{−⊗T}} Sp(V, T)

[figure: Sp(V, T); Σ_{T}; Ev_{n}; V; Ev_{n}]

Hence, for any spectrum E, E ⊗T and Σ_{T} E differ only in their structure maps. This statement carries over to the adjoints V(T, −): Sp(V, T) Sp(V, T) and Ω_{T}: Sp(V, T) Sp(V, T). The composition T Sph ^{⊂} V ^{T ⊗−} V does not factor over the inclusion T Sph ^{⊂} V. Hence the shift functor Sh: F F does not have a compatible analog in the category of spectra. But there is the shift sh: Sp(V, T) Sp(V, T) where (sh(E))_{n}: = E_{n+1}. The nth structure map of sh(E) is e_{n+1}, and the following diagram commutes for all n ≥ 0.

[figure: F; Sh; F; ev; Sp(V, T)]

[figure: Sp(V, T); ev; sh; Sp(V, T); Ev_{n}; V; Ev_{n}]

The stabilization for spectra uses that the structure maps of a spectrum E define a natural map s(E): E sh(Ω_{T} E). Let us abbreviate the composition sh ∘ Ω_{T} by S: Sp(V, T) Sp(V, T). Then s(E)_{n}: E_{n} S(E)_{n} = Ω_{T} E_{n+1} is the adjoint of e_{n}. The stabilization S^{∞}(E) of a spectrum E is the colimit of the diagram E ^{s(E)} S(E) ^{S(s(E))} S(S(E)) · · ·. Let s^{∞}: E S^{∞}(E) be the canonical map. In [8, §4], the notation ι: Id_{Sp(V,T)} Θ is used instead of s: Id_{Sp(V,T)} S, and j: Id_{Sp(V,T)} Θ^{∞} instead of s^{∞}: Id_{Sp(V,T)} S^{∞}.

Definition 6.15. A map f: E F of spectra is a stable equivalence if S^{∞}(R ∘ f) is a pointwise weak equivalence, and a stable fibration if f is a pointwise fibration and

[figure: E; R ∘ E; s^{∞}(R∘E); S^{∞}(R ∘ E); F; f; R ∘ F; R∘f; s^{∞}(R∘F); S^{∞}(R ∘ F); S^{∞}(R∘f)]

is a homotopy pullback square in the pointwise model structure.

We summarize the results [8, 4.12, 4.14, 6.5].

Theorem 6.16 (Hovey). Let V be an almost finitely generated, pointed, proper, and cellular monoidal model category. Let T be some cofibrant and V-finitely presentable object of V. Assume sequential colimits commute with pullbacks in V. Then Sp(V, T) is an almost finitely generated proper V-model category with stable equivalences as weak equivalences, stable fibrations as fibrations and cofibrations as cofibrations.

The condition that V be cellular might be weakened according to the remark after [8, 4.12]. An important input in the proof of 6.16 is the following lemma.

Lemma 6.17. If E is pointwise fibrant, then s^{∞}(E) is a stable equivalence with a stably fibrant codomain.

The natural maps ev(t^{∞}(X)) and s^{∞}(ev(X)) do not coincide. Compatibility of the two stabilization processes is therefore an issue, see 6.18 and 6.19 below.

Lemma 6.18. If v is an object of fV, let v^{∗}: F F be the functor where X X ∘ (v ⊗ −) and let γ_{v,n} be short for the coherence isomorphism

T ⊗ (v ⊗ T^{n}) ^{α_{T,V,T^{n}}} (T ⊗ v) ⊗ T^{n} ^{σ_{T,v}⊗T^{n}} (v ⊗ T) ⊗ T^{n} ^{α_{v,T,T^{n}}^{−1}} v ⊗ T^{n+1}.

Then the next diagram is commutative and natural in X.

[figure: X(v ⊗ T^{n}); t(X)(v⊗T^{n}); V(T, X(T ⊗ (v ⊗ T^{n}))) = 𝕋(X)(v ⊗ T^{n}); V(T, X(v ⊗ T^{n+1})); V(T,X(γ_{v,n})); s(ev(v^{∗}(X)))_{n}]

Proof. By definition, t(X)(v ⊗ T^{n}) is the adjoint of

T ^{η_{v⊗T^{n}} (T)} V(v ⊗ T^{n}, T ⊗ (v ⊗ T^{n})) ^{hom^{X}} V(X(v ⊗ T^{n}), X(T ⊗ (v ⊗ T^{n}))).

Likewise, the map s(ev(v^{∗}(X)))_{n} is the adjoint of

T ^{η_{T^{n}} (T)} V(T^{n}, T^{n+1}) ^{hom_{T^{n},T^{n+1}}^{v^{∗}(X)}} V(X(v ⊗ T^{n}), X(v ⊗ T^{n+1})).

Note that hom_{T^{n},T^{n+1}}^{v^{∗}(X)} = hom_{T^{n},T^{n+1}}^{X∘(v⊗−)} = hom_{v⊗T^{n},v⊗T^{n+1}}^{X} ∘ hom_{T^{n},T^{n+1}}^{v⊗−}. The claim follows from A.2.

Corollary 6.19. For all X ∈ Ob F and v ∈ Ob fV, there exists an isomorphism γ: 𝕋^{∞}(X)(v ⊗ T^{n}) S^{∞}(ev(v^{∗}(X)))_{n} and the diagram below is commutative and natural in X.

[figure: X(v ⊗ T^{n}); t^{∞}(X)(v⊗T^{n}); 𝕋^{∞}(X)(v ⊗ T^{n}); S^{∞}(ev(v^{∗}(X)))_{n}; γ; s^{∞}(ev(v^{∗}(X)))_{n}] Proof. The maps γ and s^{∞}(ev(v^{∗}(X)))_{n} are the vertical sequential compositions in the following commutative diagram.

Documenta Mathematica 8 (2003) 409–488

X(v ⊗ T^{n})  \overline{^{id_{X(v⊗T^{n})}}} X(v ⊗ T^{n})

[figure: 𝕋(X)(v ⊗ T^{n}); t(X)(v⊗T^{n}); V(T,X(γ^{1})); V(T, X(v ⊗ T^{n+1})); s(ev(v^{∗}(X)))_{n}]

[figure: 𝕋(𝕋(X))(v ⊗ T^{n}); 𝕋(t(X))(v⊗T^{n}); V(T,V(T,X(γ^{2}))); V(T, V(T, X(v ⊗ T^{n+2}))); V(T,s(ev(v^{∗}(X)))_{n+1})]

[figure: ⋮; ⋮]

Here γ^{1} is the map γ_{v,n} in 6.18. Note that the second square above is V(T, −) applied to the diagram

[figure: X(T ⊗ (v ⊗ T^{n})); t(X)(t⊗(v⊗T^{n})); 𝕋(X)(T ⊗ (v ⊗ T^{n})); X(v ⊗ T^{n+1}); V(T,X(γ_{1})); s(ev(v^{∗}(X)))_{n+1}; V(T, X(v ⊗ T^{n+1})).; V(T,X(γ_{2}))]

Let γ^{2} be defined as T ⊗(T ⊗(v ⊗T^{n})) ^{T ⊗γ_{v,n}} T ⊗(v ⊗T^{n+1}) ^{γ_{v,n+1}} v ⊗T^{n+1}. Then the lower square commutes because of naturality and 6.18. Likewise one constructs γ^{n} inductively, and puts γ to be colim_{n}γ^{n}. The result follows.

If v ∈ Ob fV, the composition 𝕀_{∗}X ∘ Φ ∘ (v ⊗ −) = ℏ(X) ∘ (v ⊗ −) determines a T -spectrum \overline{v^{∗}X} for every V-functor X. The nth term of \overline{v^{∗}X} is 𝕀_{∗}X(Φ(v⊗T^{n})) and the structure map \overline{v^{∗}X}_{n} ⊗ T \overline{v^{∗}X}_{n+1} is the composition

𝕀_{∗}X(Φ(v ⊗ T^{n})) ⊗ T ^{sw_{T}^{𝕀∗X}} 𝕀_{∗}X(Φ(v ⊗ T^{n}) ⊗ T) ^{𝕀_{∗}X(θ_{v⊗T^{n}})} 𝕀_{∗}X(Φ(v ⊗ T^{n+1}))

up to an associativity isomorphism. This construction is functorial and commutes with colimits and the closed V-module structures.

Lemma 6.20. A map f: X Y in F is a stable equivalence if and only if \overline{v^{∗}f}: \overline{v^{∗}X} \overline{v^{∗}Y} is a stable equivalence of T -spectra for all v ∈ Ob fV.

Proof. Let f: X Y be a stable equivalence, and pick v ∈ Ob fV. The map 𝕋′^{∞}(ℏ(R∘f))(v⊗T^{n}) is a weak equivalence in V by definition. The isomorphism mentioned in 6.19 implies that this map is isomorphic to S^{∞}(v^{∗}(R∘f)′)_{n}. Thus \overline{v^{∗}(R ∘ f)} is a stable equivalence of T -spectra. Since this is a pointwise fibrant replacement of \overline{v^{∗}f} in Sp(V, T), it follows that \overline{v^{∗}f} is a stable equivalence. The converse holds by running the argument backwards.

If additional conditions are satisfied, the characterization of stable equivalences can be improved in that ℏ becomes redundant. Note that the last characterization uses the axiom f4.

Corollary 6.21. A map f: X Y of pointwise fibrant homotopy functors is a stable equivalence if and only if ev(v^{∗}(f)) is a stable equivalence of spectra for every object v of fV.

Proof. By 6.4, f is a stable equivalence if and only if 𝕋^{∞}(f)(w) is a weak equivalence in V for every w. The proof proceeds as in 6.20.

Corollary 6.22. A map f of cofibrant functors is a stable equivalence if and only if ev(v^{∗}(f^{h})) is a stable equivalence of spectra for all v ∈ Ob fV.

Proof. Consider the diagram

[figure: X; ∼hf; RX^{h}; Y; f; ∼hf; RY^{h}.; Rf^{h}]

¿From 5.17 – which uses f4 – and 6.3 we have that f is a stable equivalence if and only if Rf^{h} is a stable equivalence. Corollary 6.21 shows that f is a stable equivalence if and only if ev(v^{∗}(Rf^{h})) is a stable equivalence in Sp(V, T) for all v ∈ Ob fV. Since ev(v^{∗}(Rf^{h})) = R ev(v^{∗}(f^{h})) and pointwise weak equivalences of spectra are stable equivalences, it follows that f is a stable equivalence if and only if ev(v^{∗}(f^{h})) is a stable equivalence of spectra.

6.4 The generating stable equivalences

Recall the stable model structure on Sp(V, T) from 6.16. In this structure, (Σ_{T} , Ω_{T}) is a Quillen equivalence by [8, 3.9]. In general, it is not clear whether (− ⊗ T, V(T, −)) – which is more natural to consider when viewing spectra as V-functors – is a Quillen equivalence. As explained in [8, 10.3], this holds if T is symmetric, which roughly means that the cyclic permutation on T ⊗ T ⊗ T is homotopic to the identity. We formulate a working hypothesis. Hypothesis: The adjoint functor pair (−⊗T, V(T, −)) is a Quillen equivalence for the model category Sp(V, T) described in 6.16.

Lemma 6.23. The maps in D are stable equivalences.

Proof. A map d_{v}□i in D is a cofibration of cofibrant functors, and by 6.22 a stable equivalence if and only if ev(w^{∗}(d_{v}□i)^{h}) = ev(w^{∗}(d_{v})^{h})□i is a stable equivalence of spectra for all w ∈ Ob fV. We claim the latter holds if f := ev(w^{∗}(d_{v})^{h}) is a stable equivalence. Factor f as a cofibration g followed by a pointwise acyclic fibration p. The stable model structure on spectra is a V-model structure by 6.16, hence f□i factors as a cobase change of the stable acyclic cofibration g□i, followed by the map p□i. Since V is assumed to be strongly monoidal, p□i is a pointwise weak equivalence. It remains to prove that d_{v} is a stable equivalence. This condition is equivalent to τ_{v} being a stable equivalence. The latter factors by definition as

V(T ⊗ v, −) ⊗ T ^{≅} V(T, V(v, −)) ⊗ T ^{ϵ_{T} V (v,−)} V(v, −)

where ϵ_{T} is the counit (− ⊗ T) ∘ V(T, −) Id_{F}. We are reduced to prove that ϵ_{T} V (v, −) is a stable equivalence. By 6.22, this map is a stable equivalence if and only if ev(w^{∗}(ϵ_{T} V(v, −))^{h}) is a stable equivalence of spectra for every w. Since ev commutes with the action and coaction of V, ev(w^{∗}(ϵ_{T} V(v, −))^{h}) coincides with ϵ_{T} ev(w^{∗}(V(v, −))^{h}). Let q: Q Id_{Sp(V,T)} be a cofibrant replacement functor in the category of spectra, so q(E)_{n} is an acyclic fibration in V for every spectrum E and n ≥ 0. Consider the following diagram, where the notation is simplified.

(QV(T, ev(V(v, R(w ⊗ −))))) ⊗ T \overline{^{q⊗T}} V(T, ev(V(v, R(w ⊗ −)))) ⊗ T

[figure: Qev(V(v, R(w ⊗ −))); Qϵ_{T}; q; ev(V(v, R(w ⊗ −))); ϵ_{T}]

The composition

s^{∞} ∘ q ∘ Qϵ_{T}: (QV(T, ev(V(v, R(w ⊗ −))))) ⊗ T S^{∞}(ev(V(v, R(w ⊗ −))))

is a stable equivalence of spectra by the hypothesis. The target of s^{∞} ∘ q ∘ Qϵ_{T} is stably fibrant, and its domain is cofibrant. Recall that V(T, −) commutes with filtered colimits. Up to an isomorphism, the stable weak equivalence

s^{∞} ∘ q: QV(T, ev(V(v, R(w ⊗ −)))) S^{∞}(V(T, ev(V(v, R(w ⊗ −)))))

is an adjoint of s^{∞} ∘ q ∘ Qϵ_{T}. Thus q ∘ Qϵ_{T} is a stable equivalence. The map q ⊗ T is a pointwise weak equivalence since V is strongly monoidal. Hence ϵ_{T} is a stable equivalence of spectra. This ends the proof.

To consider maps in D-cell, we need to record a property of the stable model structure of spectra.

Lemma 6.24. Let f: E F be a stable equivalence of spectra such that f_{n} is a retract of a map in Cof(V) ⊗ V-cell for every n ≥ 0. Then any cobase change of f is a stable equivalence.

Proof. Let g: E  G be a map of spectra. Factor g as i: E  T followed

by p: T ^{∼pt} [figure] G, and consider the diagram:

[figure: E; i; T; \begin{matrix}p \\ ∼pt \end{matrix}; G; F; f; F ∪_{E} T; f′; p′; F ∪_{E} G; f′′]

The stable model structure on spectra is left proper by 6.16, hence f′ is a stable equivalence. Pushouts are formed pointwise, so f′_{n} is a retract of a map belonging to Cof(V) ⊗ V-cell for every n ≥ 0. By the assumption that V is strongly left proper, the cobase change p′ of the pointwise weak equivalence p along f′ is again a pointwise weak equivalence. Hence f′′ is a stable equivalence.

Lemma 6.25. The maps in D-cell are stable equivalences.

Proof. Let f: X Y be a map in D-cell. First suppose that X is cofibrant. Then Y is automatically cofibrant. By 6.22, it suffices to prove that ev(v^{∗}f^{h}) is a stable equivalence of spectra for every v. The functors ev, v^{∗} and (−)^{h} preserve colimits, hence ev(v^{∗}f^{h}) is in ev(v^{∗}(D^{h}))-cell. Every map in ev(v^{∗}(D^{h})) is of the form considered in 6.24, so cobase changes of these are stable equivalences of spectra. The stable model structure on spectra is almost finitely generated, which implies that stable equivalences of spectra are closed under sequential compositions. This proves the lemma for maps in D-cell with cofibrant domain. For f arbitrary, we will construct a commutative diagram

[figure: X′; f′; Y′; X; ∼pt; f; Y; ∼pt]

where X′ is cofibrant and f′ is a map in P_{J} ∪ D-cell. It allows to finish the proof using the special case treated above. Without loss of generality, f is the sequential composition of X = X_{0} ^{f_{0}} X_{1} ^{f_{1}} ⋯, where f_{n} is the cobase change of a coproduct of maps in D. We construct f′ as a sequential composition. Consider a cofibrant replacement g_{0}: X′_{0} = X′ ^{∼pt} [figure] X. Assume f_{0} is the cobase change of z_{0}: sZ_{0} tZ_{0}, and let a_{0}: sZ_{0} X be the attaching map. The functor sZ_{0} is cofibrant, so a_{0} lifts to a map a′_{0}: sZ_{0} X′_{0}. Taking pushouts in the commutative diagram

[figure: tZ_{0}; z_{0}; sZ_{0}; a′_{0}; X′_{0}; tZ_{0}; id; z_{0}; sZ_{0}; Id; a_{0}; X_{0}; g_{0}]

gives a pointwise weak equivalence tZ_{0} ∪_{sZ_{0}} X′_{0} ^{∼pt} X_{1}. It factors as a map in P_{J}-cell followed by say X′_{1} ^{∼pt} [figure] X_{1}. By iterating this construction one finds pointwise acyclic fibrations g_{n}: X′_{n} ^{∼pt} [figure] X_{n} for all n ≥ 0. Taking the colimit gives f′: X′ Y := colim_{n}X′_{n}.

6.5 The main theorem

Before stating our main theorem, we summarize the list of assumptions. First, V is a weakly finitely generated monoidal sSet-model category for which the monoid axiom holds. Moreover, V is strongly monoidal as defined in 4.12, right proper and cellular. Assume that filtered colimits commute with pullbacks in V, that Δ^{1} is finitely presentable in V, and that cofibrations are monomorphisms. We require fV to satisfy

f1 Every object of fV is V-finitely presentable.

f2 The unit e is in fV, and fV is closed under the monoidal product.

f3 If tj ^{∼} sj v is a diagram in V where v ∈ Ob fV and j ∈ J′, then the pushout tj ∪_{sj} v is in fV.

f4 All objects in fV are cofibrant.

In what follows, T is a cofibrant object of fV with the property that (− ⊗ T, V(T, −)) is a Quillen equivalence in the stable model structure on Sp(V, T).

Theorem 6.26. Under the assumptions above, the classes of stable equivalences, stable fibrations and cofibrations give F = [fV, V] the structure of a weakly finitely generated model category.

Proof. The proof is analogous to the proof of 5.10, using 6.14 and 6.25. Nevertheless we give some details. The set of additional generating acyclic cofibrations is the set D, and the domains and codomains of the maps in D are finitely presentable. Lemma 6.25 then shows that relative cell complexes built from the generating acyclic cofibrations are stable equivalences. By 6.14 and 4.2, the stable acyclic fibrations are detected by the generating cofibrations. With 6.3 we get that all criteria of [7, 2.1.19] are satisfied.

We refer to the model structure in 6.26 as the stable model structure. If F is equipped with the stable model structure, we indicate this by the subscript “st”.

Lemma 6.27. The model category F_{st} is a monoidal F_{hf}-model category.

Proof. Axiom f4 implies that the homotopy functor structure is monoidal. Let d_{v}□i be a map in D, and let V(w, −) ⊗ j be a map in P_{I}. Then the pushout product map in F is isomorphic to d_{v⊗w}□(i□j), i.e. a stable equivalence.

Lemma 6.28. The stable model structure is proper.

Proof. Left properness follows since the pointwise model structure is left proper. To prove right properness, it remains by 5.13 to check that the base change of a stable equivalence of pointwise fibrant homotopy functors along an hf-fibration of pointwise fibrant homotopy functors is a stable equivalence. This follows from 6.5 since 𝕋^{∞} preserves pullbacks and pointwise fibrations of pointwise fibrant functors, and F_{pt} is right proper.

In the stable structure we have the following important result, analogous to and in fact an easy a consequence of Theorem 4.11.

Theorem 6.29. If fV satisfies f0, then smashing with a cofibrant V-functor in F preserves stable equivalences.

Proof. Factor a stable equivalence f as a stable acyclic cofibration followed by a stable acyclic fibration, i.e. a pointwise acyclic fibration. By 4.11, we may assume f is a stable acyclic cofibration. The claim follows since F_{st} is monoidal.

Lemma 6.30. Suppose fV satisfies f0. Then the monoid axiom holds in F_{st}.

Proof. The domains of the generating acyclic cofibrations D′ := P_{J} ∪H ∪D for the stable model structure on F are cofibrant. Because f0 holds, 4.11 implies that every map in D′∧F is a stable equivalence. The case of a map in D′∧F-cell follows similarly as in the proof of 6.25.

7 A Quillen equivalence

In this section, we will discuss two natural choices for the domain category fV. One of the choices gives a Quillen equivalence between the stable model structure Sp(V, T)_{st} on spectra and F_{st}.

7.1 The choices

Let fV_{max} be the category of all cofibrant V-finitely presentable objects, and fV_{min} the full subcategory of fV_{max} given by the objects v for which there exists an acyclic cofibration T^{n} ^{∼} v for some n ≥ 0. In the applications, the category of V-finitely presentable objects is equivalent to a small category, hence its subcategories are valid domain categories. Axioms f1, f2 and f4 hold in both cases. If the domains and codomains of the maps in J′ are V-finite, then f3 holds. The minimal choice satisfies a property which does not hold for the maximal choice in general.

Lemma 7.1. A map f of pointwise fibrant homotopy functors in [fV_{min}, V] is a stable equivalence if and only if ev(f) is a stable equivalence of spectra.

Proof. This follows by definition of fV_{min}.

The evaluation functor is a right Quillen functor for both choices.

Lemma 7.2. Evaluation ev: F_{st} Sp(V, T)_{st} is a right Quillen functor.

Proof. Pointwise fibrations and pointwise acyclic fibrations are preserved by ev. The characterizations of stable fibrations using homotopy pullback squares can be compared using 6.19, which implies that ev preserves stable fibrations.

To deduce that ev is the right adjoint of a Quillen equivalence for the minimal choice, we prove a property of the stable model structure of spectra which is independent of the choice of fV.

7.2 The unit of the adjunction

The following lemma is a crucial observation which depends on the hypotheses on T and the stabilization functor of spectra 6.17. Recall that i_{∗} is the left adjoint of ev.

Lemma 7.3. The canonical map

V_{T Sph}(T^{n}, −) (V(T^{n}, −))^{h} = ev((i_{∗}V_{T Sph}(T^{n}, −))^{h})

is a stable equivalence of T -spectra for all n ≥ 0.

Proof. By 6.17, this follows if V_{T Sph}(T^{n}, −) S^{∞}((V(T^{n}, −))^{h}) is a stable equivalence. The canonical map V_{T Sph}(T^{n}, −) ⊗ T^{n} V_{T Sph}(T^{0}, −) consists of isomorphisms in degree n and on, so it is a stable equivalence. Note that the map V_{T Sph}(T^{0}, −) V(T^{0}, −) is the identity, and that

V(T^{0}, −) (V(T^{0}, −))^{h} is even a pointwise weak equivalence. 6.17 shows that V(T^{0}, −)^{h} S^{∞}((V(T^{0}, −))^{h}) is a stable equivalence whose codomain is a stably fibrant T -spectrum. This uses that (V(T^{0}, −))^{h} is pointwise fibrant: T^{0} is cofibrant and the input is fibrant. Hence the composition

V_{T Sph}(T^{n}, −) ⊗ T^{n}  V_{T Sph}(T^{0}, −)  V(T^{0}, −)^{h}  S^{∞}(V(T^{0}, −)^{h})

is a stable equivalence from the n-fold T -suspension of a cofibrant T -spectrum to a stably fibrant T -spectrum. The functor − ⊗ T is assumed to be a Quillen equivalence. Thus its adjoint V_{T Sph}(T^{n}, −) V(T^{n}, S^{∞}((V(T^{0}, −))^{h})) is a stable equivalence. Since T is V-finite, the latter T -spectrum is isomorphic to S^{∞}((V(T^{n}, −))^{h}). It is straightforward to check that these observations imply the claim.

Corollary 7.4. The canonical map c_{E}: E ev((i_{∗}E)^{h}) is a stable equivalence for every cofibrant T -spectrum E.

Proof. For any T -spectrum E and A ∈ Ob V, the map c_{E⊗A} is isomorphic to c_{E} ⊗ A. Tensoring with the domains and codomains of the generating cofibrations preserves stable equivalences of spectra since the analogues statement holds for V. The cofibrant T -spectra are precisely the retracts of Sph(I)-cell complexes, where Sph(I) denotes the set of generating cofibrations {V_{T Sph}(T^{n}, −) ⊗ i}_{n≥0,i∈I} from [8, 1.8]. So it suffices to consider Sph(I)-cell complexes. Recall that ev, i_{∗} and (−)^{h} preserve colimits, and stable equivalences of T -spectra are closed under sequential compositions. This allows to use transfinite induction. The induction step follows from the diagram:

[figure: V_{T Sph}(T^{n}, −) ⊗ ti; V_{T Sph}(T^{n}, −) ⊗ si; E]

[figure: V(T^{n}, −)^{h} ⊗ ti; ∼; f; V(T^{n}, −)^{h} ⊗ si; ∼; ev((i_{∗}E)^{h}); ∼]

The right hand vertical map is a stable equivalence by the induction hypothesis, and likewise for the other vertical maps by 7.3 and the argument given above. Note that f is not necessarily a cofibration of spectra, but it is pointwise in Cof(V) ⊗ V. Finally, strong left properness of V implies that the map induced on the pushouts of the rows in the diagram is a stable equivalence.

Corollary 7.5. The functor ev: [fV_{min}, V] Sp(V, T) is the right adjoint in a Quillen equivalence.

Proof. Use 7.1 and 7.4.

Let V be the usual model category sSet_{∗} = S of pointed simplicial sets, and let T be the circle S^{1} = Δ^{1}/∂Δ^{1}. The n-sphere S^{n} is the n-fold smash product of S^{1}. Then fS_{max} is the full subcategory given by the finitely presentable pointed simplicial sets, and fS_{min} is the full subcategory of pointed simplicial sets K for which there exists an acyclic cofibration S^{n} ^{∼} K for some n ≥ 0. By [11] implies that the canonical functor [fS_{max}, S] [fS_{min}, S] is the right adjoint in a Quillen equivalence of stable model categories. This uses that all pointed simplicial sets are generated by spheres. In general, one needs to distinguish between using fV_{min} and fV_{max}. For example, if V = S and T is the coproduct S^{0} ∨S^{0}, we claim the corresponding stable model categories are different. In this case, fS_{max} is as above, while fS_{min} is the full subcategory of finitely presentable pointed simplicial sets which are weakly equivalent to a discrete pointed simplicial set with 1+2^{n} points for some n ≥ 0. To show that the resulting stable model structures are not Quillen equivalent via the restriction functor, we will describe a map f: X Y of S^{0}∨S^{0}-stably fibrant functors in SF := [fS_{max}, S] that is not a weak equivalence, although f(K) is a weak equivalence for every discrete pointed simplicial set. Let Y be the constant functor with value ∗. Let X be the stably fibrant replacement of X′, which maps K to the connected component of K containing the basepoint. Since X′(K ∧ Δ_{+}^{n}) = X′(K) ∧ Δ_{+}^{n}, X′ is enriched over S. Clearly, X′ is a homotopy functor, so using an enriched fibrant replacement functor R: S S, the S^{0} ∨ S^{0}-stably fibrant replacement X of X′ maps K to

X(K) = colim_{n}S((S^{0} ∨ S^{0})^{n}, RX′((S^{0} ∨ S^{0})^{n} ∧ K)).

If K is discrete, X(K) = ∗, hence the map X ∗ is a weak equivalence in [fS_{min}, S]. However, X(S^{1}) is weakly equivalent to a countable product of a countable coproduct of S^{1} with itself, and hence not contractible.

8 Algebraic structure

This section recalls the important algebraic structures which F_{st} supports if the monoid axiom holds. Recall that F_{st} satisfies the monoid axiom if fV satisfies f0, cp. 6.30. Fix a V which satisfies the conditions listed in the beginning of Section 5, and a small full sub-V-category 𝕀: fV ^{⊂} V which satisfies the axioms f1–f4.

8.1 Recollections

In a symmetric monoidal category like (F, ∧, 𝕀) there are notions of algebras and modules over the algebras. Recall that an 𝕀-algebra is a monoid in (F, ∧, 𝕀) or just an F-category with only one object. If A is an 𝕀-algebra, a (left) A-module M is an object in F with an appropriate action of A. It can alternatively be described as an F-functor from A to F. The category of A-modules mod_{A} is then an F-category. There is a smash product ∧_{A}: mod_{A^{op}} ∧ mod_{A} F. If M ∈ mod_{A^{op}} and N ∈ mod_{A}, then M ∧_{A} N is the coequalizer of M ∧ A ∧ N [figure] M ∧ N. Likewise, for Amodules M and N the function object mod_{A}(M, N) in F is the equalizer of F(M, N) [figure] F(A ∧ M, N). If k is a commutative 𝕀-algebra, recall that mod_{k} ≅ mod_{k^{op}} and mod_{k} is a closed symmetric monoidal category under ∧_{k} with internal morphism object mod_{k}(M, N). A k-algebra is a monoid in (mod_{k}, ∧_{k}, k) or a mod_{k}-category with one object. With this notation, notice that mod_{𝕀} is F.

8.2 The model structures

Definition 8.1. If A is an 𝕀-algebra and k is a commutative 𝕀-algebra, then a map in mod_{A} or alg_{k} is called a weak equivalence (resp. fibration) if it is so when considered in F_{st}. Cofibrations are defined by the left lifting property.

Remark 8.2. Note that we chose the stable model structure as our basis. This is fixed in the following (at least on the top level), so the missing prefix “stable” from fibrations and weak equivalences should not be a source of confusion.

The next result is due to Schwede and Shipley [15, 4.1].

Theorem 8.3. Suppose that F_{st} satisfies the monoid axiom. With the structures described above, the following is true.

• Let A ∈ F be an 𝕀-algebra. Then the category mod_{A} of (left) A-modules is a cofibrantly generated model category.

• Let k ∈ F be a commutative 𝕀-algebra. Then the category of k-modules is a cofibrantly generated monoidal model category satisfying the monoid axiom.

• Let k ∈ F be a commutative 𝕀-algebra. Then the category alg_{k} of kalgebras is a cofibrantly generated model category.

Note that we did not state the hypothesis that all objects in F are small. Since V is weakly finitely generated, F_{st} is so too. The smallness of the domains and codomains of the generating cofibrations and generating acyclic cofibrations in F_{st} carries over to the relevant smallness conditions needed to prove the theorem. See also [15, 2.4].

Lemma 8.4. Suppose that fV satisfies f0. Let A be an 𝕀-algebra. Then for any cofibrant A-module N, the functor − ∧_{A} N takes weak equivalences in mod_{A^{op}} to weak equivalences in F_{st}.

Proof. Given a weak equivalence in mod_{A}, factor it as an acyclic cofibration followed by an acyclic fibration. The only trouble is with the acyclic fibration, but this is a pointwise acyclic fibration, and the argument can be phrased in the analogous theory for A-modules built on the pointwise structure F_{pt}. In this case, the generating cofibrations in mod_{A} are of the form A∧S ^{A∧i} A∧T where S ^{i} T is a generating cofibration in F_{pt}. The argument of 4.11 goes through verbatim: smashing commutes with colimits and A ∧_{A} S ≅ S.

Lemma 8.4 and [15, 4.3] imply:

Corollary 8.5. Suppose that fV satisfies the axiom f0. Let f: A ^{∼} B be a weak equivalence of 𝕀-algebras. Then extension and restriction of scalars define the Quillen equivalence

mod_{A} [figure: B∧_{A}−; f^{∗}] mod_{B}.

If A and B are commutative, there is the Quillen equivalence

alg_{A} [figure: B∧_{A}−; f^{∗}] alg_{B}.

9 Equivariant Stable Homotopy Theory

Let S be the category of pointed simplicial sets, or spaces for short. The finitely presentable spaces are the ones with only finitely many non-degenerate simplices, thus we may call these finite. A simplicial functor in the sense of [11, 4.5] is an S-functor from the category of finite spaces to the category of all spaces. In [11], Lydakis showed how simplicial functors give rise to a monoidal model category which is Quillen equivalent to the model category of spectra. Thus simplicial functors model the stable homotopy category. The purpose of this section is to use the machinery developed in the main part of the paper to give a functor model for the equivariant stable homotopy category. For technical reasons we will only consider finite groups. Fix a finite group G with multiplication μ: G × G G.

9.1 Equivariant spaces

The category GS of G-spaces consists of pointed simplicial sets with a basepoint preserving left G-action. Note that G_{+} is a S-category with only one object and composition G_{+} ∧ G_{+} G_{+} induced by μ. One can identify GS with the category [G_{+}, S] of S-functors K : G_{+} S. We will often write (uK, a_{K}) for K to stress the underlying space uK ∈ Ob S, i.e. the value of K at the single object, and the left G-action a_{K}: G_{+} ∧ uK uK. Note that a_{K} is adjoint to hom^{K}: G_{+} S(uK, uK), where g g: uK uK. According to 2.4, GS is a closed S-module. The functor u: GS S has a left S-adjoint G_{+} ∧ −: S GS by 2.5. Consider the G-space G_{+} ∧ K: G_{+} S. Its underlying space is G_{+} ∧ K, with left G-action

G_{+} ∧ (G_{+} ∧ K) ^{≅} (G × G)_{+} ∧ K ^{μ_{+}∧K} G_{+} ∧ K.

Similarly, the right S-adjoint of u is given by K ×_{g∈G}K_{g}, the G-fold product of K where h ∈ G sends K_{g} to K_{hg} via the identity. Let (−)^{∧G}: S GS be the functor whose value on K is the G-fold smash product K^{∧G} = ∧_{g∈G}K_{g} of K, where G acts by permuting the factors as above. Another functor we consider is ct: S GS. The G-space ctK is constant, i.e. the underlying space is K and hom^{ctK}: G_{+} S(K, K) sends g to either the identity map or the trivial map. Let ΔG: G G × G be the diagonal map. The smash product K ∧ L of two G-spaces K, L : G_{+} S is given by the composition

G_{+} ^{ΔG_{+}} (G × G)_{+} ≅ G_{+} ∧ G_{+} ^{K∧L} S ∧ S ^{∧} S.

The right hand smash product uses [2, 6.2.9]. In other words, the smash product of G-spaces is defined on the underlying spaces and G acts diagonally. For this reason we denote the smash product of G-spaces by ∧. If G is commutative, another closed symmetric monoidal product of G-spaces exists by 2.6.

Proposition 9.1. The category (GS, ∧, ctS^{0}) is closed symmetric monoidal. The functors u, ct and (−)^{∧G} are strict symmetric monoidal, and G_{+} ∧ − is lax symmetric monoidal.

Lemma 9.2. Let K: G_{+} S be a G-space. The following are equivalent.

1. K is GS-finitely presentable.

2. K is finitely presentable.

3. uK is finite.

Proof. Let Fix(G, −) be the S-functor that maps a G-space (uK, a_{K}) to the subspace Fix(G, K) = \{x ∈ uK | a_{K} \overline{(g, x)} = x for all g ∈ G\} fixed under the action of G. Equivalently, Fix(G, K) is lim(K: G_{+} S). Note that Fix(G, −) commutes with filtered colimits since G_{+} is a finite index category. Fix(G, −) is the right S-adjoint of ct: S GS. In particular, the G-space ctS^{0} is finitely presentable. This proves the implication 1 ⇒ 2. Likewise, the right adjoint of u commutes with filtered colimits, thus 2 ⇒ 3. It remains to prove 3 ⇒ 1. Let D: I GS be a functor where I is filtered, and consider the canonical map f_{K}: colim_{I}GS(K, D) GS(K, colim_{I}D). Since colimits in GS are formed on underlying spaces, u(f_{K}) is the canonical map colim_{I}S(uK, u ∘ D) S(uK, colim_{I}u ∘ D). If uK is finite, u(f_{K}) is an isomorphism, which implies that f_{K} is an isomorphism since the G-action on the domain coincides with the G-action on the codomain. We are done.

The full subcategory of finitely presentable G-spaces is equivalent to a small category, which can be chosen to be closed under the smash product of 9.1. Often we will refer to finite G-spaces instead of finitely presentable ones.

Lemma 9.3. Any G-space is the filtered colimit of its finite sub-G-spaces.

9.2 Unstable equivariant homotopy theory

Theorems 4.2 and 4.4 give GS the coarse model structure, with weak equivalences and fibrations defined on underlying spaces. A cofibration is an injective map f: K L where G acts freely on the complement of f(K) in L. Hence, the cofibrant G-spaces are the G-spaces with a free G-action away from the basepoint. In the following, we will consider another model structure on GS. If H is a subgroup of G, let G/H_{+} be the pointed G-space with action g·g′H := (gg′)H. Consider the S-functor G/H_{+}∧: S_{+} GS_{+}, K G/H_{+} ∧ K, with trivial action on K. Its right S-adjoint is Fix(H, −): GS_{+} S_{+} which maps L to the space of fixed points under the action of H on L. Note that Fix(H, −) coincides with S_{GS}(G/H_{+}, −). If H and H′ are two subgroups of G, there is a natural isomorphism Fix(H′, G/H_{+} ∧ K) ≅ Fix(H′, G/H_{+}) ∧ K.

Definition 9.4. A map f in GS is a G-weak equivalence if Fix(H, f) is a weak equivalence in S for every subgroup H of G. Likewise for G-fibrations.

Theorem 9.5. There is a proper monoidal model structure on GS with G-weak equivalences as weak equivalences and G-fibrations as fibrations. Cofibrations are the injective maps. One can choose generating acyclic cofibrations and generating cofibrations with finitely presentable domains and codomains.

Proof. This result is well-known. A proof is included for completeness. To prove the existence of the model structure, we will apply [7, 2.1.19]. Let

I_{G} := {G/H_{+} ∧ (∂Δ^{n} ^{⊂} Δ^{n})_{+}}_{n≥0, H subgroup of G}

and

J_{G} := {G/H_{+} ∧ (Λ_{i}^{n} ^{⊂} Δ^{n})_{+}}_{n≥1, 0≤i≤n, H subgroup of G}.

It is clear by adjointness that a map is a G-fibration if and only if it is in J_{G}-inj, or a G-fibration and a G-weak equivalence if and only if it is in I_{G}-inj. From 9.2, the domains and codomains of the maps in I_{G} and J_{G} are finite. The natural isomorphism Fix(H′, G/H_{+} ∧ K) ≅ Fix(H′, G/H_{+}) ∧ K for subgroups H and H′ of G, shows that maps in J_{G} are G-weak equivalences. The existence of the model structure follows, if every map in J_{G}-cell is a G-weak equivalence. Since Fix(H′, −) commutes with sequential colimits, it suffices to check that the cobase change of a map in J_{G} is a G-weak equivalence. Fix a subgroup H′ and consider the pushout diagram:

[figure: A := (G/H × Λ_{i}^{n})_{+}; K; B := (G/H × Δ^{n})_{+}; ∩; L; ∩]

The induced map Fix(H′, B)∪_{Fix(H′,A)} Fix(H′, K) Fix(H′, L) is injective. Surjectivity follows since Fix(H′, −) preserves injective maps. Thus any cobase change of a map in J_{G} is a G-weak equivalence, and the model structure exists. Any map in I_{G}-cell is clearly injective. Conversely, by considering fixed point spaces it follows that any injective map is contained in I_{G}-cell. The statement about the cofibrations follows, and also left properness. Right properness holds since Fix(H, −) commutes with pullbacks and S is right proper. The pushout product map of injective maps is again injective, so consider the pushout product map

G/H_{+} ∧ (∂Δ^{n} ^{⊂} Δ^{n})_{+})□G/H′_{+} ∧ (Λ_{i}^{m} ^{⊂} Δ^{m})_{+} ≅ (G/H × G/H′)_{+} ∧ i,

where i is a weak equivalence of spaces. Since there is an isomorphism of spaces Fix(H′′, (G/H × G/H′)_{+} ∧ i) ≅ Fix(H′′, (G/H × G/H′)_{+}) ∧ i, the pushout product map of a generating cofibration and a generating acyclic cofibration is again acyclic. Hence the model structure is monoidal. The monoid axiom then holds, since all G-spaces are cofibrant.

We will refer to the model structure in 9.5 as the fine model structure. The regular representation S^{∧G} is the G-fold smash product of S^{1} = Δ^{1}/∂Δ^{1} where G acts by permuting the factors. Its geometric realization is homeomorphic – as a G-space – to the one-point compactification of the real vector space ℝ^{G} of maps G ℝ. The G-space S^{∧G} is finite, since G is a finite group.

9.3 Stable equivariant homotopy theory

Let fGS denote the full sub-GS-category given by the finite G-spaces. It is equivalent to a small GS-category. Objects of the enriched functor category GF = [fGS, GS] will be called G-simplicial functors. If G is the trivial group, then GF is Lydakis’ category of simplicial functors [11, 4.4]. Let ∧ denote the smash product of G-simplicial functors. The unit of ∧ is the inclusion 𝕊^{G} = 𝕀: fGS ^{⊂} GS. All G-spaces are cofibrant in the fine model structure.

Definition 9.6. A map f: X Y in GF is a

• pointwise weak equivalence if f(K) is a G-weak equivalence for all finite G-spaces K,

• pointwise fibration if f(K) is a G-fibration for all finite G-spaces K,

• cofibration if f has the left lifting property with respect to all pointwise acyclic fibrations.

Theorem 9.7. The category GF, equipped with the classes described in 9.6, is a monoidal proper model category satisfying the monoid axiom. Generating cofibrations and generating acyclic cofibrations can be chosen with finitely presentable domains and codomains. Finally, smashing with a cofibrant Gsimplicial functor preserves pointwise weak equivalences.

Proof. ¿From 4.2, P_{I_{G}} = {GS(K, −) ∧ i}_{i∈I_{G}, K finite} are the generating cofibrations, and P_{J_{G}} = {GS(K, −) ∧ j}_{j∈J_{G}, K finite} are the generating acyclic cofibrations. The model structure is monoidal and satisfies the monoid axiom by 4.4. Properness holds by 4.8. The functor − ∧ X preserves pointwise weak equivalences for cofibrant X since 4.11 holds, cp. 9.3.

Let us write GF_{pt} for the pointwise model structure. To define the homotopy functor model structure on GF, let ρ: Id_{GS} R denote the enriched fibrant replacement functor from 3.3.2 applied to the set J_{G}. Denote by 𝕀_{∗}X the enriched left Kan extension of X: fGS GS along 𝕀: fGS ^{⊂} GS, and by X^{h} the composition 𝕀_{∗}X ∘ R ∘ 𝕀. Then X^{h} defines an endofunctor of GF and there is a natural transformation Id_{GF} (−)^{h}.

Definition 9.8. A map f: X Y in GF is an

• hf-equivalence if f^{h} is a pointwise weak equivalence.

• hf-fibration if f is a pointwise fibration and there is a homotopy pullback square in GS

[figure: XK; XL; Y K; f(K); Y L; f(L)]

for every G-weak equivalence K ^{∼} L of finitely presentable G-spaces.

Lemma 9.9. A map of G-simplicial functors is a pointwise acyclic fibration if and only if it is an hf-fibration and an hf-equivalence.

Proof. The definition of hf-equivalences in 5.2 uses the filtered fibrant replacement functor Φ^{J_{G}} from 3.3.3. All G-spaces are cofibrant, so the canonical map ω_{K}: Φ^{J_{G}}K RK is a G-weak equivalence of fibrant G-spaces, hence a simplicial homotopy equivalence. Recall that ℏ(X) = 𝕀_{∗}X ∘ Φ^{J_{G}}. If K is finite, the induced map 𝕀_{∗}X(ω_{K}): ℏ(X)(K) X^{h}K is a simplicial homotopy equivalence by 2.11, in particular a G-weak equivalence. It follows that f^{h} is a pointwise weak equivalence if and only if ℏ(f)(K) is a G-weak equivalence for every finite G-space K. The arguments in 5.4 show that the hf-fibrations in 9.8 allow the same characterization as general hf-fibrations, cp. 5.6. Since every G-space is cofibrant, any G-weak equivalence of finite G-spaces can be factored as an acyclic cofibration of finite G-spaces and a simplicial homotopy equivalence. The lemma follows from 5.8.

Theorem 9.10. The category GF, equipped with the classes of hf-equivalences, hf-fibrations and cofibrations, is a proper monoidal model category satisfying the monoid axiom. Smashing with a cofibrant G-simplicial functor preserves hf-equivalences. One can choose generating cofibrations and generating acyclic cofibrations with finitely presentable domains and codomains.

Proof. The model structure exists according to 5.10, is monoidal by 5.12 and proper by 5.13. A factorization argument and 9.7 imply the claim concerning cofibrant G-simplicial functors. The monoid axiom follows easily. The generating acyclic cofibrations are H_{G} ∪ P_{J_{G}}. It remains to define H_{G}. A G-weak equivalence of finite G-spaces w: K L induces an hf-equivalence GS(w, −): GS(L, −) GS(K, −). The simplicial mapping cylinder gives a cofibration c_{w} : GS(L, −) C_{w}. Let i ∈ I_{G}. Then H_{G} is the set of pushout product maps {c_{w}□i}, cp. Subsection 5.2. It is also possible, without changing the model structure, to consider for w only acyclic cofibrations of finite G-spaces.

Denote the model category in 9.10 by GF_{hf}. In this category, X R ∘ X^{h} is a fibrant replacement of X. Recall the functor S: GF GF mapping X to GS(S^{∧G}, X(S^{∧G} ∧ −)) and the natural transformation s: Id_{GF} S obtained pointwise as the adjoint of sw_{S^{∧G}}^{X}(K): XK ∧ S^{∧G} X(S^{∧G} ∧ K). This map is the adjoint of

S^{∧G} ^{η_{K}S^{∧G}} GS(K, S^{∧G} ∧ K) ^{hom_{K,S^{∧G}∧K}^{X}} GS(XK, X(S^{∧G} ∧ K)).

Let S^{∞}(X) denote the colimit of X ^{s(X)} S(X) ^{S(s(X))} S(S(X)) ⋯, and write s: Id_{GF} S^{∞} for the canonically induced natural transformation.

Definition 9.11. A map f: X Y in GF is a

• stable equivalence if S^{∞}(R ∘ f^{h}) is a pointwise weak equivalence,

• stable fibration if f is an hf-fibration and

[figure: X; s(X); S(X); X; f; s(X); S(X); S(f)]

is a homotopy pullback square in GF_{pt}.

Lemma 9.12. A map of G-simplicial functors is a stable fibration and a stable equivalence if and only if it is a pointwise acyclic fibration.

Proof. ¿From the proof of 9.9, one sees that the definition of stable equivalences in 9.11 agrees with 6.2. The result follows from 6.14.

The definition of stable fibrations leads to a set of generating stable acyclic cofibrations as in section 6.4. Recall the definition of S^{∧G}-spectra in GS.

Definition 9.13. A G-spectrum E consists of a sequence E_{0}, E_{1}, ⋯ of G-

spaces, together with structure maps E_{n}∧S^{∧G}  E_{n+1}. A map f: E  F

of G-spectra is a sequence of maps f_{n}: E_{n} F_{n} making the diagram

[figure: E_{n} ∧ S^{∧G}; E_{n+1}; F_{n} ∧ S^{∧G}; f_{n}∧S^{∧G}; F_{n+1}; f_{n+1}]

commutative for every n ≥ 0.

By 2.12, Sp(GS, S^{∧G}) is isomorphic to the enriched category [S^{∧G}Sph, GS] of G-simplicial functors from the category of S^{∧G}-spheres S^{∧G}Sph to GS. Thus 4.2 gives Sp(GS, S^{∧G}) a pointwise model structure. A fibrant replacement functor in this model structure is E R ∘ E, where R is defined in 3.3.2. Recall that the adjoints of the structure maps of a G-spectrum E can be viewed as a natural map E Ω_{G}shE, where the nth structure map of the G-spectrum Ω_{G}shE is GS(S^{∧G}, E_{n+1} GS(S^{∧G}, E_{n+2})), the S^{∧G}- loops of the n + 1-th structure map of E. Denote this natural transformation by st: Id St = Ω_{G} ∘ sh, and let st^{∞}: Id St^{∞} be the colimit of Id ^{st} St ^{St(st)} St^{2} · · ·. The stable model structure on G-spectra, which has the same cofibrations as the pointwise model structure, is defined as follows.

Definition 9.14. A map f: E F of G-spectra is a stable equivalence if St^{∞}(R ∘ f) is a pointwise weak equivalence, and a stable fibration if f is a pointwise fibration such that

E \overline{^{st}} St(E)

[figure: F; f; st; St(F); St(f)]

is a homotopy pullback square in the pointwise model structure.

¿From now on, we consider G-spectra with the stable model structure. Note that geometric realization induces a functor from Sp(GS, S^{∧G}) to the category of G-prespectra, cf. [3, 3.2]. It is plausible that this functor induces an equivalence of homotopy categories. Hence the homotopy category of Sp(GS, S^{∧G}) is the G-equivariant stable homotopy category. Let us illustrate this by introducing spectra which are indexed on more general representations. A G-representation is a finite-dimensional euclidean vector space on which G acts via linear isometries. A G-representation V is irreducible if zero and V are the only sub-G-representations. Let Irr_{G} = {W_{1}, …, W_{r}} be a complete set of pairwise non-isomorphic irreducible G-representations. Every G-representation is isomorphic to a direct sum of representations in Irr_{G}. That is, given a Grepresentation V, there exist unique natural numbers (n_{1}, …, n_{r}) such that V is isomorphic to W_{1}^{⊕n_{1}} ⊕ ⋯ ⊕ W_{r}^{⊕n_{r}}.

If V is a G-representation, let S_{top}^{V} denote the one-point compactification of V, with ∞ as the (G-fixed) basepoint. This is a finite G-CW -complex. One has S_{top}^{V ⊕W} ≅ S_{top}^{V} ∧ S_{top}^{W}. Furthermore, one can choose a finite G-space S^{V} such that the geometric realization |S^{V} | is homeomorphic to S_{top}^{V}. Let Rep be the GS-category with objects smash products

S^{n_{1},...,n_{r}} := S^{W_{1}} ∧ ⋯ ∧ S^{W_{1}} ∧ S^{W_{2}} ∧ ⋯ ∧ S^{W_{r}}

and morphisms GS_{Rep}(S^{n_{1},...,n_{r}}, S^{n_{1}+k_{1},...,n_{r}+k_{r}}) := S^{k_{1},...,k_{r}}. Hence Rep contains essentially all G-representations. The G-sphere S^{∧G} does not reside in this category, but it contains a G-space S̃^{∧G} with homeomorphic realization. Consequently, the stable model categories [S^{∧G}, GS] and [S̃^{∧G}, GS] are Quillen equivalent, and there are inclusions S̃^{∧G} ^{⊂} ^{j} Rep ^{⊂} fGS inducing functors

GF = [fGS, GS] [Rep, GS] ^{j^{∗}} [S̃^{∧G}, GS].

Note that j is a full inclusion. It is straightforward to define the stabilization St_{Rep}^{∞}X of a pointwise fibrant GS-functor X: Rep GS. Since every G-representation is a direct summand of a direct sum of copies of the regular representation ℝ^{G}, j^{∗}(St_{Rep}^{∞}X) is equivalent to St^{∞}j^{∗}X. This shows that the stable model structure on [Rep, GS] is Quillen equivalent to the stable model structure on G-spectra. In particular, smashing with S^{V} is a Quillen equivalence of G-spectra for every Grepresentation V. Let us turn to the last ingredient needed in the proof of 9.16.

Proposition 9.15. The functor − ∧ S^{∧G}: Sp(GS, S^{∧G}) Sp(GS, S^{∧G}) is a Quillen equivalence.

Proof. We will show that S^{∧G} is G-weakly equivalent to a symmetric G-space. The result follows then by [8, 10.3]. A G-space K is symmetric if there exists a map H such that the following diagram commutes, where cyc: K^{∧3} K^{∧3} is the cyclic permutation map.

[figure: K^{∧3} ∧ ctS^{0}; K^{∧3}∧cti_{0}; K^{∧3} ∧ ctΔ_{+}^{1}; K^{∧3}∧cti_{1}; K^{∧3} ∧ ctS^{0}]

[figure: K^{∧3} ∧ ctS^{0}; H; cyc∧ctS^{0}; id_{K^{∧3}} ∧ctS^{0}]

Consider first the trivial group. The cyclic permutation map on the geometric realization |S^{1}|^{∧3} is homotopic to the identity. Thus the singular complex K := sing|S^{1}| is a symmetric space and there is a weak equivalence S^{1} K. We claim the induced map S^{∧G} K^{∧G} of G-fold smash products is a G-weak equivalence. To see this, choose a subgroup H of G and write the underlying set of G as the union of the cosets gH. Fix a space L, and an element x = x_{g_{1}} ∧ ⋯ ∧ x_{g_{n}} of L^{∧G} where n is the order of G and x_{g_{k}} ∈ L for every k. Note that x is invariant under the action of all h ∈ H if and only if x_{g_{k}} = x_{g_{j}} for all g_{k}, g_{j} in the same coset of H. Thus Fix(H, L^{∧G}) is – up to a natural isomorphism – the G/H-fold smash product of L. In particular, we get an expression of the “diagonal” d_{L}: ctL L^{∧G} as the adjoint of the isomorphism L Fix(G, L^{∧G}). We will use essentially three maps to define a homotopy from the identity map to the cyclic permutation map on K^{∧G} ∧ K^{∧G} ∧ K^{∧G}: (1) the natural isomorphism, cp. 9.1, of G-spaces f_{K,L}: K^{∧G} ∧ L^{∧G} (K ∧ L)^{∧G} which rearranges the factors, (2) the diagonal d = d_{Δ_{+}^{1}}: ctΔ_{+}^{1} (Δ_{+}^{1})^{∧G}, and (3) the homotopy F : K ∧K ∧K ∧Δ_{+}^{1} K ∧K ∧K from the cyclic permutation map to the identity map. Consider now the composition

K^{∧G} ∧ K^{∧G} ∧ K^{∧G} ∧ ctΔ_{+}^{1}

K^{∧G} ∧ K^{∧G} ∧ K^{∧G} ∧ (Δ_{+}^{1})^{∧G} ^{id∧d}

(K ∧ K ∧ K)^{∧G} ^{≅}

(K ∧ K ∧ K ∧ Δ_{+}^{1})^{∧G} ^{F^{∧G}}

K^{∧G} ∧ K^{∧G} ∧ K^{∧G}. ^{≅}

This is the homotopy which shows that K^{∧G} is a symmetric G-space.

Theorem 9.16. The category GF and the classes of stable equivalences, stable fibrations and cofibrations, is a proper monoidal model category satisfying the monoid axiom. One can choose generating cofibrations and generating acyclic cofibrations with finite domains and codomains. Smashing with a cofibrant Gsimplicial functor preserves stable equivalences.

Proof. The results above allow us to apply 6.26, 6.27, 6.28, 6.29 and 6.30.

Let GF_{st} refer to the stable model category in 9.16. We end this section by comparing GF_{st} with the stable model category of G-spectra. By Section 2.5, S^{∧G}Sph is a sub-GS-category of fGS. Let i_{∗} be the enriched left Kan extension along the corresponding inclusion i. It is left adjoint to pre-composition with i, which we denote by ev. The next result follows from 7.2.

Lemma 9.17. ev: GF_{st} Sp(GS, S^{∧G}) is a right Quillen functor.

Lemma 7.4 implies that the unit of the adjunction has the following property.

Lemma 9.18. The canonical map E ev(i_{∗}E)^{h} is a stable equivalence of G-spectra for every cofibrant G-spectrum E.

If G is the trivial group, Lydakis proved that a map of homotopy functors is a stable equivalence if and only its evaluation is a stable equivalence of spectra. The proof uses the Blakers-Massey theorem. We will extend this result to any finite group using Spanier-Whitehead duality, cp. [14, 17.6].

Proposition 9.19. Let K and L be finitely presentable G-spaces. The canonical map

ev(GS(K, R(−)) ∧ L) ev(GS(K, R(L ∧ −)))

of G-spectra is a stable equivalence.

Proof. Let E ⋆F denote the (closed) symmetric monoidal product in the equivariant stable homotopy category SH(G), with unit 𝕊, and let Hom(E, −) denote the right adjoint of − ⋆ E. A G-spectrum D is dualizable if the canonical map Hom(D, 𝕊) ⋆ D Hom(D, D) is an isomorphism in SH(G). It follows that the canonical map Hom(D, E) ⋆ F Hom(D, E ⋆ F) is an isomorphism for all E, F ∈ Ob SH(G) if D is dualizable [10, II, Section 1], Suspension G-spectra of finite G-spaces are dualizable in SH(G) [3, 2.C], [10, II, 2.7]. In particular, given finite G-spaces K and L, the canonical map

Hom(ev(−∧K), evR(−))⋆ev(−∧L) Hom(ev(−∧K), evR(−)⋆ev(−∧L))

is an isomorphism in SH(G). In this special situation, a map of G-spectra lifting this isomorphism can be given as

St^{∞}(R(GS(K, evR(−)) ∧ L GS(K, evR(L ∧ −)))).

In particular, this map is a stable equivalence. This finishes the proof, because the above is a stably fibrant replacement of the map in question.

Corollary 9.20. Let X be a G-simplicial functor and L be a finite G-space. The canonical map evX^{h} ∧ L evX^{h}(L ∧ −) is a stable equivalence of G-spectra. In particular, ev reflects stable equivalences of homotopy functors.

Proof. We sketch a proof, following the script for the trivial group [11, 11.7]. Consider the first statement. If X is a cofibrant G-simplicial functor, use 9.19 by attaching cells. If X is arbitrary, use a cofibrant replacement X^{c}. The second statement then follows from 6.21.

Corollary 9.21. The stable model structure on G-simplicial functors is Quillen equivalent to the stable model structure on G-spectra via the right Quillen functor ev.

The Quillen equivalence from 9.21 factors through a Quillen equivalence to the category of symmetric G-spectra. In fact, the work [8] of Hovey shows that G-spectra and symmetric G-spectra are Quillen equivalent via a zig-zag of Quillen equivalences, but not necessarily via the canonical forgetful functor. In this case, however, it is possible to conclude this by extending results of [9] to G-spaces. Here are the details.

Theorem 9.22. The forgetful functor from symmetric G-spectra to G-spectra is the right adjoint of a Quillen equivalence.

Proof. In both the categories of G-spectra and symmetric G-spectra, a map of fibrant objects is a weak equivalence if and only if it is a pointwise weak equivalence. Hence the forgetful functor U preserves and reflects weak equivalence of fibrant objects. Its left adjoint V preserves cofibrations and pointwise acyclic cofibrations. Further, if L resp. L^{Σ} is the set of maps of G-spectra resp. symmetric G-spectra that Hovey uses to localize the pointwise model structures, then V maps L to L^{Σ} (up to isomorphism). Hence V also preserves stable acyclic cofibrations by properties of Bousfield localization [6] and is thus a left Quillen functor. It remains to prove that the canonical map

E U(V (E)^{f})

is a stable equivalence of G-spectra for E cofibrant. Here (−)^{f} denotes a fibrant replacement in the stable model structure of symmetric G-spectra. In fact, since both U and V preserve colimits and homotopy pushouts, it suffices to prove this for E varying through the domains and codomains of the generating cofibrations. To do so, we use the functors Fix(H, −) on (symmetric) spectrum level. This means the following. If E is a G-spectrum, the sequence (Fix(H, E_{0}), Fix(H, E_{1}), …) is a F ix(H, S^{∧G}) = S^{|G/H|}-spectrum of spaces. That is, it has structure maps

Fix(H, S^{∧G}) ∧ Fix(H, E_{n}) ≅ Fix(H, S^{∧G} ∧ E_{n}) Fix(H, E_{n+1}).

The same construction works on the level of symmetric G-spectra, so we have a commutative diagram

Sp^{Σ}(GS, S^{∧G}) \overline{^{Fix(H,−)}} Sp^{Σ}(S, S^{|G/H|})

[figure: Sp(GS, S^{∧G}); U; Fix(H,−); Sp(S, S^{|G/H|}).; U_{H}]

The domains and codomains of the generating cofibrations of G-spectra are of the form Fr_{n}(K) (representable ∧K) for K in a certain set of G-spaces, and similarly for the S^{|G/H|}-spectra of spaces. Since Fix(H, −): GS S commutes with the smash product (up to natural isomorphism), we get a natural isomorphism

V_{H}(Fr_{n}(Fix(H, K))) ≅ V_{H}(Fix(H, Fr_{n}(K))) ≅ Fix(H, V (Fr_{n}(K))).

This isomorphism is compatible with the units of the adjunctions (V, U) and (V_{H}, U_{H}), so that

Fix(H, Fr_{n}(K)  UV Fr_{n}(K)) ≅ Fr_{n}Fix(H, K)  U_{H}V_{H}(Fr_{n}Fix(H, K)).

The categories of (symmetric) S^{|G/H|}-spectra of spaces are just slight variations of the categories of (symmetric) spectra of spaces, which implies that (V_{H}, U_{H}) is a Quillen equivalence. To conclude the same for (V, U), it is sufficient, by the above, to prove the following two facts.

• A map f of G-spectra is a stable equivalence if so is Fix(H, f) for every subgroup H.

• If j_{E}: E ^{∼} E^{f} is a stably fibrant replacement of the symmetric Gspectrum E, then Fix(H, j_{E}: E ^{∼} E^{f}) is a stably fibrant replacement of Fix(H, E).

Concerning the first fact: a map f of G-spectra is clearly a pointwise weak equivalence if and only if so is Fix(H, f) for every subgroup H. Since a stable equivalence of stably fibrant G-spectra is a pointwise weak equivalence, it suffices to prove that Fix(H, −) preserves stably fibrant replacements for G-spectra. For this purpose, we apply the small object argument to the following set. Let J_{pt} := {Fr_{n}(∧(G/H × (Λ_{i}^{m} ^{⊂} Δ^{m}))_{+}}_{n,H,m,i} be the set of generating pointwise acyclic cofibrations. Obtain J̃_{st} from L = {Fr_{n+1}S^{∧G} Fr_{n}(S^{0})}_{n} by applying the simplicial mapping cylinder, and let J_{st} be the set

J̃_{st}□{(G/H × (∂Δ^{m} ^{⊂} Δ^{m}))}_{H,m}

of pushout product maps. Finally, J = J_{pt} ∪ J_{st} is the set we may use for a fibrant replacement^{1}. Note first that Fix(H, Fr_{n+1}S^{∧G} Fr_{n}(S^{0})) ≅ Fr_{n+1}S^{|G/H|} Fr_{n}(S^{0}). Further, Fix(H, −) is compatible with the simplicial mapping cylinder construction, since it commutes with the smash product and with pushouts of diagrams containing a monomorphism. The latter fact was already used in the proof of 9.5. It follows that Fix(H, −) maps J to the corresponding set J_{H} in the category Sp(S, S^{|G/H|}). In particular, Fix(H, −) maps sequential compositions of cobase changes of maps in J to stable equivalences. To conclude that Fix(H, −) preserves the fibrant replacement, we have to show that it preserves stably fibrant objects. One can see this by arguing that it is in fact a right Quillen functor, whose left adjoint l_{H} is determined by the requirement that l_{H}Fr_{n}(K) = Fr_{n}(G/H_{+} ∧ K) for any n and any space K. Using this description, one can see that l_{H} preserves generating (acyclic) cofibrations, hence is a left Quillen functor. It follows that Fix(H, −) preserves the fibrant replacement. This proof translates to the category of symmetric G-spectra, which justifies the second fact in the list above. This finishes the proof.

1The set J is in fact a set of generating acyclic cofibrations.

Documenta Mathematica 8 (2003) 409–488

Since the resulting Quillen equivalence (j_{∗}, j^{∗}) between G-simplicial functors and symmetric G-spectra has nice monoidal properties according to 2.16, the closed symmetric monoidal structure induced by the smash product of Gsimplicial functors is the correct one. Given the above, comparisons of modules and algebras along the lines of [14] are possible. Note, however, that at present it is not clear how to compare commutative algebras [14, 0.9].

Corollary 9.23. The model categories of symmetric ring G-spectra and of 𝕀-algebras in GF are Quillen equivalent via the canonical adjoint pair (ι_{∗}, ι^{∗}). If R is a cofibrant 𝕀-algebra, the model categories of R-modules and of ι^{∗}Rmodules are Quillen equivalent. If Q is a cofibrant symmetric ring G-spectrum, the model categories of Q-modules and of j_{∗}Q-modules are Quillen equivalent.

Proof. First we observe that the model category of symmetric G-spectra satisfies the monoid axiom. Here are some details. A pushout diagram

[figure: X; Y; Z; f; Z ∪_{X} Y; g]

of symmetric G-spectra in which f is a pointwise cofibration is a homotopy pushout square. This, the fact that stable equivalences of symmetric G-spectra are closed under sequential colimits (see 3.5) and general arguments from [15] show that it suffices to prove the following. Let X be a symmetric G-spectrum and j: sj ^{∼} tj a generating acyclic cofibration, then X ∧ j is a stable equivalence and a pointwise cofibration. To see that the latter holds, note that the pushout product map of a pointwise cofibration and a cofibration is a pointwise cofibration, by comparing with the smash product of symmetric sequences of G-spaces as in [9, section 5.3]. Now one can use that the stable model structure on symmetric G-spectra is stable in the sense that suspension with S^{1} is a Quillen equivalence. So X ∧ j will be a stable equivalence if and only if X ∧ (tj/sj) is stably equivalent to a point. Note that tj/sj is cofibrant. Hence it suffices to prove that the smash product of a cofibrant symmetric G-spectrum and a pointwise weak equivalence is a pointwise weak equivalence. By arguments which already appeared in this proof, one can reduce to the case of the domains and codomains of the generating cofibrations (which are cofibrant). These, however, are gotten directly from symmetric sequences of G-spaces. A comparison like [9, proof of 5.3.7] of the smash products of symmetric G-spectra and these sequences concludes the proof of the monoid axiom. The hard work is done. Let 𝕊 denote the unit in the category of symmetric G-spectra, and recall that the unit in GF is the inclusion 𝕀. By 2.16, the canonical adjoint pair (j_{∗}, j^{∗}) induced by the inclusion j: S^{∧G}Sph^{Σ} ^{⊂} fGS induces an adjoint pair

ι_{∗}: alg_{𝕊} [figure] alg_{𝕀}: ι^{∗}.

Since the forgetful functors with domain alg_{𝕊} resp. alg_{𝕀} detect weak equivalences and fibrations, (ι_{∗}, ι^{∗}) is a Quillen adjunction. Moreover, since j^{∗} preserves and detects weak equivalences of fibrant G-simplicial functors, ι^{∗} detects weak equivalences of fibrant 𝕀-algebras. To conclude that (ι_{∗}, ι^{∗}) is a Quillen equivalence, it suffices to note that a cofibrant 𝕊-algebra is in particular a cofibrant symmetric G-spectrum [15, 4.1]. The other two cases are similar, modulo an application of 8.4.

It is desirable to compare the stable model category of G-simplicial functors also with the stable model category of orthogonal G-spectra [13]. As an intermediate step, we will use symmetric spectra of topological G-spaces. Let T denote the (closed symmetric monoidal) model category of pointed compactly generated topological spaces [7, 2.4.21], and let GT be the category of G-objects in T. The latter is closed symmetric monoidal by an analog of 9.1, and it is a monoidal model category by transferring the model structure from 9.5 to the topological situation (see [13, II.1.8 and II.1.22]). The (strict symmetric monoidal) Quillen equivalence | − |: S T given by geometric realization extends to a (strict symmetric monoidal) Quillen equivalence | − |: GS GT. In particular, we can regard GT as a GS-model category. As one can check using results from [13, II.1], the model structure on GT is cellular, so the stable model category of symmetric spectra in GT with respect to |S^{∧G}| exists by [8]. Further, we can apply [8, 9.3] to conclude that

| − |: Sp^{Σ}(GS, S^{∧G}) Sp^{Σ}(GT, |S^{∧G}|)

is a Quillen equivalence. By inspection, this Quillen functor is strict symmetric monoidal, and its right adjoint is lax symmetric monoidal. Thus to obtain a variant of 9.23, it suffices to note that the stable model category Sp^{Σ}(GT, |S^{∧G}|) satisfies the monoid axiom. A proof can be obtained by translating the proof of the monoid axiom for Sp^{Σ}(GS, S^{∧G}) to the topological situation. It remains to relate orthogonal G-spectra to symmetric G-spectra of topological spaces. For the definition of an orthogonal G-spectrum, which we take to be indexed on all G-representations, consider [13]. Any orthogonal G-spectrum X gives rise to a symmetric G-spectrum of topological spaces uX by neglect of structure, or rather – since both objects are simply enriched functors on certain domain GT-categories – restriction. The restriction takes place both on objects (from all G-representations to direct sums of the regular representation ℝ^{G}) and morphisms (from orthogonal groups to symmetric groups). See [14, 4.4] (for the non-equivariant case) and [13, II.4]. Viewed as a restriction, u has a left adjoint v by enriched Kan extension.

Theorem 9.24. The adjoint pair (v, u) is a Quillen equivalence. It induces Quillen equivalences between orthogonal ring G-spectra and symmetric ring Gspectra of topological spaces. If R is a cofibrant symmetric ring G-spectrum and P is a cofibrant orthogonal ring G-spectrum, (v, u) induces Quillen equivalences

mod_{R} [figure] mod_{vR} mod_{uP} [figure] mod_{P} .

Proof. The forgetful functor U_{T}: Sp^{Σ}(GT, |S^{∧G}|) Sp(GT, |S^{∧G}|) has a left adjoint V_{T} fitting into a commutative diagram

Sp(GS, S^{∧G}) \overline{^{|−|}} Sp(GT, |S^{∧G}|)

[figure: Sp^{Σ}(GS, S^{∧G}); V; |−|; Spe^{Σ}(GT, |S^{∧G}|); V_{T}]

in which the left vertical functor resp. the lower horizontal functor are Quillen equivalences by 9.22 resp. [8, 9.3]. The upper horizontal functor is a Quillen equivalence by [8, 5.7], hence the right vertical functor is a Quillen equivalence. Thus to conclude that (v, u) is a Quillen equivalence, it suffices to prove that the forgetful functor from orthogonal G-spectra to Sp(GT, |S^{∧G}|) is a Quillen equivalence. By [13, III.4.16], the forgetful functor from orthogonal G-spectra to Gprespectra as defined in [13, II.1.2] is a Quillen equivalence. The category of G-prespectra so far is indexed on all G-representations. However, as observed in [13, II.2.2 and V.1.10], one can index both orthogonal G-spectra and Gprespectra on a collection of G-representations with is both closed under direct sum and cofinal in the collection of all G-representations without changing the homotopy theory. An acceptable candidate is the collection of direct sums of the regular representation. Hence the restriction from orthogonal G-spectra to Sp(GT, |S^{∧G}|) is a Quillen equivalence. This proves the first statement. The other statements then follow as in the proof of 9.23, since the monoid axiom holds for orthogonal G-spectra [13, III.7.4].

Hence for the purpose of studying the homotopy theory of algebras and modules, the category of G-simplicial functors is as good as the category of orthogonal G-spectra for a finite group G. Another comparison functor can be obtained as in [14, 19.11] by passing from GF to GT-functors from an appropriate domain category (say, finite G-CW-complexes) to GT via geometric realization, and then restricting to orthogonal G-spectra. Up to geometric realization, this functor amounts to a neglect of structure.

A Calculations

This appendix looks into the proofs of the remaining claims in the main part of the paper. The structure map hom_{A,B}^{−⊗T}: V(A, B) V(A ⊗ T, B ⊗ T) of the V-functor − ⊗ T: V V is defined as the adjoint of the composition

V(A, B) ⊗ (A ⊗ T) α_{V(A,B),A,T}^{−1} ; ≅ (V(A, B) ⊗ A) ⊗ T ^{(ϵ_{A}B)⊗T} T ⊗ A

482 B. I. Dundas, O. Röndigs, P. A. Østvær

where α_{V(B,T),B,A}^{−1} is the associativity isomorphism. The next lemma shows that there are in general two different suspension functors for T -spectra. Lemma A.1. The following diagram commutes.

[figure: V(A, B) ⊗ (A ⊗ T); hom_{A,B}^{−⊗T} ⊗(A⊗T); V(A ⊗ T, B ⊗ T) ⊗ (A ⊗ T); (V(A, B) ⊗ A) ⊗ T; α_{V(A,B),A,T}^{−1}; (ϵ_{A}B)⊗T; B ⊗ T; ϵ_{A⊗T} (B⊗T)]

Proof. Use naturality and the triangular identity ϵ_{A⊗T} (V(A, B) ⊗ (A ⊗ T)) ∘ (η_{A⊗T} V(A, B)) ⊗ (A ⊗ T) = id_{V(A,B)⊗(A⊗T)}.

A similar statement is used to show that the stabilization of enriched functors and the stabilization of spectra can be compared. Lemma A.2. The following diagram commutes.

[figure: A; η_{B}A; V(B, A ⊗ B); V(B ⊗ T, A ⊗ (B ⊗ T)); η_{B⊗T} A; V(B⊗T,α_{A,B,T}^{−1} ); V(B ⊗ T, (A ⊗ B) ⊗ T); hom_{B,A⊗B}^{−⊗T}]

Proof. Similar to A.1, using ϵ_{B}(A ⊗ B) ∘ (η_{B}A) ⊗ B = id_{A⊗B}. Next we start the proof of: the two natural stabilization maps X T (X) described in 6.1 coincide. It is lengthy and perhaps not very illuminating. The map hom_{A,B}^{V(T,−)} : V(A, B) V(V(T, A), V(T, B)) is given as the adjoint of comp: V(A, B) ⊗ V(T, A) V(T, B) which, up to an associativity isomorphism, is adjoint to V(A, B) ⊗ (V(T, A) ⊗ T) ^{V(A,B)⊗ϵ_{T} A} V(A, B) ⊗ A ^{ϵ_{A}B} B. Lemma A.3. The following diagram commutes.

[figure: V(A, B); hom_{A,B}^{V(T,−)}; V(V(T, A), V(T, B)); V(V(T, A) ⊗ T, B); ≅ ; f; V(ϵ_{T} A,B)]

Proof. Here, f is V(η_{T} V(T, A), V(T, B)) ∘ hom_{V(T,A)⊗T,B}^{V(T,−)}. The diagram

[figure: V(A, B); hom_{A,B}^{V(T,−)}; V(V(T, A), V(T, B)); V(V(T, A) ⊗ T, B); V(ϵ_{T} A,B); hom_{V(T,A)⊗T,B}^{V(T,−)}; V(V(T, V(T, A) ⊗ T), V(T, B)); V(V(T,ϵ_{T} A),V(T,B))]

Documenta Mathematica 8 (2003) 409–488

commutes, because V(T, −) is a V-functor. Therefore, f ∘V(ϵ_{T} (A), B) coincides with V(η_{T} V(T, A), V(T, B)) ∘ V(V(T, ϵ_{T} (A)), V(T, B)) ∘ hom_{V(T,A)⊗T,B}^{V(T,−)} and the triangular identity V(T, ϵ_{T} (A))∘η_{T} V(T, A) = id_{V(T,A)} completes the proof.

Let C be a full sub-V-category closed under ⊗. If v ∈ Ob C and X: C V is a V-functor, one can consider the adjoint X(c) ⊗ v X(v ⊗ c) of the

map v ^{η_{c}v} V(c, v ⊗ c) ^{hom_{c,v⊗c}^{X}} V(X(c), X(v ⊗ c)). It defines a V-natural transformation sw_{v}^{X}: X ⊗ v X ∘ (v ⊗ −) (X swallows v). One of the maps X 𝕋(X) is defined using the map sw_{T}^{X}. Using a commutativity isomorphism, one can define a map v ⊗ X X ∘ (v ⊗ −) which will also be denoted sw_{v}^{X}. An interesting case is V(T, −): V V. Then sw_{A}^{V(T,−)}(B): A⊗

V(T, B) V(T, A ⊗ B) is the adjoint of (A ⊗ V(T, B)) ⊗ T ^{α_{A,V(T,B)}} A ⊗ (V(T, B) ⊗ T) ^{A⊗ϵ_{T} B} A ⊗ B.

Lemma A.4. Let A, B, T ∈ Ob V. The following diagram commutes.

[figure: A; η_{V(T,B)}A; V(V(T, B), A ⊗ V(T, B)); V(B, A ⊗ B); η_{B}A; hom_{B,A⊗B}^{V(T,−)}; V(V(T, B), V (T, A ⊗ B)); V(V(T,B),sw_{A}^{V(T,−)}(B))]

Proof. Using the definition of hom^{V(T,−)} and the description of sw_{A}^{V(T,−)} from above, one gets a large diagram which commutes by naturality and the triangular identity ϵ_{B}(A ⊗ B) ∘ (η_{B}A) ⊗ B = id_{A⊗B}.

Lemma A.5. Let A, B, T ∈ Ob V. The following diagram commutes.

[figure: V(A, B) ⊗ V(T, A); sw_{V(A,B)}^{V(T,−)}(A); V(T, V(A, B) ⊗ A); V(V(T, A), V(T, B)) ⊗ V(T, A); hom_{A,B}^{V(T,−)}⊗V(T,A); ϵ_{V(T,A)}V(T,B); V(T, B); V(T,ϵ_{A}B)]

Proof. This is similar to the proof of A.4; the relevant triangular identity is ϵ_{V(T,A)}(V(A, B)⊗V(T, A))∘(η_{V(T,A)}V(A, B))⊗V(T, A) = id_{V(A,B)⊗V(T,A)}.

Lemma A.6. Let A, B, T ∈ Ob V. The following diagram commutes.

A ⊗ V(T, B) \overline{^{η_{T} A⊗V(T,B)}} V(T, A ⊗ T) ⊗ V(T, B)

[figure: V(T, A ⊗ B); sw_{A}^{V(T,−)}(B); V(T, V(T, B) ⊗ (A ⊗ T)); sw_{V(T,B)}^{V(T,−)}(A⊗T)]

[figure: V(T, B ⊗ A); V(T,σ_{A,B}); V(T,(ϵ_{T} B)⊗A); V(T, (V(T, B) ⊗ T) ⊗ A); V(T,c)]

Here c is the composition of an associativity and a commutativity isomorphism.

Proof. Insert the description of sw^{V(T,−)} and use naturality, associativity and commutativity coherence, and ϵ_{T} (A ⊗ T) ∘ (η_{T} A) ⊗ T = id_{A⊗T}.

Lemma A.7. Let C be a full sub-V-category closed under ⊗, X: C V a V-functor, and f: V(T, V(v, w)) V(T ⊗v, w) the adjointness isomorphism, with T, v, w ∈ Ob C. The following diagram commutes.

V(T, V(v, w)) ⊗ (T ⊗ Xv)  \overline{^{α^{−1}}} (V(T, V(v, w)) ⊗ T) ⊗ Xv

[figure: V(T ⊗ v, w) ⊗ X(T ⊗ v); f⊗sw_{T}^{X}(v); V(v, w) ⊗ Xv; (ϵ_{T} V(v,w))⊗Xv]

[figure: V(X(T ⊗ v), Xw) ⊗ X(T ⊗ v); hom_{T ⊗v,w}^{X}⊗X(T ⊗v); V(Xv, Xw) ⊗ Xv; hom_{v,w}^{X}⊗Xv]

[figure: Xw; ϵ_{Xv}Xw; ϵ_{X(T ⊗v)}Xw]

Proof. The proof is divided into two steps. First we note that

V(T, V(v, w)) ⊗ (T ⊗ Xv)  \overline{^{α^{−1}}} (V(T, V(v, w)) ⊗ T) ⊗ Xv

[figure: V(T, V(Xv, Xw)) ⊗ (T ⊗ Xv); V(T,hom_{v,w}^{X})⊗(T ⊗Xv); V(v, w) ⊗ Xv; (ϵ_{T} V(v,w))⊗Xv]

[figure: V(T ⊗ Xv, Xw) ⊗ (T ⊗ Xv); g⊗(T ⊗Xv); V(Xv, Xw) ⊗ Xv; hom_{v,w}^{X}⊗Xv]

[figure: Xw; ϵ_{Xv}Xw; ϵ_{T ⊗Xv}Xw]

commutes, where g is the adjointness isomorphism

V(T ⊗ Xv, ϵ_{Xv}Xw) ∘ hom_{T,V(Xv,Xw)}^{−⊗Xv}.

Commutativity of the diagram follows from the definition of hom^{−⊗Xv}, naturality and the triangular identity ϵ_{T ⊗Xv}(−⊗(T ⊗Xv))∘(η_{T ⊗Xv})⊗(T ⊗Xv) = id applied to V(T, V(Xv, Xw)) ⊗ (T ⊗ Xv). In the second step, we prove that

V(T, V(v, w)) ⊗ T ⊗ Xv \overline{^{V(T,hom_{v,w}^{X})⊗T ⊗Xv}} V(T, V(Xv, Xw)) ⊗ T ⊗ Xv

[figure: V(T ⊗ v, w) ⊗ X(T ⊗ v); f⊗sw_{T}^{X}; V(T ⊗ Xv, Xw) ⊗ T ⊗ Xv; g⊗T ⊗Xv]

[figure: V(X(T ⊗ v), Xw) ⊗ X(T ⊗ v); hom_{T ⊗v,w}^{X} ⊗X(T ⊗v); ϵ_{X(T ⊗v)}Xw; Xw; ϵ_{T ⊗Xv}Xw]

Documenta Mathematica 8 (2003) 409–488

commutes. The adjoint of ϵ_{X(T ⊗v)}Xw ∘ (hom_{T ⊗v,w}^{X} ⊗ sw_{T}^{X}) coincides with the composition comp ∘ (hom_{T ⊗v,w}^{X} ⊗ hom_{v,T ⊗v}^{X}) ∘ V(T ⊗ v, w) ⊗ η_{v}T. Because X is a V-functor, this map is the same as hom_{v,w}^{X} ∘ comp ∘ V(T ⊗ v, w) ⊗ η_{v}T. Hence the diagram above commutes if and only if

V(T, V(v, w)) ⊗ T ⊗ Xv  \overline{^{V(T,hom_{v,w}^{X})⊗T ⊗Xv}} V(T, V(Xv, Xw)) ⊗ T ⊗ Xv

[figure: V(T ⊗ v, w) ⊗ V(v, T ⊗ v) ⊗ Xv; f⊗η_{v}T ⊗Xv; V(T ⊗ Xv, Xw) ⊗ T ⊗ Xv; g⊗T ⊗Xv]

[figure: V(v, w) ⊗ Xv; comp⊗Xv; Xw; ϵ_{T ⊗Xv}Xw]

[figure: V(Xv, Xw) ⊗ Xv; ϵ_{Xv}Xw; hom_{v,w}^{X}⊗Xv]

commutes. The shortest composition is hom_{v,w}^{X} ⊗ Xw ∘ (ϵ_{T} V(v, w)) ⊗ Xw, as the triangular identity ϵ_{T ⊗Xv}(−⊗(T ⊗Xv))∘(η_{T ⊗Xv})⊗(T ⊗Xv) = id_{−⊗(T ⊗Xv)} evaluated at V(T, V(Xv, Xw)) shows. Therefore it remains to prove that the map comp ∘ (f ⊗ η_{v}T) coincides with the map ϵ_{T} V(v, w). It is equivalent to switch to the adjoints (under tensoring with v), and here naturality and the triangular identity ϵ_{T ⊗v}(− ⊗ (T ⊗ v)) ∘ (η_{T ⊗v}) ⊗ (T ⊗ v) = id_{−⊗(T ⊗v)} evaluated at V(T, V(v, w)) give the desired identification.

In the proof of 6.1, we used the next result.

Proposition A.8. The two maps X 𝕋(X) coincide.

Proof. Most for notational convenience, we will often leave out associativity and commutativity constraints. The two maps in question are determined

by Xv ^{η_{T} Xv} V (T, Xv ⊗ T) ^{V (T,sw_{T}^{X}(v))} V(T, X(T ⊗ v)) and (up to Yoneda isomorphism) τ_{v}(w): V(T ⊗ v, w) ⊗ T ^{f⊗T} V(T, V(v, w)) ⊗ T ^{ϵ_{T} V(v,w)} V(v, w). The Yoneda isomorphism

Xv ^{≅} ∫_{Ob C} V(V(v, w), Xw)

is induced by the natural transformation y_{v}^{X}(w): Xv V(V(v, w), Xw), that is, the composition

Xv  \overline{^{η_{V(v,w)}Xv}} V(V(v, w), Xv ⊗ V(v, w))
\overline{^{V(V(v,w),Xv⊗hom_{v,w}^{X})}} V(V(v, w), Xv ⊗ V(Xv, Xw))
\overline{^{V(V(v,w),σ_{Xv,V(Xv,Xw)})}} V(V(v, w), V(Xv, Xw) ⊗ Xv)
\overline{^{V(V(v,w),ϵ_{Xv}Xw)}} V(V(v, w), Xw).

Documenta Mathematica 8 (2003) 409–488

Hence it suffices to prove that, for every w ∈ Ob C, the composition

a(w) := y_{T ⊗v}^{V(T,X)}(w) ∘ V(T, sw_{T}^{X}(v)) ∘ η_{T} Xv

coincides with the composition b(w) := h ∘ V(τ_{v}, Xw) ∘ y_{v}^{X}(w). Here h denotes the adjointness isomorphism V(V(T ⊗ v, w) ⊗ T, Xw) V(V(T ⊗ v, w), V(T, Xw)). The isomorphism f: V(T, V(v, w)) V(T ⊗ v, w) will be used in the proof. The diagram

[figure: V(V(v, w), Xw); hom^{V(T,−)}; V(V(T, V(v, w)), V(T, Xw)); V(V(T ⊗ v, w)) ⊗ T, Xw); V(τ_{v},Xw); V(f⊗T,Xw); V(V(T, V(v, w)) ⊗ T, Xw); ≅]

commutes by A.3, where the vertical map on the right hand side is the adjointness isomorphism. Then by naturality and A.4, b(w) coincides with the composition V(V(T, V(v, w)), V(T, ϵ_{Xv}Xw∘hom_{v,w}^{X})∘sw_{Xv}^{V(T,−)})∘η_{V(T,V(v,w))}Xv. The diagram

Xv  \overline{^{η_{V(T,V(v,w))}Xv}} V(V(T, V(v, w)), Xv ⊗ V(T, V(v, w)))

[figure: V(T, Xv ⊗ T); η_{T} Xv; V(V(T, V(v, w)), V(T, Xv ⊗ T) ⊗ V(T, V(v, w))); V(V(T,V(v,w)),(η_{T} Xv) ⊗V(T,V(v,w)))]

[figure: V(T, X(T ⊗ v)); V(T, sw_{T}^{X}(v)); η_{V(T,V(v,w))}; V(V(T, V(v, w)), V(T, X(T ⊗ v)) ⊗ V(T, V(v, w))); V(V(T,V(v,w)),V(T,sw_{T}^{X}(v)) ⊗V(T,V(v,w)))]

commutes by naturality. Hence the maps a(w) and b(w) coincide if

V(T, V(v, w)) ⊗ Xv \overline{^{V(T,V(v,w))⊗η_{T} Xv}} V(T, V(v, w)) ⊗ V(T, Xv ⊗ T)

[figure: V(T, Xv ⊗ V(v, w)); sw_{Xv}^{V(T,−)}; V(T, V(v, w)) ⊗ V(T, X(T ⊗ v)); V(T,V(v,w))⊗ V(T,sw_{T}^{X} (v))]

[figure: V(T, Xv ⊗ V(Xv, Xw)); V(T,Xv⊗ hom_{v,w}^{X}); V(T ⊗ v, w) ⊗ V(T, X(T ⊗ v)); f⊗ V(T,X(T ⊗v))]

[figure: V(T, V(Xv, Xw) ⊗ Xv); V(T,σ); V(X(T ⊗ v), Xw) ⊗ V(T, X(T ⊗ v)); hom^{X} ⊗ V(T,X(T ⊗v))]

[figure: V(T, Xw); V(T,ϵ_{Xv}Xw); ϵ_{V(T,X(T ⊗v))}; V(V(T, X(T ⊗ v)), V (T, Xw)) ⊗ V (T, X(T ⊗ v)); hom^{V(T,−)}⊗ V(T,X(T ⊗v))]

commutes for all w. Now use A.5 (with A = X(T ⊗v) and B = Xw), naturality of the map sw^{V(T,−)} and the isomorphism f to replace the composition from the upper corner on the right hand side to the lower corner on the left hand side. The result is the diagram

V(T, V(v, w)) ⊗ Xv \overline{^{V(T,V(v,w))⊗η_{T} Xv}} V(T, V(v, w)) ⊗ V(T, Xv ⊗ T)

[figure: V(T, Xv ⊗ V(v, w)); sw_{Xv}^{V(T,−)}; V(T,Xv⊗ϵ_{T} V(v,w)); V(T, V(T, V(v, w)) ⊗ X(T ⊗ v)); sw_{V(T,V(v,w))}^{V(T,−)}]

[figure: V(T, Xv ⊗ V(Xv, Xw)); V(T,Xv⊗ hom_{v,w}^{X}); V(T, V(T ⊗ v, w) ⊗ X(T ⊗ v)); V(T,f⊗X(T ⊗v))]

[figure: V(T, V(Xv, Xw) ⊗ Xv); V(T,σ); V(T, V(X(T ⊗ v), Xw) ⊗ X(T ⊗ v)); V(T,hom^{X} ⊗X(T ⊗v))]

[figure: V(T, Xw).; V(T,ϵ_{X(T ⊗v)}Xw); V(T,ϵ_{Xv}Xw)]

The upper part commutes by A.6 (with A = Xv and B = V(v, w)), the lower part is V(T, −) applied to a diagram which commutes by A.7. This completes the proof.

References

[1] J. F. Adams. Stable homotopy and generalised homology. University of Chicago Press, Chicago, Ill., 1974. Chicago Lectures in Mathematics.

[2] Francis Borceux. Handbook of categorical algebra. 2. Cambridge University Press, Cambridge, 1994. Categories and structures.

[3] Gunnar Carlsson. A survey of equivariant stable homotopy theory. Topology, 31(1):1–27, 1992.

[4] Brian Day. On closed categories of functors. In Reports of the Midwest Category Seminar, IV, pages 1–38. Springer, Berlin, 1970.

[5] Bjørn Ian Dundas, Oliver Röndigs, and Paul Arne Østvær. Motivic functors. Documenta Math., 8:489–525, 2003.

[6] Philip S. Hirschhorn. Model categories and their localizations. American Mathematical Society, Providence, RI, 2003.

[7] Mark Hovey. Model categories. American Mathematical Society, Providence, RI, 1999.

[8] Mark Hovey. Spectra and symmetric spectra in general model categories. J. Pure Appl. Algebra, 165(1):63–127, 2001.

[9] Mark Hovey, Brooke Shipley, and Jeff Smith. Symmetric spectra. J. Amer. Math. Soc., 13(1):149–208, 2000.

[10] L. Gaunce Lewis, Jr., J. Peter May, and Mark Steinberger. Equivariant stable homotopy theory. Volume 1213 of Lecture Notes in Math.. Springer, Berlin, 1986.

[11] Manos Lydakis. Simplicial functors and stable homotopy theory. Preprint 98-049, SFB 343, Bielefeld, June 1998.

[12] Saunders MacLane. Categories for the working mathematician. Springer-Verlag, New York, 1971. Graduate Texts in Mathematics, Vol. 5.

[13] Michael A. Mandell and J. Peter May Equivariant orthogonal spectra and S-modules Mem. Amer. Math. Soc.. 159(755):x+108, 2002.

[14] Michael A. Mandell, J. Peter May, Stefan Schwede, and Brooke E. Shipley. Model categories of diagram spectra. Proc. London Math. Soc. (3), 82(2):441–512, 2001.

[15] Stefan Schwede and Brooke E. Shipley. Algebras and modules in monoidal model categories. Proc. London Math. Soc. (3), 80(2):491–511, 2000.

[16] Vladimir Voevodsky. 𝔸^{1}-homotopy theory. In Proceedings of the International Congress of Mathematicians, Vol. I (Berlin, 1998), pages 579–604 (electronic), 1998.

Bjørn Ian Dundas

Department of Mathematical Sciences The Norwegian University of Science and Technology Trondheim, Norway dundas@math.ntnu.no Oliver Röndigs Department of Mathematics The University of Western Ontario London, Ontario, Canada oroendig@uwo.ca

Paul Arne Østvær

Department of Mathematics University of Oslo Oslo, Norway paularne@math.uio.no